### v2

# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

Abstract.
We consider the configuration space of points on the two-dimensional sphere that satisfy a specific system of quadratic equations. We construct periodic orbits in this configuration space using elliptic theta functions and show that they satisfy semi-discrete analogues of mKdV and sine-Gordon equations. The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle. Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system generates an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229. All authors are partially supported by JST CREST Grant No. JPMJCR1911.
1. Introduction
Given a simple finite graph $(V,E)$ with edge weights $\ell:E\to\mathbb{R}$, its $d$-dimensional realisation is a map $\phi:V\to\mathbb{R}^{d}$ that satisfies the condition $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$. This mapping embeds the graph into $d$-dimensional Euclidean space while preserving the prescribed edge lengths. The collection of all possible realisations forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$. This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [10, 21, 25]. The primary motivation for this work is rooted in the analysis of the motion of a particular class of linkage mechanisms called Kaleidocycles (Fig. 1 (Left)). These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring. A state of a Kaleidocycle corresponds to a realisation of the graph formed by the vertices and edges of the tetrahedra (Eq. (2.1)), where $\ell$ is determined by the shape of the tetrahedron. Its motion, disregarding self-collision, is described by a path in the configuration space. Although this object has been studied by several authors [8, 28, 30], even its existence beyond simple classical cases has not been clarified. Our work addresses this gap by providing an explicit construction of the object and its motion.
Our key observation is to associate each state of a Kaleidocycle with a closed spatial polygonal curve. This curve is formed by connecting the midpoints of the hinge edges. The angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the curvature angle (§2.1). In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle. We can then identify the configuration space of a given Kaleidocycle with a subspace of a direct product of two-spheres (2.2). Note that it is not clear if there exists $\ell$ (the shape of the tetrahedra) for a given $k$ such that the configuration space is non-empty; that is, if a Kaleidocycle exists. We construct a one-dimensional subspace of the configuration space using elliptic theta functions, proving that there is a Kaleidocycle for any $k\geq 6$ that can move. The idea is to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
It was known in [19] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations [19] are defined on the configuration space (see §3.2). Therefore, a one-dimensional orbit in the configuration space is generated through the time evolution defined by the system of ordinary differential equations. The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [11]. We show that our construction of the motion of the Kaleidocycle (§5) indeed satisfies the system of the ordinary differential equations described in [19]. The sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 16].
Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, Kaleidocycle. It provides an interesting example, where a flow generated by an integrable system on a real-algebraic set is explicitly integrated using elliptic theta functions.
The structure of the paper is as follows. Section 2 introduces the definition of Kaleidocycles. As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame. The state of a Kaleidocycle is identified with a solution to a system of quadratic equations corresponding to the geometric constraints on the binormal vectors. Section 3 reviews the connection between the theory of curves and integrable systems. The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented. A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled. Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions. The resulting curve is shown to have a constant torsion angle. A condition for the resulting curve to be closed is also discussed; for a given $k$, the appropriate choice of parameters yields a Kaleidocycle. By introducing an additional parameter, a family of time-parametrised Kaleidocycles is constructed in Section 5. We then study the time evolution equations satisfied by the vertex positions and curvature angle of the Kaleidocycles and show connections to previous research. Section 6 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures. We present two conjectures that are suggested by numerical experiments. The Kaleidocycles constructed in this paper seem to be Möbius Kaleidocycles (Conjecture 6.1).
Most of the proofs of our theorems and propositions are deferred to the appendix since they require technically-involved computations regarding elliptic theta functions.
2. Algebraic and geometric models of Kaleidocycle
2.1. Configuration space of Kaleidocycle
A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring. It can be folded from a sheet of paper and makes an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid. The origin of Kaleidocycle is obscure but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R. For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [29] and the references therein).
We first define a generalisation of Kaleidocycle following [18]. We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring. Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces. Let $l_{01},l_{12},l_{20}$ be the lengths of the sides of the triangle face. A state of a $k$-Kaleidocycles of type $(l_{01},l_{12},l_{20})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k-1})\in(\mathbb{R}^{3})^{2k}$ of the quadratic equations
(2.1) $|x_{2n}-x_{2n+1}|^{2}$ $=l^{2}_{01},$
$|x_{2n}-x_{2n+2}|^{2}$ $=|x_{2n+1}-x_{2n+3}|^{2}=l^{2}_{02},$
$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12}.$
Note that collision between tetrahedra is ignored. The solution space of (2.1) admits an action of the Euclidean transformation group $E(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $63\cdot 2k-5k-\dim(E(3))=k-$ 6. This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [8]. The Bricard 6R is a case of $k=6$. Its configuration space has dimension one, which is greater than the expected dimension of 00, giving an example of an overconstrained  [8]. When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [20].Some of the findings of [20] are also published in [30]. For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of underconstrained linkage mechanisms whose configuration space has less dimension than the expected one.
The state of a Kaleidocycle can be equivalently described by a spatial polygon or a closed discrete spatial curve. Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k-1)$ be the mid-point of the hinge edge. We set $\gamma_{k}=\gamma_{0}$. Since the tetrahedron is equifacial, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $\widetilde{B}_{n}$ and $\widetilde{B}_{n+1}$, where $\widetilde{B}_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$ is the hinge direction. By reordering the vertices, we can assume the inner product $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ is constant for $1\leq n<k$. Denote this constant by $\cos\lambda$ with $0<\lambda<\pi$, which we call the torsion angle. We have two cases $\langle\widetilde{B}_{k-1},\widetilde{B}_{0}\rangle=\pm\cos\lambda$. We say the Kaleidocycle is oriented (resp. non-oriented) when the value is $+\cos\lambda$ (resp. $-\cos\lambda$). We set $\widetilde{B}_{k}=\pm\widetilde{B}_{0}$ so that $\langle\widetilde{B}_{k-1},\widetilde{B}_{k}\rangle=\cos\lambda$. Furthermore, since $\widetilde{B}_{n+1}\times\widetilde{B}_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, we have $\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$. Conversely, the vertex coordinates $x_{n}$ can be reconstructed from $\{\gamma_{n}\}$, $\{\widetilde{B}_{n}\}$ and $(l_{01},l_{12},l_{20})$ up to translation. Therefore,
Definition 1.
The configuration space of the $k$-Kaleidocycles with the torsion angle $\lambda$ is the set $\{(\widetilde{B}_{0},\ldots,\widetilde{B}_{k})\in(S^{2})^{k+1}\}$ satisfying the following conditions:
(2.2) $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ $=\cos\lambda,\quad(1\leq n\leq k)$
$\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})$ $=0,$
$\widetilde{B}_{k}$ =±$=\pm\widetilde{B}_{0}.$
A motion of a Kaleidocycle is identified with a one-dimensional (time-parametrised) orbit in the space defined by (2.1) or (2.2).
Figure 1. (Left) The classical 6-Kaleidocycle. (Right) Kaleidocycle as a discrete curve with its framing.
2.2. Discrete spatial curve of a constant torsion angle
A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals. Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle\neq 0$ for all $n\in\mathbb{Z}$. A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows. Define the tangent $T:\mathbb{Z}\rightarrow S^{2}$ and the normal $\widetilde{N}:\mathbb{Z}\rightarrow S^{2}$ by
(2.3) $T_{n}=\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|},\quad\widetilde{N}_{n}=\widetilde{B}_{n}\times T_{n}.$
Furthermore, given $\gamma_{0}\in\mathbb{R}^{3}$, define $\gamma:\mathbb{Z}\to\mathbb{R}^{3}$ by
(2.4) $\gamma_{n+1}=\gamma_{n}+T_{n}=\gamma_{n}+\frac{\widetilde{B}_{n+1}\times\widetilde{B}_{n}}{\left|\widetilde{B}_{n+1}\times\widetilde{B}_{n}\right|}.$
The signed curvature angle $\kappa$ and the torsion angle are defined by
(2.5) $\begin{split}\langle T_{n},T_{n-1}\rangle&=\cos\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
The discrete Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]\in\mathrm{SO}(3)$ satisfies the discrete Frenet-Serret formula
(2.6) $\Phi_{n+1}=\Phi_{n}L_{n},\quad L_{n}=M_{1}(-\lambda_{n+1})M_{3}(\kappa_{n+1}),$
where
(2.7) $M_{1}(\theta)=\left[\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\\ \end{array}\right],\quad M_{3}(\theta)=\left[\begin{array}[]{ccc}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{array}\right].$
Remark 2.1.
Note that our Frenet frame coincides with the usual one only up to sign. We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes the state when the curvature vanishes.
As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles. Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen. The fact that a Kaleidocycle is closed enforces interesting topological constraints. For example, the self-linking number must be a half-integer, and hence, stay constant during the motion. By Călugăreanu-Fuller-White’s theorem [1], the writhing number of the curve should stay constant as well. The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
3. Curves and integrable systems
There are deep connections between differential geometry and the theory of integrable systems, where the integrable differential or difference equations arise as compatibility conditions of the geometric objects (see, for example, [27]). In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations. The compatibility condition yields so-called the Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example. In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [9, 23, 27]. The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves. For example, continuous deformations of discrete plane and spatial curves have been considered in [5], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations. Discrete deformations of the discrete plane and spatial curves have been studied in [16]. In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
The theta functions naturally appear in so-called the quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions. Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 24]. Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [11].
Since the dynamics of the Kaleidocycles are modelled by the integrable deformation of the closed discrete spatial curves with a constant torsion angle which preserves the segment-length and torsion angle [19], we study the theta function solutions to the corresponding integrable semi-discrete equations, where the theta functions are nothing but the $\tau$ functions. Therefore we construct the explicit formula of the Kaleidocycles based on the method of direct reconstruction of the discrete spatial curves in terms of the $\tau$ functions.
Hereafter, a (discrete) curve always indicates a polygonal spatial curve.
3.1. Expression of a curve by the $\tau$ functions
We construct a discrete curve $\gamma_{n}$ and its Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [11]. Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n}$
(3.2) $\beta F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1}$
(3.3) $\beta F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1}$
(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n}$
(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*}$
(3.6) $\left|G_{n}\right|^{2}=\widetilde{G}_{n}^{2},$
where $\beta\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [12]. These functions are called the $\tau$ functions. Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}f_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0}$
(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vspace{2mm}\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vspace{2mm}\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vspace{2mm}\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
Similarly to [11, §3], we obtain:
Proposition 3.1.
$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying (2.6). Its curvature and torsion angles are given by
(3.23) $\lambda_{n}$ $=\left.\frac{1}{i}\log\left(\frac{G_{n}^{*}G_{n-1}}{\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0}=\left.\frac{1}{2i}\log\left(\frac{G_{n}^{*}G_{n-1}}{G_{n}G_{n-1}^{*}}\right)\right|_{z=0},$
$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta F_{n}}\right)\right|_{z=0}.$
3.2. Motion of Kaleidocycle on the osculating plane
Let $\gamma$ be a Kaleidocycle, that is, a closed curve with a constant torsion angle and a unit segment length. A motion of a Kaleidocycle is identified with a time-parametrised family $\gamma:\mathbb{Z}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{3}$ such that $\gamma(t)$ is a Kaleidocycle for all $t\geq 0$. In [19], a particular motion of a Kaleidocycle satisfying the following conditions is considered:
(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
where $\tilde{\rho}>0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$. These conditions mean that the velocity of each vertex is restricted within the osculating plane and has a constant magnitude. It is shown in [19] that the Frenet-Serret formula (2.6) together with (3.24) demand that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
or the semi-discrete sine-Gordon equation
(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
The motions corresponding to these equations differ in general. However, the explicit construction discussed in §5 satisfies both equations simultaneously.
Remark 3.2.
As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (2.2) is generally $k-6$, and hence, (3.24) may look too restrictive. However, the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one regardless of $k$. An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (2.2). Therefore, (2.2) specifies a particular yet essential motion of Kaleidocycles.
4. An explicit formula of closed discrete curve with a constant torsion angle in terms of elliptic theta functions
Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle. Then, we give sufficient conditions for the curve to be closed and derive an explicit expression for the motion of Kaleidocycles.
4.1. A family of open curves with a constant torsion angle
Throughout this section, $j$ runs through $\{1,2,3,4\}$. Fix $w\in\mathbb{C}$ with $\mathrm{Im}(w)>0$. The elliptic theta functions $\vartheta_{j}:\mathbb{C}\to\mathbb{C}$, are defined as follows [26, 22]:
(4.1) $\vartheta_{1}(v;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(v+\frac{1}{2}\right)},$
$\vartheta_{2}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)v},$
$\vartheta_{3}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inv},$
$\vartheta_{4}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(v+\frac{1}{2}\right)}.$
We often denote $\vartheta_{j}(v;w)$ by $\vartheta_{j}(v)$ when $w$ is fixed.
We construct the $\tau$ functions satisfying (3.1)-(3.6) using these elliptic theta functions such that the torsion angle given by Proposition 3.1 is a constant. The following observations are useful insights for constructing solutions. In order to make $F$ a real-valued function, it is convenient to use a property of elliptic theta functions, shown in Lemma K.1. Therefore, $w$ is taken to be a pure imaginary number. Additionally, to make $\gamma$ a closed curve, it is convenient to use the pseudo-period of elliptic theta functions (A.12), and therefore the space variable is taken to be a pure imaginary number. A real parameter $r$ is introduced to prevent the curve from being a planar curve. Then, the solution to satisfy (3.1)-(3.6) can be constructed from the identities of elliptic theta functions (e.g. (A.2)), and the result is as shown in Theorem 4.1. For $x\in\mathbb{C}$ and $y>0$, we define $d_{j}(x;y)=\frac{\vartheta^{{}^{\prime}}_{j}(x;iy)}{\vartheta_{j}(x;iy)}$, where ${}^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions. For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(4.2) $\alpha_{2}=\sqrt{\vartheta_{3}(2r;iy)\vartheta_{3}(0;iy)},\quad\beta=\frac{\vartheta_{3}(iv;iy)}{\vartheta_{3}(0;iy)},\quad s_{2}=\vartheta_{3}(2r;iy)\vartheta_{1}(iv;iy),$
$u_{2}=\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right),$
$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),$
$R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r;iy\right)}.$
Noting $\vartheta_{j}(z;iy)^{*}=\vartheta_{j}(z^{*};iy)$ when $y>0$, we have
(4.3) $\begin{split}&\alpha_{2}^{*}=\alpha_{2},\quad\beta^{*}=\beta,\quad s_{2}^{*}=-s_{2},\quad u_{2}^{*}=u_{2},\\ &\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}R_{j}^{*}=1.\end{split}$
Theorem 4.1.
Put $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$ and define
(4.4) $F_{n}(z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
$f_{n}(z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
$g_{n}(z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
$G_{n}(z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
$H_{n}(z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
$\widetilde{G}_{n}(z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right).$
Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(4.5) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(0)G_{n-1}(0)}{\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right).$
In particular, this does not depend on $n$, and $\gamma$ has a constant torsion angle $\lambda_{n}=\lambda_{0}$. Moreover, the signed curvature angle is given by
(4.6) $\begin{split}&\kappa_{n}=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$.
A proof of Theorem 4.1 is given in Appendix B.
4.2. Sufficient conditions for closedness
The constant torsion-angle and unit segment-length curves given in Theorem 4.1 are not necessarily closed. We derive a sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle. We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}=\gamma_{n}$ for all $n$. Rewriting this condition using the $\tau$ functions, we obtain
(4.7) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
$n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
It is difficult to give general solutions to (4.7), but we can utilise the pseudo-period of elliptic theta functions (A.12) to obtain particular solutions. For a fixed $k$, choose $m\in\mathbb{Z}\setminus{\{0\}}$ and set $v=my/k$. Then, from (A.12), we see
(4.8) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
$F_{n+k}=F_{n}\exp\left(k\frac{\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\exp\left[-\pi i\left\{2m\left(\mu_{n}-\frac{1}{2}iv\right)+m^{2}w\right\}\right],$
where $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Using (4.8), (4.7) can be rewritten as
(4.9) $\Lambda:=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
$\Xi:=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
Note that since
(4.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi$
holds, (4.9) can also be written as
(4.11) $\Lambda=0,\quad\frac{\partial}{\partial r}\Lambda=0.$
From the above discussion we obtain the following theorem.
Theorem 4.2.
For a fixed $k$, let $m\in\mathbb{Z}\setminus{\{0\}}$, $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$ satisfy $kv=my$ and (4.9). Then, the curve defined by Theorem 4.1 is closed and forms a $k$-Kaleidocycle with a torsion angle given by (4.5).
Proof.
The closure condition (4.7) holds for any $n\in\mathbb{Z}$ when the parameters $v,r,y$ satisfy $kv=my$ and (4.9). Thus, the conditions $kv=my$ and (4.9) give sufficient conditions for the curve given by Theorem 4.1 to be a closed curve. ∎
Hereafter, we assume that the parameters $v,r,y$ are taken to satisfy the sufficient conditions for closedness.
Example 4.3.
The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [8]) is obtained by $(v,r,y,m)\approx(0.640,0.250,1.279,3)$.
Remark 4.4.
By using (4.9) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. From this, we also see that $m$ and $Lk$ have the same parity. Furthermore, under the transformation $\widetilde{B}_{n}\rightarrow(-1)^{n+1}\widetilde{B}_{n}$, $\widetilde{N}_{n}\rightarrow(-1)^{n+1}\widetilde{N}_{n}$, the signed curvature angle becomes $(-1)^{n+1}\kappa_{n}$ and the torsion angle becomes $\pi+\lambda$. This indicates that this transformation corresponds to the one described in Proposition 5.16.
5. Construction of the motion of a Kaleidocycle
By introducing the deformation parameter to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem 5.1. We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature. Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [16] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [19] are discussed and summarised in Table 2.
5.1. Explicit formula of the discrete curve with deformation parameter
We introduce the deformation parameter so that only the signed curvature angle changes with a constant torsion angle and while satisfying (3.1)-(3.6). Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve. The following theorem describes a Kaleidocycle together with its motion explicitly in terms of elliptic theta functions.
Theorem 5.1.
Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$.
(5.1) $F_{n}(t,z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$f_{n}(t,z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
$g_{n}(t,z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
$G_{n}(t,z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$H_{n}(t,z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
$\widetilde{G}_{n}(t,z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
where $C,\Gamma\in\mathbb{R}$ are constants. Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(5.2) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)}{\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right),$
In particular, this depends neither on $n$ nor on $t$, and $\gamma$ has a constant torsion angle $\lambda_{n}(t)=\lambda_{0}(0)$. Moreover, the signed curvature angle is given by
(5.3) $\begin{split}&\kappa_{n}\left(t\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$. If $r,v,y$ are chosen as in Theorem 4.2, $\gamma$ is closed for every $t$.
A proof of Theorem 5.1 is given in Appendix B.
Proposition 5.2.
The parameter $C$ contributes only to the parallel translation of the curve along the $Z$-axis, whereas $\Gamma$ contributes only to the rotation of the curve around the $Z$-axis.
Proof.
For $A\in\mathbb{C}$ and $B:\mathbb{C}\to\mathbb{C}$, we have
(5.4) $\begin{split}&\left(\begin{array}[]{ccc}\cos\Gamma t&-\sin\Gamma t\\ \sin\Gamma t&\cos\Gamma t\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\frac{1}{2}\left(\begin{array}[]{ccc}e^{i\Gamma t}+e^{-i\Gamma t}&-\frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)\\ \frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)&e^{i\Gamma t}+e^{-i\Gamma t}\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\left(\begin{array}[]{ccc}e^{i\Gamma t}A+e^{-i\Gamma t}A^{*}\\ \frac{1}{i}\left(e^{i\Gamma t}A-e^{-i\Gamma t}A^{*}\right)\end{array}\right)\end{split}$
and
(5.5) $\begin{split}2\frac{\partial\log\left(e^{\frac{C}{2}tz}B\right)}{\partial z}=Ct+2\frac{\partial\log B}{\partial z}.\end{split}$
The assertion follows from the definition (3.10). ∎
In Theorem 5.9, we will make convenient choices of $C$ and $\Gamma$ to derive the deformation equation considered in the previous studies presented in §3.2.
Table 1 summarises the relationship between various variables. The first column lists the parameters that determine the curve $\gamma$ through the variables listed in the first row. A circle in the table indicates that the variable shown in the first row is dependent on the parameter shown in the first column.
Table 1. Dependency of variables on the parameters $v,r,y,k,m,t,C,\Gamma$.
$\gamma$ $\kappa$ $\lambda$ $\Lambda$ $\Xi$ $u_{2}$ $\beta$ $s_{2}$ $R_{j}$ $\Delta_{j}$ $\alpha_{2}$ $\beta_{1,2}$ $\epsilon_{1,2}$ $\omega$ $\delta_{1,2,3}$
$v$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$r$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$y$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$k$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$m$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$t$circ $\circ$
$\Gamma$circ $\circ$
$C$circ
(3.10) (5.3) (5.2) (4.9) (4.9) (4.2) (4.2) (4.2) (4.2) (4.2) (4.2) (5.6) (5.6) (5.6) (5.6)
The constructed motion has a periodic curvature. The next proposition follows from (A.12).
Proposition 5.3.
We have $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right)$. In other words, $2y$ is a period of the motion.
Remark 5.4.
There are certain relations between the curves constructed in Theorem 5.1 with different sets of parameters. The transformation of parameters $(v,m)\rightarrow(-v,-m)$ preserves the closure condition. We see that $\kappa_{n}\left(t;-v\right)=-\kappa_{n}\left(-t;v\right),\lambda(-v,r)=-\lambda(v,r)$. In what follows, We always choose $(v,m)$ or $(-v,-m)$ so that $\lambda\geq 0$.
5.2. Motion of vertex positions
The construction in Theorem 5.1 involves parameters $C$ and $\Gamma$ that contribute only to the global rigid transformation of the curve. In general, the velocity of the vertex position $\dot{\gamma_{n}}$ has a binormal component, and its magnitude is not constant as described in Theorem 5.6. However, with the right choice of $C$ and $\Gamma$, curves that satisfy (3.24) are obtained as described in Theorem 5.9.
We derive bilinear equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
For $\Gamma\in\mathbb{R},y>0,r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
(5.6) $\omega=\frac{\vartheta_{1}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)}{\vartheta_{3}\left(2r;iy\right)\vartheta_{1}\left(iv;iy\right)},$
$\delta_{1}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)^{3}},$
$\delta_{2}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{1}\left(iv;iy\right)^{2}\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)},$
$\delta_{3}=\frac{-iu_{2}}{s_{2}\vartheta_{3}(0;iy)\vartheta_{1}(iv;iy)}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right.$
$\left.+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right],$
$\epsilon_{1}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0;iy)}{\vartheta_{3}(0;iy)},$
$\epsilon_{2}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0;iy)^{2}}{\vartheta_{1}(iv;iy)^{2}},$
$\beta_{1}=\frac{1}{2}\left(\delta_{3}+\delta_{3}^{*}\right)+\frac{1}{2}i\Gamma\left(\delta_{2}-\delta_{2}^{*}\right),$
$\beta_{2}=\frac{1}{2}i\left(\delta_{3}-\delta_{3}^{*}\right)-\frac{1}{2}\Gamma\left(\delta_{2}+\delta_{2}^{*}\right).$
Then, it follows that
(5.7) $\omega^{*}=-\omega,\quad\delta_{1}^{*}=\delta_{1},\quad\epsilon_{1}^{*}=\epsilon_{1},\quad\epsilon_{2}^{*}=\epsilon_{2},\quad\beta_{1}^{*}=\beta_{1},\quad\beta_{2}^{*}=\beta_{2}.$
A proof of the following proposition is given in Appendix C.
Proposition 5.5.
The $\tau$ functions (5.1) satisfy the following bilinear equations.
(5.8) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}+\omega G_{n}G_{n}^{*},$
(5.9) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\omega G_{n}G_{n}^{*},$
(5.10) $\delta_{2}f_{n}^{*}g_{n-1}-\delta_{2}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
(5.11) $\delta_{1}F_{n}F_{n}+\delta_{2}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
(5.12) $\left(\beta-\delta_{1}\right)F_{n}F_{n}+\left(\omega-\delta_{2}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
(5.13) $D_{t}H_{n}\cdot F_{n}=\left(\delta_{3}+i\Gamma\delta_{2}\right)f_{n}^{*}g_{n-1}+\left(\delta_{3}^{*}-i\Gamma\delta_{2}^{*}\right)f_{n-1}^{*}g_{n},$
(5.14) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\epsilon_{1}+C\right)F_{n}F_{n}+\epsilon_{2}G_{n}G_{n}^{*}.$
By using these equations, we obtain the evolution equation of the vertex positions of the curve.
Theorem 5.6.
The curve $\gamma$ constructed in Theorem 5.1 satisfies the following deformation equation.
(5.15) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta\left[\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right).\end{split}$
A proof of Theorem 5.6 is given in Appendix D. This is a deformation equation similar to the one presented in [16]. Starting from (5.15), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
Proposition 5.7.
When the curve with a unit segment length and a constant torsion angle deforms according to the deformation equation (5.15), the following holds.
  1. (1)
    The segment length is preserved.
  2. (2)
    $\kappa$ satisfies the following semi-discrete mKdV equation:
    (5.16) $\dot{\kappa}_{n}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
Furthermore, if the left-hand-side of the equation (5.27) does not depend on $n$, the following holds:
  1. (3)
    The torsion angle is preserved.
Proof.
The proofs of (i) and (ii) are given by the same calculation as in Section 5 of [16]. We prove (iii). From the deformation equation (5.15) and the discrete Frenet-Serret formula (2.6), we see that
(5.17) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta\beta_{1}\sin\lambda-2\beta\beta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
holds, which vanishes because of (5.27). Therefore $\dot{\lambda}=0$.
Remark 5.8.
Proposition 5.7 (iii) does not hold in general without the assumption on (5.27). We can construct a motion of a Kaleidocycle satisfying (5.15) such that the left-hand side of (5.27) depends both on $n$ and $t$.
By adjusting the global motion of $\gamma$ by $C$ and $\Gamma$, the deformation equation of $\gamma$ can be made to satisfy the equation considered in the previous studies presented in §3.2.
Theorem 5.9.
Set $C,\Gamma$ to be
(5.18) $\begin{split}\Gamma_{+}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)+\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{+}=&-\epsilon_{1}-i\Gamma_{+}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega},\\ \rm or\\ \Gamma_{-}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)-\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{-}=&-\epsilon_{1}-i\Gamma_{-}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega}.\end{split}$
Then, the curve constructed in Theorem 5.1 satisfies the following deformation equations.
(5.19) $\begin{split}&\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2},\end{split}$
where $\rho^{2}>0$ is given by
(5.20) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2}$
when $(C,\Gamma)=(C_{+},\Gamma_{+})$ and
(5.21) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2}$
when $(C,\Gamma)=(C_{-},\Gamma_{-})$.
A proof of Theorem 5.9 is given in Appendix D.
Remark 5.10.
A discrete analogue of the constant negative curvature surface, called the discrete K-surface, is introduced in [2]. It is discussed that certain discrete-time deformation of discrete curves with a constant torsion angle generates discrete K-surfaces in [16]. Deformation equation (3.24) can be seen as a continuous-time analogue of the discrete-time deformation. Theorem 5.9 shows that specialising $(C,\Gamma)$ to $(C_{-},\Gamma_{-})$ or $(C_{+},\Gamma_{+})$, the curve constructed in Theorem 5.1 deforms according to (3.24). Therefore, its sweeping surface forms a semi-discrete analogue of the discrete K-surface (see figures in §6).
Since the curvature and torsion angles uniquely determine a curve up to a rigid transformation, Proposition 5.3 shows that the motion constructed in Theorem 5.1 is periodic up to a global rigid transform. Here, we control the global orientation by choosing the parameters $C$ and $\Gamma$ appropriately so that the motion becomes periodic in terms of vertex positions.
Proposition 5.11.
Let $l\in\mathbb{Z}$. If we set
(5.22) $C$ $=\frac{4\pi i}{y}\frac{1}{\Delta_{3}-\Delta_{1}},$
(5.23) $\Gamma$ $=-\frac{\pi}{y}\left(4r+2l+1\right),$
we have
(5.24) ${\gamma_{n}}(t+y)={\gamma_{n}}(t),$
for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$.
Proof.
First, we prove the periodicity for the $Z$-component. Denote the $Z$-component of $\gamma_{n}$ by ${\gamma_{n}}_{Z}$. From the pseudo-period of elliptic theta functions (A.12), we see
(5.25) $\begin{split}&{\gamma_{n}}_{Z}(t+y)-{\gamma_{n}}_{Z}(t)\\ &=-2\frac{\partial}{\partial z}\log\left\{\exp\left(\frac{C}{2}yz\right)\frac{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it+iy\right)}{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it\right)}\right\}\\ &=-2\frac{\partial}{\partial z}\log\exp\left\{\frac{C}{2}yz-\pi i\left(2iv\left(n-\frac{1}{2}\right)+\frac{2z}{\Delta_{3}-\Delta_{1}}+2it\right)+\pi y\right\}\\ &=-2\left(\frac{C}{2}y-\frac{2\pi i}{\Delta_{3}-\Delta_{1}}\right)=0.\end{split}$
For the $X$ and $Y$ components, from (3.10) we have to show $\frac{H_{n}(t+y)}{F_{n}(t+y)}\frac{F_{n}(t)}{H_{n}(t)}=1$.
If we write $P_{n}=iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it$, the left-hand side can be computed using (A.12) as
(5.26) $\begin{split}&\exp\left(i\Gamma y\right)\frac{\vartheta_{4}\left(P_{n}-2r+iy\right)\vartheta_{2}\left(P_{n}\right)}{\vartheta_{2}\left(P_{n}+iy\right)\vartheta_{4}\left(P_{n}-2r\right)}\\ &=-\exp\left(i\Gamma y\right)\frac{\exp\left\{-\pi i\left(2P_{n}-4r+iy\right)\right\}}{\exp\left\{-\pi i\left(2P_{n}+iy\right)\right\}}\\ &=-\exp\left\{i\left(\Gamma y+4\pi r\right)\right\}=1.\end{split}$
5.3. Evolution of curvature
The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$.
Proposition 5.12.
The signed curvature angle of the curve obtained in Theorem 5.1 satisfies the following recurrence equation for any $t$ and $n$:
(5.27) $\tan\frac{\kappa_{n+1}}{2}\tan\frac{\kappa_{n-1}}{2}-\tan^{2}\frac{\kappa_{n}}{2}+\tan\frac{\kappa_{n+1}}{2}\tan^{2}\frac{\kappa_{n}}{2}\tan\frac{\kappa_{n-1}}{2}=-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}.$
The right-hand side of this equation depends neither on $n$ nor on $t$. Therefore, the quantity on the left-hand side can be considered a conserved quantity of the signed curvature angle.
A proof of Proposition 5.12 is given in Appendix E. The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.7.
Theorem 5.13.
The signed curvature angle (5.3) satisfies
(5.28) $\dot{\kappa}_{n}=2\beta^{2}\eta_{3}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right),$
where
(5.29) $\eta_{3}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv)-d_{3}(iv)\right).$
Combined with Proposition 5.7, (5.6), and the fact that $e^{i\lambda}\delta_{2}\in\mathbb{R}$, we see
(5.30) $2\beta^{2}\eta_{3}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right).$
A proof of Theorem 5.13 is given in Appendix F.
5.4. Potential function of curvature
A family of functions $\Theta_{n}$ parameterised by $n\in\mathbb{Z}$ is said to be a potential of the curvature $\kappa$ if the equation $2\kappa_{n}=\Theta_{n+1}-\Theta_{n-1}$ holds. When $\gamma$ deforms by (5.19), the potential function $\Theta$ is known to satisfy the semi-discrete potential mKdV equation or the semi-discrete sine-Gordon equation [19]. Here, we derive the evolution equations satisfied by the potential function for the curve constructed by Theorem 5.1.
We construct a potential function which simultaneously satisfy (3.25) and (3.26) in terms of elliptic theta functions.
Theorem 5.14.
For fixed $v,y$, define the function $\Theta$ as follows.
(5.31) $\Theta_{n}\left(t\right)=\Theta_{n}\left(t;v,y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}.$
Then we see that $\kappa$ defined in (5.3) satisfies
(5.32) $\kappa_{n}=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}.$
A proof of Theorem 5.14 is given in Appendix G. The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation.
Theorem 5.15.
Let $y>0$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$. The potential function $\Theta$ defined by (5.31) simultaneously satisfies the semi-discrete potential mKdV and sine-Gordon equations
(5.33) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\frac{L_{1}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
(5.34) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\frac{L_{2}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right),$
where $Q_{1},Q_{2},L_{1},L_{2}\in\mathbb{R}$ are defined by
(5.35) $Q_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad Q_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
(5.36) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)+Q_{2}(v,y),$
(5.37) $L_{2}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)-Q_{2}(v,y).$
A proof of Theorem 5.15 is given in Appendix H.
The following shows two equations are related by a parameter shift.
Proposition 5.16.
By the transformation of parameters $v\rightarrow v+y$ and $r\rightarrow\frac{1}{2}-r$, the signed curvature angle $\kappa$, the torsion angles $\lambda$, the potential functions $\Theta$, the coefficients of the semi-discrete mKdV equation $L_{1}/y$ and of the semi-discrete sine-Gordon equation $L_{2}/y$, the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ are transformed as
(5.38) $\kappa_{n}\left(t\right)$ $\rightarrow(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2}\right),\quad\lambda\rightarrow\pi+\lambda,$
(5.39) $\Theta_{n}\left(t\right)$ $\rightarrow(-1)^{n}\Theta_{n}\left(t-\frac{y}{2}\right),$
(5.40) $\frac{L_{1}}{y}$ $\rightarrow\frac{L_{2}}{y},\quad\frac{L_{2}}{y}\rightarrow\frac{L_{1}}{y},$
(5.41) $\widetilde{N}_{n}(t)$ $\rightarrow(-1)^{n+1}\widetilde{N}_{n}\left(t-\frac{y}{2}\right),\quad\widetilde{B}_{n}(t)\rightarrow(-1)^{n+1}\widetilde{B}_{n}\left(t-\frac{y}{2}\right).$
A proof of Proposition 5.16 is given in Appendix I. The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{n}\Theta_{n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.33) with one of the semi-discrete sine-Gordon equation (5.34). Proposition 5.16 shows that this equation-level transformation is also achieved at the solution level by transforming the parameters.
Remark 5.17.
Rewriting (5.28) in terms of the potential function $\Theta$, we obtain
(5.42) $\frac{d}{dt}\Theta_{n}=4\beta^{2}\eta_{3}\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
which is also called the semi-discrete potential mKdV equation [14, 13, 6]. In other words, the potential function $\Theta$ defined by (5.31) is a simultaneous solution of three differential-difference equations, (5.33), (5.34) and (5.42). It does not usually happen that these equations are simultaneously solved. For example, it can be checked that the 2 and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [6, 7] are not solutions of the semi-discrete potential mKdV equation.
The following describes an explicit relation between our construction and the previous work [19] reviewed in §3.2.
Proposition 5.18.
When $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), we have
(5.43) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
Similarly, when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), we have
(5.44) $\frac{L_{2}(v,y)}{y}\frac{1}{1-\cos\lambda}=-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}=\rho_{-}.$
A proof of Proposition 5.18 is given in Appendix J. Note the difference in approaches of [19] and this work; [19] starts with a geometric property observed in the motion of Kaleidocycles, and derive evolution equations, whereas we aim to construct a particular motion of Kaleidocycles and investigate the equations satisfied by them. They turn out to arrive at the same place; the latter gives an explicit solution to the equations derived in the former.
Table 2 summarises the relationships among the three deformation equations.
Table 2. Relation between the three deformation equations
Values of
$C,\Gamma$
Deformation equation of the curve Equation for potential function
$C,\Gamma\in\mathbb{R}$
$\dot{\gamma}_{n}=\footnotesize\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right)\normalsize$
$+2\beta\left(\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
$\dot{\Theta}_{n}=\left(4\beta\beta_{1}\cos\lambda+4\beta\beta_{2}\sin\lambda\right)$
×$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
$\downarrow$ Specialise the deformation equation by choosing rigid transformation parameters $C,\Gamma$.
$C_{+},\Gamma_{+}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{+}$
$\frac{\dot{\Theta}_{n+1}+\dot{\Theta}_{n}}{2}=\left(1+\cos\lambda\right)\rho_{+}$
×$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
$\updownarrow$ By replacing the parameters of the solution with $v\rightarrow v+y,r\rightarrow 1/2-r$,
these two deformation equations transform to each other.
$C_{-},\Gamma_{-}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{-}$
$\frac{\dot{\Theta}_{n+1}-\dot{\Theta}_{n}}{2}=\left(1-\cos\lambda\right)\rho_{-}$
×$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
6. Examples
In this section, we present examples of Kaleidocycles constructed by Theorem 4.2 and the semi-discrete K-surfaces obtained as their motion trajectory as explained in Remark 5.10.
Numerical experiments using our computer codes [17] suggest that the $8$-, $9$- and $15$-Kaleidocycles in Figs. 2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, $9$-, and $15$-Kaleidocycles, respectively [18]. In other words, they are the Möbius Kaleidocycles [20].
For Fig. 5, the torsion angle is larger than the above minimum one. However, it has a self-linking number $7.5$. As the self-linking number of a Kaleidocycle is a half-integer, these $15$-Kaleidocycles depicted in Figs. 4 and 5 belong to different connected components of the configuration space of $15$-Kaleidocycles. We numerically find that the latter attains the local minima of the torsion angle. We present the following two conjectures based on our extensive numerical experiments:
Conjecture 6.1.
Any Kaleidocycle obtained by Theorem 4.2 has a minimal or maximal torsion angle. In other words, the projection from the space defined by (2.2) onto the $\lambda$-coordinate attains a critical value.
Conjecture 6.2.
The self-linking number of the closed curve $\gamma$ constructed in Theorem 4.2 is $m/2$ when the curve is unknotted.
As briefly discussed in §1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [18, 20, 30]. Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
Figure 2. (Left) An example of 8-Kaleidocycle with $(v,r,y,m)\approx(0.400,0.717,1.067,3)$ and $\cos\lambda\approx 0.4700$. (Middle) 8-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 3. (Left) An example of 9-Kaleidocycle with $(v,r,y,m)\approx(0.342,0.289,1.027,3)$ and $\cos\lambda\approx 0.5852$. (Middle) 9-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 4. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.189,0.300,0.947,3)$ and $\cos\lambda\approx 0.8533$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 5. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.401,0.689,1.204,5)$ and $\cos\lambda\approx 0.6497$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Appendix
In the appendix, we prove the theorems and propositions in the paper by using the properties of elliptic theta functions listed in Appendix A.
Appendix A Functional identities of elliptic theta functions
We collect some properties of elliptic theta functions which we rely heavily on our calculations.
A.1. Four term identities
Let
(A.1) $\begin{split}&X_{1}=\frac{1}{2}(X+Y+U+V),\quad Y_{1}=\frac{1}{2}(X+Y-U-V),\\ &U_{1}=\frac{1}{2}(X-Y+U-V),\quad V_{1}=\frac{1}{2}(X-Y-U+V),\end{split}$
then for any $X,Y,U,V\in\mathbb{C}$, elliptic theta functions satisfy the following identities [26, 22].
(A.2) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})-\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{2}(X)\vartheta_{2}(Y)\vartheta_{3}(U)\vartheta_{3}(V)-\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.3) $\vartheta_{4}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{2}(V_{1})-\vartheta_{2}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{2}(V)-\vartheta_{2}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{4}(V),$
(A.4) $\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.5) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})+\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)+\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.6) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{1}(U)\vartheta_{1}(V),$
(A.7) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})$
$=\vartheta_{4}(X)\vartheta_{4}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.8) $\vartheta_{4}(X_{1})\vartheta_{4}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})+\vartheta_{2}(X_{1})\vartheta_{2}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{2}(U)\vartheta_{2}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{4}(U)\vartheta_{4}(V),$
(A.9) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{1}(U_{1})\vartheta_{1}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{3}(U_{1})\vartheta_{3}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{1}(U)\vartheta_{1}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{3}(U)\vartheta_{3}(V),$
(A.10) $\vartheta_{3}(X_{1})\vartheta_{3}(Y_{1})\vartheta_{4}(U_{1})\vartheta_{4}(V_{1})-\vartheta_{1}(X_{1})\vartheta_{1}(Y_{1})\vartheta_{2}(U_{1})\vartheta_{2}(V_{1})$
$=\vartheta_{3}(X)\vartheta_{3}(Y)\vartheta_{4}(U)\vartheta_{4}(V)-\vartheta_{1}(X)\vartheta_{1}(Y)\vartheta_{2}(U)\vartheta_{2}(V).$
A.2. Shifts by periods and half-periods
For any $X\in\mathbb{C}$, $n\in\mathbb{Z}$ and $y>0$, the following holds [26, 22].
(A.11) $\vartheta_{1}\left(X+n\right)=(-1)^{n}\vartheta_{1}\left(X\right),\quad\vartheta_{4}\left(X+n\right)=\vartheta_{4}\left(X\right).$
(A.12) $\begin{split}&\vartheta_{1}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{1}\left(X;iy\right),\\ &\vartheta_{2}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{3}\left(X+niy;iy\right)=\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{4}\left(X+niy;iy\right)=(-1)^{n}\exp\{-\pi i\left(2nX+n^{2}iy\right)\}\vartheta_{4}\left(X;iy\right).\end{split}$
(A.13) $\begin{split}&\vartheta_{1}\left(X+\frac{1}{2}\right)=\vartheta_{2}\left(X\right),\quad\vartheta_{2}\left(X+\frac{1}{2}\right)=-\vartheta_{1}\left(X\right),\\ &\vartheta_{3}\left(X+\frac{1}{2}\right)=\vartheta_{4}\left(X\right),\quad\vartheta_{4}\left(X+\frac{1}{2}\right)=\vartheta_{3}\left(X\right).\end{split}$
(A.14) $\begin{split}&\vartheta_{1}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{4}\left(X;iy\right),\\ &\vartheta_{2}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{3}\left(X;iy\right),\\ &\vartheta_{3}\left(X+\frac{iy}{2};iy\right)=\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{2}\left(X;iy\right),\\ &\vartheta_{4}\left(X+\frac{iy}{2};iy\right)=i\exp\left\{-\pi i\left(X+\frac{iy}{4}\right)\right\}\vartheta_{1}\left(X;iy\right).\end{split}$
A.3. Identity for theta constants
For any $y>0$, elliptic theta functions satisfy the following identity [26, 22].
(A.15) $\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)=\pi\vartheta_{2}\left(0;iy\right)\vartheta_{3}\left(0;iy\right)\vartheta_{4}\left(0;iy\right).$
A.4. Modular transformations
For any $X\in\mathbb{C}$ and $y>0$, elliptic theta functions satisfy the following identities [26, 22].
(A.16) $\begin{split}&\vartheta_{1}\left(\frac{X}{y};i\frac{2}{y}\right)=-i\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{1}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{2}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{4}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{3}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{3}\left(\frac{iX}{2};i\frac{y}{2}\right),\\ &\vartheta_{4}\left(\frac{X}{y};i\frac{2}{y}\right)=\sqrt{\frac{y}{2}}\exp\left(-\pi\frac{X^{2}}{2y}\right)\vartheta_{2}\left(\frac{iX}{2};i\frac{y}{2}\right).\end{split}$
(A.17) $\frac{\vartheta_{1}\left(X;iy\right)}{\vartheta_{2}\left(X;iy\right)}=-i\frac{\vartheta_{1}\left(\frac{iX}{y};\frac{i}{y}\right)}{\vartheta_{4}\left(\frac{iX}{y};\frac{i}{y}\right)}.$
A.5. Bilinear identities connecting theta functions with $y$ and $y/2$
For any $X,Y\in\mathbb{C}$, and $y>0$, elliptic theta functions satisfy the following identities [26, 22].
$2\vartheta_{1}\left(X+Y;iy\right)\vartheta_{4}\left(X-Y;iy\right)$
(A.18) $=\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)+\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right),$
$2\vartheta_{2}\left(X+Y;iy\right)\vartheta_{3}\left(X-Y;iy\right)$
(A.19) $=\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)-\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right).$
(A.20) $\begin{split}&\vartheta_{3}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{4}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{4}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ &\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{1}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ -&\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}+\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}=2\vartheta_{2}\left(X;iy\right)\vartheta_{3}\left(0;iy\right).\\ \end{split}$
Appendix B Proofs of Theorem 4.1 and Theorem 5.1
Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (3.1)-(3.6). Theorem 4.1 follows from Theorem 5.1 by setting $t=0$. We assume $z=0$ without loss of generality except for the proof of (3.5). Putting
(B.5) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+r+it,\\ \vspace{2mm}Y=ivn-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
in (A.2) yields (3.1). Putting
(B.10) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=ivn+r+it,\\ \vspace{2mm}U=iv,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
in (A.2) yields (3.2). Putting
(B.15) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=iv,\\ V=ivn-r+it,\\ \end{array}\right.$
in (A.3) yields (3.3). Putting
(B.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=-\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.4) yields
(B.21) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
Also, putting
(B.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.4) yields
(B.27) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
Subtracting (B.27) from (B.21) and substituting
(B.28) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
in the equation, we obtain
(B.29) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn-r+it\right),\end{split}$
from which we see (3.4) holds. Putting
(B.34) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+z+r+it,\\ \vspace{2mm}Y=ivn+z-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r+h,\\ V=\frac{1}{2}iv-r+h,\\ \end{array}\right.$
in (A.2) yields
(B.35) $\begin{split}&\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
Differentiating (B.35) with $h$ and then putting $h=0$, we obtain
(B.36) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\\ &+\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right).\end{split}$
Using the equation (B.35) with $h=0$ and substituting (B.36), we obtain
(B.37) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right.\\ &\left.+\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\right.\\ &\left.\times\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right\}\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}+\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z-r+it\right),\end{split}$
from which we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows.
Appendix C Proof of Proposition 5.5
For the proof of Proposition 5.5, we use the following lemma:
Lemma C.1 ([11, §3]).
The following equation holds:
(C.1) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}.$
C.1. Proof of Proposition 5.5
Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (5.8)-(5.14). We assume $z=0$ without loss of generality except for the proof of (5.14). Putting
(C.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=\frac{1}{2}iv-r,\\ \vspace{2mm}U=ivn+r+it,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
in (A.8) yields (5.8). Also, (5.9) follows from (C.1) and (5.8). Putting
(C.11) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r+h,\\ \vspace{2mm}U=\frac{1}{2}iv-r+h,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
in (A.3) yields
(C.12) $\begin{split}&\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it-h\right)\\ &=\vartheta_{4}\left(ivn-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(iv\left(n-1\right)-r+it\right)\\ &-\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(iv\left(n-1\right)-r+it\right),\end{split}$
After differentiating (C.12) by $h$ and putting $h=0$, we obtain (5.13) by rearrangement. Also, rearranging the equation with $h=0$ in (C.12) yields (5.10). Putting
(C.17) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r,\\ \vspace{2mm}U=iv\left(n-\frac{1}{2}\right)+it,\\ V=-iv\left(n-\frac{1}{2}\right)-it,\\ \end{array}\right.$
in (A.5) yields (5.11). Also, (5.12) follows from (5.8) and (5.11). Putting
(C.22) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}Y=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}U=h,\\ V=h,\\ \end{array}\right.$
in (A.2) yields
(C.23) $\begin{split}&\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}-h\right)\\ &=-\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}\\ &+\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
After differentiating (C.23) twice with $h$ and putting $h=0$, we obtain (5.14).
Appendix D Proofs of Theorem 5.6 and Theorem 5.9
For the proofs of Theorem 5.6 and Theorem 5.9, we need the following lemmas.
Lemma D.1 ([11, §3]).
Using (3.2) and (3.3), we see that the following equation holds.
(D.1) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
Lemma D.2.
The following relation holds.
(D.2) $\omega=\delta_{2}-\delta_{2}^{*}.$
Proof.
Putting
(D.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
in (A.9) yields (D.2). ∎
Lemma D.3.
The following identity holds.
(D.8) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=2\delta_{2}\delta_{3}^{*}+2\delta_{2}^{*}\delta_{3}+\epsilon_{2}=$ 0 .
Proof.
Putting
(D.13) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r+h,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r-h,\\ \vspace{2mm}U=\frac{1}{2}iv+r-h,\\ V=-\frac{1}{2}iv+r+h,\\ \end{array}\right.$
in (A.9) yields
(D.14) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
After differentiating (D.14) by $h$ and setting $h=0$, we obtain (D.8). ∎
For simplicity, we define $p$ and $q$ as follows
(D.15) $p_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
(D.16) $q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
Then we rewrite $D_{t}H_{n}\cdot F_{n}$ by using (5.6), (D.15) and (D.16) as follows
(D.17) $D_{t}H_{n}\cdot F_{n}=\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}.$
Also, we see the following hold.
(D.18) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}q_{n}^{*}+q_{n}\\ \vspace{2mm}\frac{1}{i}\left(q_{n}^{*}-q_{n}\right)\\ -2\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\end{array}\right),$
(D.19) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left(p_{n}-p_{n}^{*}\right)\\ \vspace{2mm}p_{n}+p_{n}^{*}\\ -\frac{2}{i}\omega|G_{n}|^{2}\end{array}\right).$
We have the following lemma.
Lemma D.4.
The following relations hold.
(D.20) $p_{n}p_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\omega^{2}|G_{n}|^{4},$
(D.21) $\begin{split}p_{n}q_{n}^{*}-p_{n}^{*}q_{n}=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
(D.22) $\begin{split}&q_{n}q_{n}^{*}=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+\left\{1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
Proof.
By definition of $p$, we see that
(D.23) $\begin{split}p_{n}p_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\end{split}$
holds. Substituting (D.1) and (5.9) into (D.23) yields (D.20). By definition of $p$ and $q$, we see that
(D.24) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\\ &=2\left(g_{n}^{*}f_{n-1}f_{n}^{*}g_{n-1}-g_{n}f_{n-1}^{*}f_{n}g_{n-1}^{*}\right)\\ \end{split}$
holds. On the other hand, from (C.1) and (5.9), the following holds.
(D.25) $\begin{split}&\beta\omega F_{n}^{2}|G_{n}|^{2}\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right).\\ \end{split}$
Using (D.25), we obtain
(D.26) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
Substituting (5.9), (5.11) and (5.12) into (D.26) gives
(D.27) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\omega|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
thus (D.21). Also from (D.1), (5.11) and (5.12), we see that
(D.28) $\begin{split}&q_{n}q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\delta_{1}F_{n}^{2}+\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta-\delta_{1}\right)F_{n}^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\\ &=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}\end{split}$
holds. From (D.28), we get (D.22). ∎
Using Proposition 5.5, Lemma D.2, Lemma D.3 and Lemma D.4, Theorem 5.6 and Theorem 5.9 can be proved as follows.
D.1. Proof of Theorem 5.6
By differentiating $\gamma$ with respect to $t$, we get
(D.29) $\begin{split}&\dot{\gamma}_{n}(t)=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -\left(\epsilon_{1}+C\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
On the other hand, from (D.18) and (D.19), we see
(D.30) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}+2i\beta_{2}\omega|G_{n}|^{2}\end{array}\right)\end{split}$
holds. From (D.8), (D.30) can be rewritten as follows
(D.31) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
Thus (5.15) is obtained.
D.2. Proof of Theorem 5.9
We compute the condition satisfied by $C$ and $\Gamma$ when
(D.32) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
(D.33) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
where $\rho\in\mathbb{R}$ is a constant. From (D.19) and (D.29) we see that
(D.34) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(p_{n}-p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)\\ &+\left(p_{n}+p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ &+2\omega|G_{n}|^{2}\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)\right\}\\ &+2\omega|G_{n}|^{4}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\end{split}$
holds. For (D.32) to be satisfied, it is sufficient that the following equations hold.
(D.35) $2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)=0,\$
(D.36) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=$ 0 .
By comparing (D.36) and (D.8) we see that (D.35) is a necessary and sufficient condition for the realization of (D.32). By substituting (5.6) into (D.35), it can be rewritten in terms of $C$ and $\Gamma$ as follows.
(D.37) $i\Gamma\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)+\epsilon_{1}+C+\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)-\frac{\delta_{3}-\delta_{3}^{*}}{\omega}=0.$
Also, since
(D.38) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)^{2}\\ &-\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)^{2}\\ &+\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\beta_{1}^{2}\delta_{1}\left(\beta-\delta_{1}\right)+\left(\epsilon_{1}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)\right.\\ &\left.-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}\right),\end{split}$
holds, we see that for (D.33) holds, it is sufficient that
(D.39) $4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)=0,$
(D.40) $-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}=0,$
hold. Since
(D.41) $\begin{split}&4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}+8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)-4\beta_{2}^{2}\omega^{2}-\epsilon_{2}^{2}\\ =&\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}-\epsilon_{2}\right\},\end{split}$
holds, (D.40) is identically satisfied. Therefore, (D.39) is a necessary and sufficient condition for the realization of (D.33). By using (D.8), (5.6) and (D.37), (D.39) can be written in terms of $\Gamma$ as follows.
(D.42) $i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)\Gamma^{2}+2\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\Gamma+i\left(\delta_{2}{\delta_{3}^{*}}^{2}-\delta_{2}^{*}\delta_{3}^{2}\right)=0.$
(D.42) has the following solutions.
(D.43) $\Gamma=\frac{-\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\pm\sqrt{\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}}}{i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)}.$
Since
(D.44) $\begin{split}\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\end{split}$
holds, (D.43) can be rewritten as
(D.45) $\Gamma=\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)\pm\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega}.$
Now, we prove equations (5.20) and (5.21). From (4.2) and (5.6), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
(D.46) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
(D.47) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
From (D.38), (D.46) and (D.47), we see that
(D.48) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2}$
holds. (D.48) is equivalent to (5.20). From (4.2) and (5.6), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
(D.49) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
(D.50) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{4}(r)^{4}-\vartheta_{2}(r)^{4}\right)^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}.$
From (D.38), (D.49) and (D.50), we see that
(D.51) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2}$
holds. (D.51) is equivalent to (5.21).
Appendix E Proof of Proposition 5.12
Here we prove that the signed curvature angle $\kappa$ satisfies (5.27). First, we see
Lemma E.1.
The $\tau$ functions (5.1) satisfy the following bilinear equations.
(E.1) $\widetilde{G}_{n}^{2}+\beta^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
(E.2) $\beta^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
Proof.
Using (3.1)-(3.3), we see that (E.1) hold. Putting
(E.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=0,\\ \vspace{2mm}Y=0,\\ \vspace{2mm}U=iv\left(n+\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ V=-iv\left(n-\frac{3}{2}\right)-it-\frac{z}{\Delta_{3}-\Delta_{1}},\\ \end{array}\right.$
in (A.10), we obtain (E.2).
E.1. Proof of Proposition 5.12
Using (E.1) and (E.2), we see that (5.27) holds.
Appendix F Proof of Theorem 5.13
Here, we show that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28). For the proof of Theorem 5.13, we first see
Lemma F.1.
The $\tau$ functions (5.1) satisfy the following bilinear equation.
(F.1) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\eta_{3}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right).$
Proof.
We assume $z=0$ without loss of generality. After differentiating the equation (A.3) with
(F.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n+\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=iv+h,\\ \vspace{2mm}U=-iv+h,\\ V=iv\left(n-\frac{3}{2}\right)+it\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(F.7) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}.\end{split}$
After rearranging (F.7), we obtain (F.1).
F.1. Proof of Theorem 5.13
Using (E.1) and (F.1), we see that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28).
Appendix G Proof of Theorem 5.14
From (A.17), (A.18) and (A.19), we see that
(G.1) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}\end{split}$
holds.
Appendix H Proof of Theorem 5.15
We first construct the solution to the semi-discrete potential mKdV equation (5.33) in two stages by proving Lemma H.1 and Lemma H.2. Then we construct the solution to the semi-discrete sine-Gordon equation (5.34) in two stages by proving Lemma H.3 and Lemma H.4. This is based on the construction of the solution to the mKdV equation using the theory of $\tau$ functions [15].
For fixed $v,y$, define the following real-valued functions
(H.1) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
Then the potential function can be written as follows
(H.2) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$. $\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy.
Lemma H.1.
Assuming that the solution of (5.33) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.33).
(H.3) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
(H.4) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
where $\eta_{1}$ is a real constant.
Proof.
Since
(H.5) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
(H.6) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right)\end{split}$
hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.33) is that
(H.7) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\eta_{1}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
is satisfied, where $\eta_{1}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(H.8) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
The real part of this equation yields (H.3), and the imaginary part yields (H.4). ∎
In Lemma H.2, we construct a solution of (H.3) and (H.4) in terms of elliptic theta functions.
Lemma H.2.
The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.3) and (H.4) with
(H.9) $\eta_{1}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)+Q_{4}(v,y)}{L_{1}(v,y)}.$
Proof.
In this proof, all elliptic theta functions have $w=2i/y$ and we denote $\vartheta_{j}(v;2i/y)$ by $\vartheta_{j}(v)$. After differentiating the equation in (A.6) with
(H.14) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.15) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.21) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.27) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
(H.32) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
by $h$ and putting $h=0$, we obtain
(H.33) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Adding (H.15) and (H.21) yields (H.3), and subtracting (H.27) from (H.33) yields (H.4). ∎
Next, we construct the solution to the semi-discrete sine-Gordon equation (5.34).
Lemma H.3.
Assuming that the solution of (5.34) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.34).
(H.34) $D_{t}A_{n+1}\cdot A_{n}-D_{t}B_{n+1}\cdot B_{n}=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right),$
(H.35) $D_{t}A_{n+1}\cdot B_{n}+D_{t}B_{n+1}\cdot A_{n}=\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right),$
where $\eta_{2}$ is a real constant.
Proof.
Since
(H.36) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}-\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}\right),\end{split}$
(H.37) $\begin{split}&\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}\right)\end{split}$
hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.34) is that
(H.38) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}=-\frac{L_{2}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}+\frac{L_{2}}{2y}\eta_{2}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n},\end{split}$
is satisfied, where $\eta_{2}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(H.39) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}-B_{n+1}\cdot B_{n}\right)+iD_{t}\left(A_{n+1}\cdot B_{n}+B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right)\\ &+i\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right).\end{split}$
The real part of (H.39) yields (H.34), and the imaginary part yields (H.35). ∎
In Lemma H.4, we construct a solution of (H.34) and (H.35) in terms of elliptic theta functions.
Lemma H.4.
The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.34) and (H.35) with
(H.40) $\eta_{2}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)-Q_{4}(v,y)}{L_{2}(v,y)}.$
Proof.
In this proof, we set $w=2i/y$ similarly to Lemma H.2. Subtracting (H.21) from (H.15) yields (H.34), and adding (H.33) and (H.27) yields (H.35). ∎
Appendix I Proof of Proposition 5.16
We first show (5.40) and (5.39) by using the periodicity of elliptic theta functions (A.11). Consider replacing the parameter with $\tilde{v}=v+y$. Then we see
(I.1) $\begin{split}&Q_{1}(\tilde{v},y)=-Q_{1}(v,y),\quad Q_{2}(\tilde{v},y)=-Q_{2}(v,y),\\ &d_{1}(\tilde{v},y)=d_{1}(v,y),\quad d_{4}(\tilde{v},y)=d_{4}(v,y),\\ &L_{1}(\tilde{v},y)=L_{2}(v,y).\end{split}$
and
(I.2) $\begin{split}&\Theta_{n}\left(t;\tilde{v},y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}{\vartheta_{4}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}\\ &=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}\\ &=(-1)^{n}4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}\\ &=(-1)^{n}\Theta_{n}\left(t-\frac{y}{2};v,y\right).\end{split}$
From the relation between the signed curvature angle $\kappa$ and the potential function $\Theta$, we see
(I.3) $\begin{split}&\kappa_{n}\left(t;\tilde{v}\right)=(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2};v\right).\end{split}$
Next, we check the transformation of $\lambda$. From (A.13) and (A.14), we have
(I.4) $\begin{split}&\lambda\left(\tilde{v},\frac{1}{2}-r\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}{\vartheta_{1}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{2}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}{\vartheta_{2}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}\frac{\vartheta_{4}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}{\vartheta_{4}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(-\frac{\vartheta_{3}\left(-\frac{1}{2}iv-r\right)}{\vartheta_{3}\left(\frac{1}{2}iv-r\right)}\frac{\vartheta_{1}\left(\frac{1}{2}iv-r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv-r\right)}\right)\\ &=\pi+\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &=\pi+\lambda\left(v,r\right).\\ \end{split}$
Finally, we see that the transformation of the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ follows from the Frenet-Serret formula (2.6).
Appendix J Proof of Proposition 5.18
First, we calculate $L_{1}(v,y)/y$ and $L_{2}(v,y)/y$. From (5.36), (A.15), (A.16) and (A.20), we see
(J.1) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
Similarly, from (5.37), (A.15), (A.16) and (A.20), we see
(J.2) $\begin{split}\frac{L_{2}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}-i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)-\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
On the other hand, from (4.2) and (5.2), we see
(J.3) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}},\end{split}$
(J.4) $\begin{split}&1-\cos\lambda\\ &=-\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)-\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}}.\end{split}$
From (J.1) and (J.3), we see that (5.43) holds. Similarly, from (J.2) and (J.4), we see that (5.44) holds.
Appendix K Special property of elliptic theta functions when $w$ is a pure imaginary number
Lemma K.1.
The following holds for any $z\in\mathbb{C}$ when $w=iy$.
(K.1) $\vartheta_{j}(z)^{*}=\vartheta_{j}(z^{*}),\quad j=1,2,3,4.$
Proof.
Consider the following theta function with characteristics.
(K.2) $\vartheta_{a,b}(z):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z+b)+\pi i(n+a)^{2}\omega\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
then,
(K.3) $\vartheta_{1}(z)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(z),\quad\vartheta_{2}(z)=\vartheta_{\frac{1}{2},0}(z),\quad\vartheta_{3}(z)=\vartheta_{0,0}(z),\quad\vartheta_{4}(z)=\vartheta_{0,\frac{1}{2}}(z)$
hold. Since $w^{*}=-w$, we have
(K.4) $\begin{split}\vartheta_{a,b}(z)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(z^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(z^{*}).\end{split}$
This lemma guarantees that some $\tau$ functions, such as $F_{n}$, are real-valued functions.

### v3

@@
-footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)
+An explicit construction of Kaleidocycles by elliptic theta functions
 Abstract.
-We consider the configuration space of points on the two-dimensional sphere that satisfy a specific system of quadratic equations.
-We construct periodic orbits in this configuration space using elliptic theta functions and show that they satisfy semi-discrete analogues of mKdV and sine-Gordon equations.
+We consider the configuration space of ordered points on the two-dimensional sphere that satisfy a specific system of quadratic equations.
+We construct periodic orbits in this configuration space using elliptic theta functions and show that they simultaneously satisfy semi-discrete analogues of mKdV and sine-Gordon equations.
 The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle.
-Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system generates an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
-The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229. All authors are partially supported by JST CREST Grant No. JPMJCR1911.
+A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra greater than five.
+Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system is explicitly solved to generate an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
+Key words and phrases:
+Kinematic chain, Configuration space, Discrete differential geometry, Integrable systems, Elliptic theta function
+2000 Mathematics Subject Classification:
+Primary 53A04, 53A70; Secondary 53A17, 70B15, 37K25, 37K10, 35Q53
+The first-named author was partially supported by KAKENHI, Grant-in-Aid for Scientific Research (B) 25K00921 and (S) 25H00399.
+The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229.
 1. Introduction
-Given a simple finite graph $(V,E)$ with edge weights $\ell:E\to\mathbb{R}$, its $d$-dimensional realisation is a map $\phi:V\to\mathbb{R}^{d}$ that satisfies the condition $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$.
-This mapping embeds the graph into $d$-dimensional Euclidean space while preserving the prescribed edge lengths.
-The collection of all possible realisations forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$.
-This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [10, 21, 25].
-The primary motivation for this work is rooted in the analysis of the motion of a particular class of linkage mechanisms called Kaleidocycles (Fig.
+Given a natural number $k$ and a real number $1<\cos\lambda<1$, we consider the following space of configurations of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
+(1.1) $\mathcal{C}^{\pm}_{k,\lambda}=\{(b_{0},\ldots,b_{k})\in(S^{2})^{k+1}\mid$ $\langle b_{n-1},b_{n}\rangle=\cos\lambda,\,(1\leq n\leq k),$
+$\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=0,$
+$b_{k}=\pm b_{0}\}.$
+Our main result states that for any $k\geq 6$, there exists $\lambda$ such that a circle can be embedded in $\mathcal{C}^{\pm}_{k,\lambda}$.
+Moreover, we explicitly construct a circle in terms of elliptic theta functions.
+This result implies the existence of a particular class of linkage mechanisms called Kaleidocycles (Fig.
 1 (Left)).
 These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring.
-A state of a Kaleidocycle corresponds to a realisation of the graph formed by the vertices and edges of the tetrahedra (Eq.
-(2.1)), where $\ell$ is determined by the shape of the tetrahedron.
-Its motion, disregarding self-collision, is described by a path in the configuration space.
-Although this object has been studied by several authors [8, 28, 30], even its existence beyond simple classical cases has not been clarified.
-Our work addresses this gap by providing an explicit construction of the object and its motion.
-Our key observation is to associate each state of a Kaleidocycle with a closed spatial polygonal curve.
-This curve is formed by connecting the midpoints of the hinge edges.
-The angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the curvature angle (§2.1).
-In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle.
-We can then identify the configuration space of a given Kaleidocycle with a subspace of a direct product of two-spheres (2.2).
-Note that it is not clear if there exists $\ell$ (the shape of the tetrahedra) for a given $k$ such that the configuration space is non-empty; that is, if a Kaleidocycle exists.
-We construct a one-dimensional subspace of the configuration space using elliptic theta functions, proving that there is a Kaleidocycle for any $k\geq 6$ that can move.
-The idea is to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
-It was known in [19] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations [19] are defined on the configuration space (see §3.2).
-Therefore, a one-dimensional orbit in the configuration space is generated through the time evolution defined by the system of ordinary differential equations.
-The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [11].
-We show that our construction of the motion of the Kaleidocycle (§5) indeed satisfies the system of the ordinary differential equations described in [19].
-The sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 16].
-Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, Kaleidocycle.
-It provides an interesting example, where a flow generated by an integrable system on a real-algebraic set is explicitly integrated using elliptic theta functions.
+Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial ones has remained an open problem.
+Since an element of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a Kaleidocycle with $k$-tetrahedra in a specific position (Section 2.1) and its motion is described by a path in the configuration space, our main result provides an explicit construction and its periodic motion for Kaleidocycles of any $k\geq 6$.
+The key observation is to identify a Kaleidocycle with a closed spatial polygonal curve formed by connecting the midpoints of the hinge edges.
+Then, the angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the signed curvature angle (§2.1).
+An element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a sequence of binormals.
+In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle $\lambda$.
+In our construction, it plays the central role to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
+It was known in [21] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (see §3.2).
+The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12]5).
+Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 18].
+Both views combined, a flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ is explicitly integrated using elliptic theta functions.
+Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, the Kaleidocycle.
 The structure of the paper is as follows.
 Section 2 introduces the definition of Kaleidocycles.
+They are identified with a certain class of discrete spatial curves.
 As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame.
-The state of a Kaleidocycle is identified with a solution to a system of quadratic equations corresponding to the geometric constraints on the binormal vectors.
 Section 3 reviews the connection between the theory of curves and integrable systems.
@@
 The resulting curve is shown to have a constant torsion angle.
-A condition for the resulting curve to be closed is also discussed; for a given $k$, the appropriate choice of parameters yields a Kaleidocycle.
-By introducing an additional parameter, a family of time-parametrised Kaleidocycles is constructed in Section 5.
-We then study the time evolution equations satisfied by the vertex positions and curvature angle of the Kaleidocycles and show connections to previous research.
+In Section 5, a family of time-parametrised curves is constructed by introducing an additional parameter $t$.
+We then study the time evolution equations satisfied by the vertex positions and curvature angle of the curve and show connections to previous research.
+Section 6 investigates a necessary and sufficient condition for the resulting curve to be closed for every $t$.
+We prove there exists a set of parameters that satisfies the closure condition for a given $k\geq 6$, which proves the existence of Kaleidocycles.
 Section 7 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures.
-We present two conjectures that are suggested by numerical experiments.
-The Kaleidocycles constructed in this paper seem to be Möbius Kaleidocycles (Conjecture 6.1).
-Most of the proofs of our theorems and propositions are deferred to the appendix since they require technically-involved computations regarding elliptic theta functions.
+We present a conjecture (Conjecture 7.1) suggested by numerical experiments: the Kaleidocycles constructed in this paper seem to be so-called Möbius Kaleidocycles.
 2. Algebraic and geometric models of Kaleidocycle
-2.1. Configuration space of Kaleidocycle
+2.1. Configuration space of a Kaleidocycle
 A Kaleidocycle consists of (typically six) identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints to form a ring.
-It can be folded from a sheet of paper and makes an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid.
-The origin of Kaleidocycle is obscure but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R.
+It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid.
+The origin of Kaleidocycle is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R.
 For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein).
-We first define a generalisation of Kaleidocycle following [18].
+However, there is no literature that discusses the existence of such objects for arbitrary $k$.
+We first define a generalisation of the Kaleidocycle following [20].
 We consider $k$-copies of an equifacial tetrahedron in $\mathbb{R}^{3}$ connected by opposing edges to form a ring.
 Denote the vertex set by $\{0,1,\ldots,2k-1\}$ such that the vertices $(2n,2n+1)$ form a hinge edge and the vertices $(2n,2n+1,2n+2)$ form one of the triangle faces.
-Let $l_{01},l_{12},l_{20}$ be the lengths of the sides of the triangle face.
-A state of a $k$-Kaleidocycles of type $(l_{01},l_{12},l_{20})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k-1})\in(\mathbb{R}^{3})^{2k}$ of the quadratic equations
+Let $l_{01},l_{12},l_{02}$ be the lengths of the sides of the triangle face, with the edge of length $l_{01}$ serving as a hinge (see Figure 1 Right).
+Now, a natural way to model a Kaleidocycle is as a realisation of an edge-weighted graph.
+Let $(V,E)$ be the graph formed by the vertices and edges of the tetrahedra in a Kaleidocycle and $\ell:E\to\mathbb{R}$ be the edge length.
+A realisation of $(V,E,\ell)$ in $\mathbb{R}^{3}$ is a map $\phi:V\to\mathbb{R}^{3}$ satisfying $|\phi(u)-\phi(v)|^{2}=\ell(uv)^{2}$ for all edges $uv\in E$.
+In general, the collection of all possible realisations of an edge-weighted finite graph $(V,E,\ell)$ in $\mathbb{R}^{d}$ forms a real algebraic set in $\mathbb{R}^{d|V|}$, known as the configuration space of $(V,E,\ell)$.
+This space has been the subject of extensive research across various domains, including algebraic geometry, robotics, and mechanism theory [11, 23, 27].
+More concretely, a state of a $k$-Kaleidocycle of type $(l_{01},l_{12},l_{02})$ is identified by the real solution $(x_{0},x_{1},\ldots,x_{2k+1})\in(\mathbb{R}^{3})^{2k+2}$ of the quadratic equations
 (2.1) $|x_{2n}-x_{2n+1}|^{2}$ $=l^{2}_{01},$
 $|x_{2n}-x_{2n+2}|^{2}$ $=|x_{2n+1}-x_{2n+3}|^{2}=l^{2}_{02},$
-$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12}.$
-Note that collision between tetrahedra is ignored.
-The solution space of (2.1) admits an action of the Euclidean transformation group $E(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $63\cdot 2k-5k-\dim(E(3))=k-$ 6.
+$|x_{2n+1}-x_{2n+2}|^{2}$ $=|x_{2n}-x_{2n+3}|^{2}=l^{2}_{12},$
+for $0\leq n\leq k-2$ and $(x_{2k},x_{2k+1})=(x_{0},x_{1})$ or $(x_{1},x_{0})$. To remove the translational freedom, we demand $x_{0}+x_{1}=0$.
+Notice that we have two combinatorially distinct cases: length $l_{02}$ edges form a $k$-cycle ($x_{2k}=x_{0}$) or $2k$-cycle, going around the ring twice ($x_{2k}=x_{1}$).
+Remark 2.1.
+The solution space of (2.1) admits an action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $3\cdot 2k-5k-3-\dim(O(3))=k-6$, where $5k$ is the number of edges and $3$ comes from the equation $x_{0}+x_{1}=0$.
 This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [9].
+However, this estimate is sometimes violated in interesting examples.
 The Bricard 6R is a case of $k=6$.
-Its configuration space has dimension one, which is greater than the expected dimension of 00, giving an example of an overconstrained  [8].
-When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [20].Some of the findings of [20] are also published in [30].
-For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of underconstrained linkage mechanisms whose configuration space has less dimension than the expected one.
-The state of a Kaleidocycle can be equivalently described by a spatial polygon or a closed discrete spatial curve.
-Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k-1)$ be the mid-point of the hinge edge.
-We set $\gamma_{k}=\gamma_{0}$.
-Since the tetrahedron is equifacial, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $\widetilde{B}_{n}$ and $\widetilde{B}_{n+1}$, where $\widetilde{B}_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$ is the hinge direction.
-By reordering the vertices, we can assume the inner product $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ is constant for $1\leq n<k$.
-Denote this constant by $\cos\lambda$ with $0<\lambda<\pi$, which we call the torsion angle.
-We have two cases $\langle\widetilde{B}_{k-1},\widetilde{B}_{0}\rangle=\pm\cos\lambda$.
-We say the Kaleidocycle is oriented (resp.
-non-oriented) when the value is $+\cos\lambda$ (resp.
-$-\cos\lambda$).
-We set $\widetilde{B}_{k}=\pm\widetilde{B}_{0}$ so that $\langle\widetilde{B}_{k-1},\widetilde{B}_{k}\rangle=\cos\lambda$.
-Furthermore, since $\widetilde{B}_{n+1}\times\widetilde{B}_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, we have $\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$.
-Conversely, the vertex coordinates $x_{n}$ can be reconstructed from $\{\gamma_{n}\}$, $\{\widetilde{B}_{n}\}$ and $(l_{01},l_{12},l_{20})$ up to translation.
-Therefore,
-Definition 1.
-The configuration space of the $k$-Kaleidocycles with the torsion angle $\lambda$ is the set $\{(\widetilde{B}_{0},\ldots,\widetilde{B}_{k})\in(S^{2})^{k+1}\}$ satisfying the following conditions:
-(2.2) $\langle\widetilde{B}_{n-1},\widetilde{B}_{n}\rangle$ $=\cos\lambda,\quad(1\leq n\leq k)$
-$\sum_{n=0}^{k-1}(\widetilde{B}_{n+1}\times\widetilde{B}_{n})$ $=0,$
-$\widetilde{B}_{k}$ =±$=\pm\widetilde{B}_{0}.$
-A motion of a Kaleidocycle is identified with a one-dimensional (time-parametrised) orbit in the space defined by (2.1) or (2.2).
-Figure 1. (Left) The classical 6-Kaleidocycle. (Right) Kaleidocycle as a discrete curve with its framing.
+Its configuration space has dimension one, which is greater than the expected dimension of 0, giving an example of an overconstrained system  [9].
+When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [22].
+[^1] Some of the findings of [22] are also published in [31].
+For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of an underconstrained linkage mechanism whose configuration space has a smaller dimension than the expected one.
+We use the following elementary geometric property of equifacial tetrahedron:
+Lemma 2.2.
+An equifacial tetrahedron whose faces are congruent triangles with edge-lengths $(l_{01},l_{12},l_{02})$ may be realised within a rectangular cuboid of side lengths
+$\sqrt{\frac{l_{01}^{2}+l_{12}^{2}-l_{02}^{2}}{2}},\quad\sqrt{\frac{l_{12}^{2}+l_{02}^{2}-l_{01}^{2}}{2}},\quad\sqrt{\frac{l_{02}^{2}+l_{01}^{2}-l_{12}^{2}}{2}}$
+so that tetrahedron edges are diagonals of the faces. Moreover, the twist angle between the pair of opposite edges of length $l_{01}$ is given by the relation
+$\cos(\lambda)=\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}.$
+We will see the solution space of Equation 2.1 is identified with $\mathcal{C}^{\pm}_{k,\lambda}$.
+Proposition 2.3.
+The space of real solutions to Equation 2.1 is $O(3)$-equivariantly homeomorphic to $\mathcal{C}^{\pm}_{k,\lambda}$ with $\lambda=\arccos\left(\frac{|l_{12}^{2}-l_{02}^{2}|}{l_{01}^{2}}\right)$.
+Proof.
+Given $(x_{0},\ldots,x_{2k+1})$ satisfying Equation 2.1, we define $b_{n}=\dfrac{x_{2n+1}-x_{2n}}{|x_{2n+1}-x_{2n}|}$.
+By 2.2, the inner product $\langle b_{n-1},b_{n}\rangle$ is constant $\cos\lambda$ for $1\leq n<k$.
+Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge.
+There exists $l>0$ such that $l|b_{n+1}\times b_{n}|=|\gamma_{n+1}-\gamma_{n}|$.
+By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$.
+Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$.
+This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
+Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=l\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron.
+We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$.
+Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
+It is clear that the above constructions are mutually inverse. ∎
+While the shape of the tetrahedron is determined by the triangle edge lengths $(l_{01},l_{12},l_{02})$, the configuration space depends only on the hinge angle $\lambda$.
+We call $\lambda$ the torsion angle.
+We call the (possible non-flat) polygon, or closed discrete spatial curve, defined by $\gamma_{n}$ the central curve of the Kaleidocycle.
+Depending on the sign of $b_{k}=\pm b_{0}$, we say the Kaleidocycle is oriented (resp. non-oriented).
+Figure 1. (Left) The classical 6-Kaleidocycle. (Centre) Kaleidocycle as a discrete curve with its framing (Right) Net of a Kaleidocycle with labelled vertices and edge lengths.
 2.2. Discrete spatial curve of a constant torsion angle
+We saw that a discrete spatial curve $\gamma_{n}$ is associated with a Kaleidocycle. We review some basics on discrete curves.
 A discrete spatial curve with non-vanishing torsion angles can be described in terms of binormals.
-Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle\neq 0$ for all $n\in\mathbb{Z}$.
+Let $\widetilde{B}:\mathbb{Z}\rightarrow S^{2}$ satisfy $\left|\widetilde{B}_{n}\times\widetilde{B}_{n-1}\right|\neq 0$ for all $n\in\mathbb{Z}$.
 A discrete spatial curve of unit segment-length having $\widetilde{B}$ as the binormals can be constructed as follows.
@@
 The signed curvature angle $\kappa$ and the torsion angle are defined by
-(2.5) $\begin{split}\langle T_{n},T_{n-1}\rangle&=\cos\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
+(2.4) $\begin{split}\langle T_{n},T_{n-1}\rangle=\cos\kappa_{n},\quad\langle\widetilde{N}_{n},T_{n-1}\rangle&=-\sin\kappa_{n},\quad-\pi<\kappa_{n}<\pi,\\ \langle\widetilde{B}_{n},\widetilde{B}_{n-1}\rangle&=\cos\lambda_{n},\quad 0<\lambda_{n}<\pi.\end{split}$
 The discrete Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]\in\mathrm{SO}(3)$ satisfies the discrete Frenet-Serret formula
@@
 Note that our Frenet frame coincides with the usual one only up to sign.
-We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes the state when the curvature vanishes.
+We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes through the state where the curvature vanishes.
 As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles.
 Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen.
+In this language, a Kaleidocycle is equivalent to a discrete closed ($\gamma_{k}=\gamma_{0})$ spatial curve with constant torsion angle $\lambda$ determined by the binormal $\widetilde{B}_{n}=b_{n}$.
+Our approach is to utilise this identification.
 The fact that a Kaleidocycle is closed enforces interesting topological constraints.
 For example, the self-linking number must be a half-integer, and hence, stay constant during the motion.
-By Călugăreanu-Fuller-White’s theorem [1], the writhing number of the curve should stay constant as well.
+By the Călugăreanu-Fuller-White theorem [1], the writhing number of the curve must remain constant as well.
 The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
@@
 In the context of the deformation of spatial and plane curves, the Frenet frame and its deformation are described by the system of linear partial differential equations.
-The compatibility condition yields so-called the Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example.
+The compatibility condition yields the so-called Ablowitz-Kaup-Newell-Segur (AKNS) hierarchy for the curvature and the torsion, including the nonlinear Schrödinger (NLS) equation as a typical example.
 In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29].
@@
 In principle, for a given solution to the nonlinear differential/difference equation for the curvature and torsion, one can reconstruct the curve by solving the associated system of linear differential/difference equations for the Frenet frame and integrating the tangent vector.
-The theta functions naturally appear in so-called the quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions.
+The theta functions naturally appear in so-called quasi-periodic solutions to the nonlinear integrable equations, and the corresponding solutions to the associated linear problem are nothing but the Baker-Akhiezer functions.
 Reconstruction of the deformation of the spatial curve with constant torsion described by the theta function solutions to the mKdV equation has been studied in this framework in [4, 26].
 Also, an explicit formula of the spatial curves in terms of the $\tau$ functions of the two-component KP hierarchy based on Hirota’s bilinear formalism has been presented in [12].
-Since the dynamics of the Kaleidocycles are modelled by the integrable deformation of the closed discrete spatial curves with a constant torsion angle which preserves the segment-length and torsion angle [19], we study the theta function solutions to the corresponding integrable semi-discrete equations, where the theta functions are nothing but the $\tau$ functions.
-Therefore we construct the explicit formula of the Kaleidocycles based on the method of direct reconstruction of the discrete spatial curves in terms of the $\tau$ functions.
-Hereafter, a (discrete) curve always indicates a polygonal spatial curve.
 3.1. Expression of a curve by the $\tau$ functions
-We construct a discrete curve $\gamma_{n}$ and its Frenet frame $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [11].
+We introduce a special class of discrete curves $\gamma_{n}$ and their Frenet frames $\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ indexed by $n\in\mathbb{Z}$ using a modification of the $\tau$ functions presented in [12].
 Let $F_{n},\widetilde{G}_{n}:\mathbb{C}\to\mathbb{R}$ and $f_{n},g_{n},G_{n},H_{n}:\mathbb{C}\to\mathbb{C}$ be functions indexed by $n\in\mathbb{Z}$ with a complex variable $z$ that satisfy the following equations
-(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n}$
-(3.2) $\beta F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1}$
-(3.3) $\beta F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1}$
-(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n}$
-(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*}$
+(3.1) $f_{n}f_{n}^{*}+g_{n}g_{n}^{*}=F_{n+1}F_{n},$
+(3.2) $\beta_{1}F_{n}f_{n}+G_{n}g_{n}^{*}=F_{n+1}f_{n-1},$
+(3.3) $\beta_{1}F_{n}g_{n}-G_{n}f_{n}^{*}=F_{n+1}g_{n-1},$
+(3.4) $H_{n+1}F_{n}-H_{n}F_{n+1}=f_{n}^{*}g_{n},$
+(3.5) $D_{z}F_{n+1}\cdot F_{n}:=\frac{dF_{n+1}}{dz}F_{n}-F_{n+1}\frac{dF_{n}}{dz}=g_{n}g_{n}^{*},$
 (3.6) $\left|G_{n}\right|^{2}=\widetilde{G}_{n}^{2},$
-where $\beta\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [12].
+where $\beta_{1}\in\mathbb{R}$ is some constant and $D_{z}$ is Hirota’s bilinear differential operator [13], and the symbol * denotes the complex conjugate.
 These functions are called the $\tau$ functions.
 Let $\varrho_{n}=\dfrac{1}{2}\dfrac{G_{n}}{\widetilde{G}_{n}}$ and
-(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}f_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0}$
-(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vspace{2mm}\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
-(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vspace{2mm}\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0}$
-(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vspace{2mm}\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vspace{2mm}\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
-Similarly to [11, §3], we obtain:
+(3.10) $T_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054ptf_{n}^{*}g_{n}+f_{n}g_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left(f_{n}^{*}g_{n}-f_{n}g_{n}^{*}\right)\\ f_{n}f_{n}^{*}-g_{n}g_{n}^{*}\end{array}\right)\right|_{z=0},$
+(3.14) $\widetilde{N}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \vskip 5.69054pt\frac{1}{i}\left[\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}-\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ -2\left(f_{n}^{*}g_{n}^{*}\varrho_{n}+f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
+(3.18) $\widetilde{B}_{n}$ $=\left.\frac{1}{f_{n}f_{n}^{*}+g_{n}g_{n}^{*}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left[-\left\{(f_{n}^{*})^{2}-(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}-(g_{n})^{2}\right\}\varrho_{n}^{*}\right]\\ \vskip 5.69054pt\left\{(f_{n}^{*})^{2}+(g_{n}^{*})^{2}\right\}\varrho_{n}+\left\{(f_{n})^{2}+(g_{n})^{2}\right\}\varrho_{n}^{*}\\ \frac{2}{i}\left(f_{n}^{*}g_{n}^{*}\varrho_{n}-f_{n}g_{n}\varrho_{n}^{*}\right)\end{array}\right)\right|_{z=0},$
+(3.22) $\gamma_{n}$ $=\left.\left(\begin{array}[]{c}\vskip 5.69054pt\dfrac{H_{n}+H_{n}^{*}}{F_{n}}\\ \vskip 5.69054pt\dfrac{1}{i}\dfrac{H_{n}-H_{n}^{*}}{F_{n}}\\ n-2\dfrac{\partial\log F_{n}}{\partial z}\end{array}\right)\right|_{z=0}.$
+Similarly to [12, §3], we obtain:
 Proposition 3.1.
-$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying (2.6). Its curvature and torsion angles are given by
-(3.23) $\lambda_{n}$ $=\left.\frac{1}{i}\log\left(\frac{G_{n}^{*}G_{n-1}}{\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0}=\left.\frac{1}{2i}\log\left(\frac{G_{n}^{*}G_{n-1}}{G_{n}G_{n-1}^{*}}\right)\right|_{z=0},$
-$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta F_{n}}\right)\right|_{z=0}.$
-3.2. Motion of Kaleidocycle on the osculating plane
+$\Phi_{n}=[T_{n},\widetilde{N}_{n},\widetilde{B}_{n}]$ defines the discrete Frenet frame for the curve $\gamma$ satisfying the discrete Frenet-Serret formula (2.5).
+Its curvature and torsion angles are given by
+(3.23) $\lambda_{n}$ $=\left.\arccos\left(\frac{G_{n}^{*}G_{n-1}+G_{n}G_{n-1}^{*}}{2\widetilde{G}_{n}\widetilde{G}_{n-1}}\right)\right|_{z=0},$
+$\kappa_{n}$ $=\left.2\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)\right|_{z=0},$
+where we take the principal value for $\arccos$ in $0,\pi]$ and for $\arctan$ in $(-\pi/2,\pi/2)$.
+3.2. Motion of a Kaleidocycle on the osculating plane
 Let $\gamma$ be a Kaleidocycle, that is, a closed curve with a constant torsion angle and a unit segment length.
@@
 (3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
-where $\tilde{\rho}>0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$.
-These conditions mean that the velocity of each vertex is restricted within the osculating plane and has a constant magnitude.
-It is shown in [19] that the Frenet-Serret formula (2.6) together with (3.24) demand that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
-(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
-or the semi-discrete sine-Gordon equation
-(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\tilde{\rho}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
-The motions corresponding to these equations differ in general. However, the explicit construction discussed in §5 satisfies both equations simultaneously.
+where $\rho\neq 0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$.
+These conditions imply that the velocity of each vertex is restricted to the osculating plane and has a constant magnitude.
+It is shown in [21] that the Frenet-Serret formula (2.5), together with (3.24), requires that there exists a function $\Theta_{n}$, called the potential function, with $\kappa_{n}=\dfrac{\Theta_{n+1}-\Theta_{n-1}}{2}$ that satisfies either the semi-discrete potential mKdV equation:
+(3.25) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\left(1+\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
+or the semi-discrete sine-Gordon equation:
+(3.26) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\left(1-\cos\lambda\right)\rho\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
+The motions corresponding to these equations differ in general. However, we will construct a motion in §5 that satisfies both equations simultaneously.
 Remark 3.2.
-As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (2.2) is generally $k-6$, and hence, (3.24) may look too restrictive.
-However, the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one regardless of $k$.
+As noted in §2.1, the dimension of the configurations of a Kaleidocycle defined by (1.1) is generally k6k-6.
+However, for the Möbius Kaleidocycles, which exhibit critical torsion angles, the dimension of the configurations is conjectured to be one, regardless of $k$.
 An extensive numerical simulation suggests that the motion of the Möbius Kaleidocycle satisfies (3.24).
 Therefore, (3.24) specifies a particular yet essential motion of Kaleidocycles.
-4. An explicit formula of closed discrete curve with a constant torsion angle in terms of elliptic theta functions
-Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle.
-Then, we give sufficient conditions for the curve to be closed and derive an explicit expression for the motion of Kaleidocycles.
-4.1. A family of open curves with a constant torsion angle
+4. An explicit formula of a discrete curve with a constant torsion angle in terms of elliptic theta functions
+Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in Theorem 4.2.
+4.1. A family of curves with a constant torsion angle
 Throughout this section, $j$ runs through $\{1,2,3,4\}$.
@@
 The elliptic theta functions $\vartheta_{j}:\mathbb{C}\to\mathbb{C}$, are defined as follows [28, 24]:
-(4.1) $\vartheta_{1}(v;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(v+\frac{1}{2}\right)},$
-$\vartheta_{2}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)v},$
-$\vartheta_{3}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inv},$
-$\vartheta_{4}(v;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(v+\frac{1}{2}\right)}.$
-We often denote $\vartheta_{j}(v;w)$ by $\vartheta_{j}(v)$ when $w$ is fixed.
-We construct the $\tau$ functions satisfying (3.1)-(3.6) using these elliptic theta functions such that the torsion angle given by Proposition 3.1 is a constant.
-The following observations are useful insights for constructing solutions.
-In order to make $F$ a real-valued function, it is convenient to use a property of elliptic theta functions, shown in Lemma K.1.
-Therefore, $w$ is taken to be a pure imaginary number.
-Additionally, to make $\gamma$ a closed curve, it is convenient to use the pseudo-period of elliptic theta functions (A.12), and therefore the space variable is taken to be a pure imaginary number.
-A real parameter $r$ is introduced to prevent the curve from being a planar curve.
-Then, the solution to satisfy (3.1)-(3.6) can be constructed from the identities of elliptic theta functions (e.g.
-(A.2)), and the result is as shown in Theorem 4.1.
-For $x\in\mathbb{C}$ and $y>0$, we define $d_{j}(x;y)=\frac{\vartheta^{{}^{\prime}}_{j}(x;iy)}{\vartheta_{j}(x;iy)}$, where ${}^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions.
+(4.1) $\vartheta_{1}(u;w)$ $=-\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)\left(u+\frac{1}{2}\right)},$
+$\vartheta_{2}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iw\left(n+\frac{1}{2}\right)^{2}+2\pi i\left(n+\frac{1}{2}\right)u},$
+$\vartheta_{3}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi inu},$
+$\vartheta_{4}(u;w)$ $=\sum_{n=-\infty}^{\infty}e^{\pi iwn^{2}+2\pi in\left(u+\frac{1}{2}\right)}.$
+We often denote $\vartheta_{j}(u;w)$ by $\vartheta_{j}(u)$ when $w$ is fixed.
+These functions satisfy various identities, collected in Appendix A, that are convenient to construct the $\tau$ functions satisfying (3.1)–(3.6).
+We see the torsion angle given by Proposition 3.1 is a constant.
+The following property is useful for examining whether a $\tau$ function constructed in this paper is real-valued.
+Lemma 4.1.
+The following holds for any $u\in\mathbb{C}$ when $w=iy$ $(y>0)$.
+(4.2) $\vartheta_{j}(u)^{*}=\vartheta_{j}(u^{*}).$
+Proof.
+Consider the following theta function with characteristics.
+(4.3) $\vartheta_{a,b}(u):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u+b)+\pi i(n+a)^{2}w\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
+then,
+(4.4) $\vartheta_{1}(u)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(u),\quad\vartheta_{2}(u)=\vartheta_{\frac{1}{2},0}(u),\quad\vartheta_{3}(u)=\vartheta_{0,0}(u),\quad\vartheta_{4}(u)=\vartheta_{0,\frac{1}{2}}(u),$
+hold. Since $w^{*}=-w$, we have
+(4.5) $\begin{split}\vartheta_{a,b}(u)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(u^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(u^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(u^{*}).\end{split}$
+For $u\in\mathbb{C}$ and $y>0$, we define $d_{j}(u;y)=\frac{\vartheta^{{}^{\prime}}_{j}(u;iy)}{\vartheta_{j}(u;iy)}$, where $^{{}^{\prime}}$ denotes the derivative with respect to the first variable of elliptic theta functions.
 For $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
-(4.2) $\alpha_{2}=\sqrt{\vartheta_{3}(2r;iy)\vartheta_{3}(0;iy)},\quad\beta=\frac{\vartheta_{3}(iv;iy)}{\vartheta_{3}(0;iy)},\quad s_{2}=\vartheta_{3}(2r;iy)\vartheta_{1}(iv;iy),$
-$u_{2}=\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right),$
-$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),$
-$R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r;iy\right)}.$
-Noting $\vartheta_{j}(z;iy)^{*}=\vartheta_{j}(z^{*};iy)$ when $y>0$, we have
-(4.3) $\begin{split}&\alpha_{2}^{*}=\alpha_{2},\quad\beta^{*}=\beta,\quad s_{2}^{*}=-s_{2},\quad u_{2}^{*}=u_{2},\\ &\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}R_{j}^{*}=1.\end{split}$
+(4.6) $u_{1}=\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right),$
+$\Delta_{j}=d_{j}\left(-\frac{1}{2}iv+r;y\right)-d_{j}\left(\frac{1}{2}iv+r;y\right),\quad R_{j}=\frac{\vartheta_{j}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{j}\left(\frac{1}{2}iv+r\right)},$
+where $\vartheta_{j}(u;iy)$ is denoted by $\vartheta_{j}(u)$. Using Lemma4.1, we see
+(4.7) $\begin{split}u_{1}>0,\quad\Delta_{j}^{*}=-\Delta_{j},\quad R_{j}^{*}=R_{j}^{-1}.\end{split}$
+Note that from (4.7), we see $R_{1}R_{3}^{*}=R_{1}R_{3}^{-1}$ and is a complex number of magnitude $1$. Thus
+(4.8) $\begin{split}\frac{R_{1}R_{3}^{*}+(R_{1}R_{3}^{*})^{*}}{2}=\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2},\end{split}$
+is real and lies in [1,1][-1,1]. Also, from (4.7) and
+(4.9) $\Delta_{3}-\Delta_{1}=-\frac{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}^{{}^{\prime}}\left(0\right)}{u_{1}},$
+we see that $\Delta_{3}-\Delta_{1}\neq 0$.
 Theorem 4.2.
 Put $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$ and define
-(4.4) $F_{n}(z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
-$f_{n}(z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
-$g_{n}(z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
-$G_{n}(z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
-$H_{n}(z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
-$\widetilde{G}_{n}(z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right).$
-Then, they satisfy (3.1)-(3.6) to form $\tau$ functions.
-The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$.
-The torsion angle is given by
-(4.5) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(0)G_{n-1}(0)}{\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right).$
+(4.10) $F_{n}(z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv\right),$
+$f_{n}(z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{2}\left(\mu_{n}+r\right),$
+$g_{n}(z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-r\right),$
+$G_{n}(z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
+$H_{n}(z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r\right),$
+$\widetilde{G}_{n}(z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv\right),$
+where
+$\alpha_{1}=\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)},\quad\alpha_{2}=\vartheta_{3}\left(-\frac{1}{2}iv+r\right),\quad\alpha_{3}=\vartheta_{1}\left(\frac{1}{2}iv+r\right),$
+$\alpha_{4}=\frac{\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}{\vartheta_{3}(0)},\quad\alpha_{5}=\frac{u_{1}}{\vartheta_{3}(2r)\vartheta_{1}(iv)\sqrt{\vartheta_{3}(2r)\vartheta_{3}(0)}}.$
+Then, they satisfy (3.1)-(3.6) with
+(4.11) $\beta_{1}=\frac{\vartheta_{3}(iv)}{\vartheta_{3}(0)},$
+and form $\tau$ functions. The corresponding discrete curve $\gamma$ defined by (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
+(4.12) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(0)G_{n-1}(0)+G_{n}(0)G_{n-1}(0)^{*}}{2\widetilde{G}_{n}(0)\widetilde{G}_{n-1}(0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$
 In particular, this does not depend on $n$, and $\gamma$ has a constant torsion angle $\lambda_{n}=\lambda_{0}$. Moreover, the signed curvature angle is given by
-(4.6) $\begin{split}&\kappa_{n}=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right),\end{split}$
-where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$.
-A proof of Theorem 4.1 is given in Appendix B.
-4.2. Sufficient conditions for closedness
-The constant torsion-angle and unit segment-length curves given in Theorem 4.1 are not necessarily closed.
-We derive a sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle.
-We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}=\gamma_{n}$ for all $n$.
-Rewriting this condition using the $\tau$ functions, we obtain
-(4.7) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
-$n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
-It is difficult to give general solutions to (4.7), but we can utilise the pseudo-period of elliptic theta functions (A.12) to obtain particular solutions.
-For a fixed $k$, choose $m\in\mathbb{Z}\setminus{\{0\}}$ and set $v=my/k$.
-Then, from (A.12), we see
-(4.8) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
-$F_{n+k}=F_{n}\exp\left(k\frac{\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\exp\left[-\pi i\left\{2m\left(\mu_{n}-\frac{1}{2}iv\right)+m^{2}w\right\}\right],$
-where $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Using (4.8), (4.7) can be rewritten as
-(4.9) $\Lambda:=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
-$\Xi:=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
-Note that since
-(4.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi$
-holds, (4.9) can also be written as
-(4.11) $\Lambda=0,\quad\frac{\partial}{\partial r}\Lambda=0.$
-From the above discussion we obtain the following theorem.
-Theorem 4.2.
-For a fixed $k$, let $m\in\mathbb{Z}\setminus{\{0\}}$, $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$ satisfy $kv=my$ and (4.9).
-Then, the curve defined by Theorem 4.1 is closed and forms a $k$-Kaleidocycle with a torsion angle given by (4.5).
-Proof.
-The closure condition (4.7) holds for any $n\in\mathbb{Z}$ when the parameters $v,r,y$ satisfy $kv=my$ and (4.9).
-Thus, the conditions $kv=my$ and (4.9) give sufficient conditions for the curve given by Theorem 4.1 to be a closed curve.
-∎
-Hereafter, we assume that the parameters $v,r,y$ are taken to satisfy the sufficient conditions for closedness.
-Example 4.3.
-The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [8]) is obtained by $(v,r,y,m)\approx(0.640,0.250,1.279,3)$.
+(4.13) $\begin{split}\kappa_{n}&=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta_{1}F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right).\end{split}$
+This theorem follows as a corollary of Theorem 5.1 and will be proven in §5.1.
 Remark 4.3.
-By using (4.9) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$.
-From this, we also see that $m$ and $Lk$ have the same parity.
-Furthermore, under the transformation $\widetilde{B}_{n}\rightarrow(-1)^{n+1}\widetilde{B}_{n}$, $\widetilde{N}_{n}\rightarrow(-1)^{n+1}\widetilde{N}_{n}$, the signed curvature angle becomes $(-1)^{n+1}\kappa_{n}$ and the torsion angle becomes $\pi+\lambda$.
-This indicates that this transformation corresponds to the one described in Proposition 5.16.
-5. Construction of the motion of a Kaleidocycle
-By introducing the deformation parameter to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem 5.1.
-We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature.
-Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [16] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [19] are discussed and summarised in Table 2.
-5.1. Explicit formula of the discrete curve with deformation parameter
-We introduce the deformation parameter so that only the signed curvature angle changes with a constant torsion angle and while satisfying (3.1)-(3.6).
+Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discritised curvature (see, for example, [14]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
+[…]
+where $\tilde{K}=\pi\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}/2$ is the complete elliptic integral of the first kind and $\tilde{\xi}=\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}/\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}$ is the elliptic modulus.
+5. Deformation of a discrete curve with a constant torsion angle
+By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$).
+We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature (§5.2, §5.3, §5.4).
+Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [21] are discussed.
+This is also summarised in Table 1 in §6.
+5.1. Explicit formula of the curve with deformation parameter
+We introduce the deformation parameter $t$ so that only the signed curvature angles change while the torsion angle is kept constant.
 Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve.
@@
 Theorem 5.1.
-Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$.
-(5.1) $F_{n}(t,z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
-$f_{n}(t,z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
-$g_{n}(t,z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
-$G_{n}(t,z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
-$H_{n}(t,z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
-$\widetilde{G}_{n}(t,z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
+Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$:
+(5.1) $F_{n}(t,z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
+$f_{n}(t,z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
+$g_{n}(t,z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
+$G_{n}(t,z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
+$H_{n}(t,z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
+$\widetilde{G}_{n}(t,z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
 where $C,\Gamma\in\mathbb{R}$ are constants.
-Then, they satisfy (3.1)-(3.6) to form $\tau$ functions.
+Then, they satisfy (3.1)-(3.6) with $\beta_{1}$ given by (4.11) and form $\tau$ functions.
 For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$.
 The torsion angle is given by
-(5.2) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)}{\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right),$
+(5.2) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)+G_{n}(t,0)G_{n-1}^{*}(t,0)}{2\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$
 In particular, this depends neither on $n$ nor on $t$, and $\gamma$ has a constant torsion angle $\lambda_{n}(t)=\lambda_{0}(0)$. Moreover, the signed curvature angle is given by
-(5.3) $\begin{split}&\kappa_{n}\left(t\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
-where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$. If $r,v,y$ are chosen as in Theorem 4.2, $\gamma$ is closed for every $t$.
-A proof of Theorem 5.1 is given in Appendix B.
-Proposition 5.2.
+(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right).\end{split}$
+Proof.
+Putting
+[…]
+in (A.2) yields (3.1). Putting
+[…]
+in (A.2) yields (3.2). Putting
+[…]
+in (A.3) yields (3.3). Putting
+[…]
+in (A.4) yields
+(5.24) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
+Also, putting
+[…]
+in (A.4) yields
+(5.30) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
+Subtracting (5.30) from (5.24) and substituting
+(5.31) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
+in the equation, we obtain
+(5.32) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right),\end{split}$
+from which we see (3.4) holds. Putting
+[…]
+in (A.2) yields
+(5.38) $\begin{split}&\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
+Differentiating (5.38) with $h$ and then putting $h=0$, we obtain
+(5.39) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\\ &=-\Delta_{3}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\\ &-\Delta_{1}\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
+Using the equation (5.38) with $h=0$ and substituting (5.39), we obtain
+(5.40) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right.\\ &\left.+\Delta_{3}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\left(\Delta_{3}-\Delta_{1}\right)\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
+From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Theorem 4.2 follows from Theorem 5.1 by setting $t=0$. ∎
+Remark 5.2.
 The parameter $C$ contributes only to the parallel translation of the curve along the $z$-axis, whereas $\Gamma$ contributes only to the rotation of the curve around the $z$-axis.
-Proof.
-For $A\in\mathbb{C}$ and $B:\mathbb{C}\to\mathbb{C}$, we have
-(5.4) $\begin{split}&\left(\begin{array}[]{ccc}\cos\Gamma t&-\sin\Gamma t\\ \sin\Gamma t&\cos\Gamma t\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\frac{1}{2}\left(\begin{array}[]{ccc}e^{i\Gamma t}+e^{-i\Gamma t}&-\frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)\\ \frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)&e^{i\Gamma t}+e^{-i\Gamma t}\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\left(\begin{array}[]{ccc}e^{i\Gamma t}A+e^{-i\Gamma t}A^{*}\\ \frac{1}{i}\left(e^{i\Gamma t}A-e^{-i\Gamma t}A^{*}\right)\end{array}\right)\end{split}$
-and
-(5.5) $\begin{split}2\frac{\partial\log\left(e^{\frac{C}{2}tz}B\right)}{\partial z}=Ct+2\frac{\partial\log B}{\partial z}.\end{split}$
-The assertion follows from the definition (3.10). ∎
 In Theorem 5.11, we will make convenient choices of $C$ and $\Gamma$ to derive the deformation equation considered in the previous studies presented in §3.2.
-Table 1 summarises the relationship between various variables.
-The first column lists the parameters that determine the curve $\gamma$ through the variables listed in the first row.
-A circle in the table indicates that the variable shown in the first row is dependent on the parameter shown in the first column.
-Table 1. Dependency of variables on the parameters $v,r,y,k,m,t,C,\Gamma$.
-$\gamma$ $\kappa$ $\lambda$ $\Lambda$ $\Xi$ $u_{2}$ $\beta$ $s_{2}$ $R_{j}$ $\Delta_{j}$ $\alpha_{2}$ $\beta_{1,2}$ $\epsilon_{1,2}$ $\omega$ $\delta_{1,2,3}$
-$v$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
-$r$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
-$y$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
-$k$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
-$m$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
-$t$circ $\circ$
-$\Gamma$circ $\circ$
-$C$circ
-(3.10) (5.3) (5.2) (4.9) (4.9) (4.2) (4.2) (4.2) (4.2) (4.2) (4.2) (5.6) (5.6) (5.6) (5.6)
-The constructed motion has a periodic curvature. The next proposition follows from (A.12).
-Proposition 5.3.
-We have $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right)$. In other words, $2y$ is a period of the motion.
-Remark 5.4.
-There are certain relations between the curves constructed in Theorem 5.1 with different sets of parameters.
-The transformation of parameters $(v,m)\rightarrow(-v,-m)$ preserves the closure condition.
-We see that $\kappa_{n}\left(t;-v\right)=-\kappa_{n}\left(-t;v\right),\lambda(-v,r)=-\lambda(v,r)$.
-In what follows, We always choose $(v,m)$ or $(-v,-m)$ so that $\lambda\geq 0$.
 5.2. Motion of vertex positions
@@
 However, with the right choice of $C$ and $\Gamma$, curves that satisfy (3.24) are obtained as described in Theorem 5.11.
-We derive bilinear equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
-For $\Gamma\in\mathbb{R},y>0,r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, we define
-(5.6) $\omega=\frac{\vartheta_{1}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)}{\vartheta_{3}\left(2r;iy\right)\vartheta_{1}\left(iv;iy\right)},$
-$\delta_{1}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)^{3}},$
-$\delta_{2}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)^{2}}{\vartheta_{1}\left(iv;iy\right)^{2}\vartheta_{3}\left(2r;iy\right)\vartheta_{3}\left(0;iy\right)},$
-$\delta_{3}=\frac{-iu_{2}}{s_{2}\vartheta_{3}(0;iy)\vartheta_{1}(iv;iy)}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right.$
-$\left.+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}\right],$
-$\epsilon_{1}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0;iy)}{\vartheta_{3}(0;iy)},$
-$\epsilon_{2}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0;iy)^{2}}{\vartheta_{1}(iv;iy)^{2}},$
-$\beta_{1}=\frac{1}{2}\left(\delta_{3}+\delta_{3}^{*}\right)+\frac{1}{2}i\Gamma\left(\delta_{2}-\delta_{2}^{*}\right),$
-$\beta_{2}=\frac{1}{2}i\left(\delta_{3}-\delta_{3}^{*}\right)-\frac{1}{2}\Gamma\left(\delta_{2}+\delta_{2}^{*}\right).$
-Then, it follows that
-(5.7) $\omega^{*}=-\omega,\quad\delta_{1}^{*}=\delta_{1},\quad\epsilon_{1}^{*}=\epsilon_{1},\quad\epsilon_{2}^{*}=\epsilon_{2},\quad\beta_{1}^{*}=\beta_{1},\quad\beta_{2}^{*}=\beta_{2}.$
-A proof of the following proposition is given in Appendix C.
+We first derive equations satisfied by the $\tau$ functions given by Theorem 5.1 to calculate $\dot{\gamma_{n}}$.
+Lemma 5.3 ([12, §3]).
+Using (3.2) and (3.3), we see that the following equations hold.
+(5.41) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n},$
+(5.42) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
 Proposition 5.4.
-The $\tau$ functions (5.1) satisfy the following bilinear equations.
-(5.8) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}+\omega G_{n}G_{n}^{*},$
-(5.9) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\omega G_{n}G_{n}^{*},$
-(5.10) $\delta_{2}f_{n}^{*}g_{n-1}-\delta_{2}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
-(5.11) $\delta_{1}F_{n}F_{n}+\delta_{2}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
-(5.12) $\left(\beta-\delta_{1}\right)F_{n}F_{n}+\left(\omega-\delta_{2}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
-(5.13) $D_{t}H_{n}\cdot F_{n}=\left(\delta_{3}+i\Gamma\delta_{2}\right)f_{n}^{*}g_{n-1}+\left(\delta_{3}^{*}-i\Gamma\delta_{2}^{*}\right)f_{n-1}^{*}g_{n},$
-(5.14) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\epsilon_{1}+C\right)F_{n}F_{n}+\epsilon_{2}G_{n}G_{n}^{*}.$
-By using these equations, we obtain the evolution equation of the vertex positions of the curve.
+The $\tau$ functions (5.1) satisfy the following equations.
+(5.43) $f_{n}f_{n-1}^{*}+g_{n}^{*}g_{n-1}=\beta_{1}F_{n}F_{n}+\beta_{2}G_{n}G_{n}^{*},$
+(5.44) $g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}=f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}=\beta_{2}G_{n}G_{n}^{*},$
+(5.45) $\beta_{4}f_{n}^{*}g_{n-1}-\beta_{4}^{*}f_{n-1}^{*}g_{n}=H_{n}F_{n},$
+(5.46) $\beta_{3}F_{n}F_{n}+\beta_{4}G_{n}G_{n}^{*}=f_{n}f_{n-1}^{*},$
+(5.47) $\left(\beta_{1}-\beta_{3}\right)F_{n}F_{n}+\left(\beta_{2}-\beta_{4}\right)G_{n}G_{n}^{*}=g_{n}^{*}g_{n-1},$
+(5.48) $D_{t}H_{n}\cdot F_{n}=\left(\beta_{5}+i\Gamma\beta_{4}\right)f_{n}^{*}g_{n-1}+\left(\beta_{5}^{*}-i\Gamma\beta_{4}^{*}\right)f_{n-1}^{*}g_{n},$
+(5.49) $D_{t}D_{z}F_{n}\cdot F_{n}=\left(\beta_{6}+C\right)F_{n}F_{n}+\beta_{7}G_{n}G_{n}^{*}.$
+where
+(5.50) $\beta_{2}=\frac{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}{\vartheta_{3}\left(2r\right)\vartheta_{1}\left(iv\right)},$
+$\beta_{3}=\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)^{3}},\quad\beta_{4}=\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)^{2}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)^{2}}{\vartheta_{1}\left(iv\right)^{2}\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)},$
+$\beta_{5}=\frac{-iu_{1}}{\vartheta_{3}(0)\vartheta_{3}(2r)\vartheta_{1}(iv)^{2}}\left[\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}+\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}^{{}^{\prime}}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right],$
+$\beta_{6}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{3}^{{}^{\prime\prime}}(0)}{\vartheta_{3}(0)},\quad\beta_{7}=\frac{2i}{\Delta_{3}-\Delta_{1}}\frac{\vartheta_{1}^{{}^{\prime}}(0)^{2}}{\vartheta_{1}(iv)^{2}}.$
+Proof.
+Putting
+[…]
+in (A.8) yields (5.43). Also, (5.44) follows from (5.41) and (5.43). Putting
+[…]
+in (A.3) yields
+(5.61) $\begin{split}&\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(\mu_{n-1}-r+it\right)\\ &-\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(\mu_{n-1}-r+it\right),\end{split}$
+After differentiating (5.61) by $h$ and putting $h=0$, we obtain (5.48) by rearrangement. Also, rearranging the equation with $h=0$ in (5.61) yields (5.45). Putting
+[…]
+in (A.5) yields (5.46). Also, (5.47) follows from (5.43) and (5.46). Putting
+[…]
+in (A.2) yields
+(5.72) $\begin{split}&\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\\ &=-\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}+\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
+After differentiating (5.72) twice with $h$ and putting $h=0$, we obtain (5.49). ∎
+Coefficients $\beta_{2}$ and $\beta_{4}$ in Proposition 5.4 satisfy the following relation.
+Lemma 5.5.
+The following relation holds.
+(5.73) $\beta_{2}=\beta_{4}-\beta_{4}^{*}.$
+Proof.
+Putting
+[…]
+in (A.9) yields (5.73). ∎
+Furthermore, to simplify the calculation of $\dot{\gamma_{n}}$, we define $P$ and $Q$ as follows
+(5.79) $P_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
+(5.80) $Q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
+Then the following expressions hold
+(5.81) $D_{t}H_{n}\cdot F_{n}=\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*},$
+(5.82) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054ptQ_{n}^{*}+Q_{n}\\ \vskip 5.69054pt\frac{1}{i}\left(Q_{n}^{*}-Q_{n}\right)\\ -2\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\end{array}\right),$
+(5.83) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\frac{1}{i}\left(P_{n}-P_{n}^{*}\right)\\ \vskip 5.69054ptP_{n}+P_{n}^{*}\\ -\frac{2}{i}\beta_{2}|G_{n}|^{2}\end{array}\right),$
+where
+(5.84) $\eta_{1}=\frac{1}{2}\left(\beta_{5}+\beta_{5}^{*}\right)+\frac{1}{2}i\Gamma\left(\beta_{4}-\beta_{4}^{*}\right),\quad\eta_{2}=\frac{1}{2}i\left(\beta_{5}-\beta_{5}^{*}\right)-\frac{1}{2}\Gamma\left(\beta_{4}+\beta_{4}^{*}\right).$
+We have used (3.2) and (3.3) for (5.83). Then we have the following lemmas.
+Lemma 5.6.
+The following relations hold.
+(5.85) $2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}=2\beta_{4}\beta_{5}^{*}+2\beta_{4}^{*}\beta_{5}+\beta_{7}=0,$
+(5.86) $4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}=0.$
+Proof.
+By the definition of $\eta_{1}$ and $\eta_{2}$ and Lemma 5.5, we can see that the leftmost and middle part of (5.85) match. Putting
+[…]
+in (A.9) yields
+(5.92) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
+After differentiating (5.92) by $h$ and setting $h=0$, we obtain (5.85). From
+(5.93) $\begin{split}&4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}+8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)-4\eta_{2}^{2}\beta_{2}^{2}-\beta_{7}^{2}\\ =&\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}-\beta_{7}\right\},\end{split}$
+we obtain (5.86). ∎
+Lemma 5.7.
+The following relations hold.
+(5.94) $P_{n}P_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\beta_{2}^{2}|G_{n}|^{4},$
+(5.95) $\begin{split}P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
+(5.96) $\begin{split}&Q_{n}Q_{n}^{*}=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
+Proof.
+By definition of $P$, we see that
+(5.97) $\begin{split}P_{n}P_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right),\end{split}$
+holds. Substituting (5.42) and (5.44) into (5.97) yields (5.94). By definition of $P$ and $Q$, we see that
+[…]
+holds. On the other hand, from (5.41) and (5.44), the following holds.
+[…]
+Using (5.99), we obtain
+(5.100) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
+Substituting (5.44), (5.46) and (5.47) into (5.100) gives
+(5.101) $\begin{split}&P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}\\ &=\beta_{2}|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
+thus (5.95). Also from (5.42), (5.46) and (5.47), we see that
+(5.102) $\begin{split}&Q_{n}Q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\beta_{3}F_{n}^{2}+\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta_{1}-\beta_{3}\right)F_{n}^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}\right)\\ &=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4},\end{split}$
+holds. Thus we get (5.96). ∎
+By using Proposition 5.4, Lemma 5.5 Lemma 5.6 and Lemma 5.7, we obtain the evolution equation of the vertex positions of the curve.
 Theorem 5.8.
 The curve $\gamma$ constructed in Theorem 5.1 satisfies the following deformation equation.
-(5.15) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta\left[\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right).\end{split}$
-A proof of Theorem 5.6 is given in Appendix D.
-This is a deformation equation similar to the one presented in [16].
-Starting from (5.15), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
+(5.103) $\begin{split}\dot{\gamma}_{n}(t)=&2\beta_{1}\left[\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right]\\ &+\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right).\end{split}$
+Proof.
+By differentiating $\gamma$ with respect to $t$, we get
+(5.104) $\begin{split}\dot{\gamma}_{n}(t)&=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -\left(\beta_{6}+C\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
+On the other hand, from (5.82) and (5.83), we see
+(5.105) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}+2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{2}+2i\eta_{2}\beta_{2}|G_{n}|^{2}\end{array}\right)\end{split}$
+holds. From (5.85), (5.105) can be rewritten as follows
+(5.106) $\begin{split}&\eta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta_{1}F_{n}^{2}}\left(\begin{array}[]{c}\vskip 5.69054pt\displaystyle\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\\ \vskip 5.69054pt\displaystyle\frac{1}{i}\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ -2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)F_{n}^{2}-\beta_{7}|G_{n}|^{2}\end{array}\right).\end{split}$
+Thus (5.103) is obtained. ∎
+This is a deformation equation similar to the one presented in [18]. Starting from (5.103), the compatibility condition of the Frenet frame yields the semi-discrete mKdV equation.
 Proposition 5.9.
@@
     $\kappa$ satisfies the following semi-discrete mKdV equation:
-    (5.16) $\dot{\kappa}_{n}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
+    (5.107) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
 Furthermore, if the left-hand-side of the equation (5.136) does not depend on $n$, the following holds:
@@
 Proof.
-The proofs of (i) and (ii) are given by the same calculation as in Section 5 of [16]. We prove (iii). From the deformation equation (5.15) and the discrete Frenet-Serret formula (2.6), we see that
-(5.17) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta\beta_{1}\sin\lambda-2\beta\beta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
+The proofs of (i) and (ii) follow from the same calculation as in Section 5 of [18]. We prove (iii). From the deformation equation (5.103) and the discrete Frenet-Serret formula (2.5), we see that
+(5.108) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta_{1}\eta_{1}\sin\lambda-2\beta_{1}\eta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
 holds, which vanishes because of (5.136). Therefore $\dot{\lambda}=0$.
@@
 Proposition 5.9 (iii) does not hold in general without the assumption on (5.136).
-We can construct a motion of a Kaleidocycle satisfying (5.15) such that the left-hand side of (5.27) depends both on $n$ and $t$.
+Indeed, there is a motion of a Kaleidocycle satisfying (5.103) such that the left-hand side of (5.136) depends both on $n$ and $t$.
 By adjusting the global motion of $\gamma$ by $C$ and $\Gamma$, the deformation equation of $\gamma$ can be made to satisfy the equation considered in the previous studies presented in §3.2.
 Theorem 5.11.
-Set $C,\Gamma$ to be
-(5.18) $\begin{split}\Gamma_{+}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)+\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{+}=&-\epsilon_{1}-i\Gamma_{+}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega},\\ \rm or\\ \Gamma_{-}=&\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)-\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega},\\ C_{-}=&-\epsilon_{1}-i\Gamma_{-}\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)\\ &-\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)+\frac{\delta_{3}-\delta_{3}^{*}}{\omega}.\end{split}$
-Then, the curve constructed in Theorem 5.1 satisfies the following deformation equations.
-(5.19) $\begin{split}&\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2},\end{split}$
-where $\rho^{2}>0$ is given by
-(5.20) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2}$
+Set $C$ and $\Gamma$ to be
+(5.109) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0,\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.\end{split}$
+Then, the curve constructed in Theorem 5.1 deforms according to the equation (3.24). The magnitude of $\dot{\gamma}_{n}$ is a constant and is given by
+(5.110) $\rho^{2}=\rho_{+}^{2}:=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}\right)^{2},$
 when $(C,\Gamma)=(C_{+},\Gamma_{+})$ and
-(5.21) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2}$
+(5.111) $\rho^{2}=\rho_{-}^{2}:=\left(-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}\right)^{2},$
 when $(C,\Gamma)=(C_{-},\Gamma_{-})$.
-A proof of Theorem 5.9 is given in Appendix D.
+Proof.
+We determine the necessary and sufficient condition for $C$ and $\Gamma$ to satisfy
+(5.112) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
+(5.113) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
+where $\rho\in\mathbb{R}$ is a constant. From (5.83) and (5.104) we see
+(5.114) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(P_{n}-P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)\\ &+\left(P_{n}+P_{n}^{*}\right)\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)\\ &+2\beta_{2}|G_{n}|^{2}\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\}\\ &+2\beta_{2}|G_{n}|^{4}\left\{2\eta_{1}\left(\beta_{4}+\beta_{4}^{*}\right)+2i\beta_{2}\eta_{2}+\beta_{7}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)\right\},\end{split}$
+where we have used (5.85). Therefore, we obtain the following as a necessary and sufficient condition for (5.112) to hold.
+(5.115) $2\eta_{1}\left(2\beta_{3}-\beta_{1}\right)\beta_{2}+2i\eta_{2}+\beta_{2}\left(\beta_{6}+C\right)=0,\\$
+Using (5.84), we rewrite this in terms of $C$ and $\Gamma$ to obtain
+(5.116) $i\Gamma\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)+\beta_{6}+C+\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.$
+Next is the calculation of (5.113). The following holds
+(5.117) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}+\eta_{1}Q_{n}-i\eta_{2}P_{n}\right)^{2}\\ &-\left(\eta_{1}Q_{n}^{*}+i\eta_{2}P_{n}^{*}-\eta_{1}Q_{n}+i\eta_{2}P_{n}\right)^{2}\\ &+\left\{\left(\beta_{6}+C\right)F_{n}^{2}+\beta_{7}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\eta_{1}^{2}\left(\beta_{4}+\beta_{4}^{*}\right)^{2}-8i\eta_{1}\eta_{2}\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)+4\eta_{2}^{2}\beta_{2}^{2}+\beta_{7}^{2}\right)\\ =&F_{n}^{4}\left\{16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)\right\},\end{split}$
+where we have used (5.86). Therefore, we obtain the following as a necessary and sufficient condition for (5.113) to hold.
+(5.118) $4\eta_{1}^{2}\left(1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right)-8i\eta_{1}\eta_{2}\beta_{2}\left(2\beta_{3}-\beta_{1}\right)+4\eta_{2}^{2}+2\beta_{7}\left(\beta_{6}+C\right)=0.$
+By using (5.85) and (5.116), (5.118) can be written in terms of $\Gamma$ as follows.
+(5.119) $i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)\Gamma^{2}+2\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\Gamma+i\left(\beta_{4}{\beta_{5}^{*}}^{2}-\beta_{4}^{*}\beta_{5}^{2}\right)=0.$
+This is a quadratic equation in $\Gamma$ and has the following solutions.
+(5.120) $\Gamma_{\pm}=\frac{-\beta_{4}\beta_{4}^{*}\left(\beta_{5}+\beta_{5}^{*}\right)\pm\sqrt{\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}}}{i\beta_{4}\beta_{4}^{*}\left(\beta_{4}-\beta_{4}^{*}\right)}.$
+Since
+(5.121) $\begin{split}&\beta_{4}\beta_{4}^{*}\left(\beta_{4}\beta_{5}^{*}+\beta_{4}^{*}\beta_{5}\right)^{2}=\left(\beta_{4}\beta_{4}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\\ &\beta_{4}-\beta_{4}^{*}=\frac{\vartheta_{1}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\\ &\beta_{5}+\beta_{5}^{*}=-i\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)}{\vartheta_{3}(2r)\vartheta_{1}(iv)},\end{split}$
+holds, $\Gamma_{\pm}$ can be rewritten as
+(5.122) $\Gamma_{\pm}=\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)\pm\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)}.$
+Substituting this into (5.116) yields the expression for $C_{\pm}$ as well. From the above calculations, we see that when $C_{\pm}$ and $\Gamma_{\pm}$ are chosen as in (5.109), the following holds
+(5.123) $\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}.$
+We calculate this to compute $\rho^{2}$. From (5.84), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
+(5.124) $16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
+From (5.115) and (5.85), we see that when $(C,\Gamma)=(C_{+},\Gamma_{+})$, the following holds
+(5.125) $\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
+Adding the above equations yields
+(5.126) $\rho^{2}=16\eta_{1}^{2}\beta_{3}\left(\beta_{1}-\beta_{3}\right)+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2},$
+where we have used
+(5.127) $\vartheta_{3}(r)^{4}+\vartheta_{1}(r)^{4}=\vartheta_{3}(2r)\vartheta_{3}(0)^{3}.$
+In the same way, we see that when $(C,\Gamma)=(C_{-},\Gamma_{-})$,
+(5.128) $\rho^{2}=16\eta_{1}^{2}\left(\beta_{1}-\beta_{3}\right)\beta_{3}+\left(\beta_{6}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{1}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2},$
+holds. ∎
 Remark 5.12.
@@
 Therefore, its sweeping surface forms a semi-discrete analogue of the discrete K-surface (see figures in §7).
-Since the curvature and torsion angles uniquely determine a curve up to a rigid transformation, Proposition 5.3 shows that the motion constructed in Theorem 5.1 is periodic up to a global rigid transform.
-Here, we control the global orientation by choosing the parameters $C$ and $\Gamma$ appropriately so that the motion becomes periodic in terms of vertex positions.
-Proposition 5.11.
-Let $l\in\mathbb{Z}$. If we set
-(5.22) $C$ $=\frac{4\pi i}{y}\frac{1}{\Delta_{3}-\Delta_{1}},$
-(5.23) $\Gamma$ $=-\frac{\pi}{y}\left(4r+2l+1\right),$
-we have
-(5.24) ${\gamma_{n}}(t+y)={\gamma_{n}}(t),$
-for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$.
+5.3. Evolution of curvature
+The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$. To obtain the recurrence equation, we use the following Lemma.
+Lemma 5.13.
+The $\tau$ functions (5.1) satisfy the following equations.
+(5.129) $\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
+(5.130) $\beta_{1}^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta_{1}^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
 Proof.
-First, we prove the periodicity for the $Z$-component. Denote the $Z$-component of $\gamma_{n}$ by ${\gamma_{n}}_{Z}$. From the pseudo-period of elliptic theta functions (A.12), we see
-(5.25) $\begin{split}&{\gamma_{n}}_{Z}(t+y)-{\gamma_{n}}_{Z}(t)\\ &=-2\frac{\partial}{\partial z}\log\left\{\exp\left(\frac{C}{2}yz\right)\frac{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it+iy\right)}{\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it\right)}\right\}\\ &=-2\frac{\partial}{\partial z}\log\exp\left\{\frac{C}{2}yz-\pi i\left(2iv\left(n-\frac{1}{2}\right)+\frac{2z}{\Delta_{3}-\Delta_{1}}+2it\right)+\pi y\right\}\\ &=-2\left(\frac{C}{2}y-\frac{2\pi i}{\Delta_{3}-\Delta_{1}}\right)=0.\end{split}$
-For the $X$ and $Y$ components, from (3.10) we have to show $\frac{H_{n}(t+y)}{F_{n}(t+y)}\frac{F_{n}(t)}{H_{n}(t)}=1$.
-If we write $P_{n}=iv\left(n-\frac{1}{2}\right)+\frac{z}{\Delta_{3}-\Delta_{1}}+it$, the left-hand side can be computed using (A.12) as
-(5.26) $\begin{split}&\exp\left(i\Gamma y\right)\frac{\vartheta_{4}\left(P_{n}-2r+iy\right)\vartheta_{2}\left(P_{n}\right)}{\vartheta_{2}\left(P_{n}+iy\right)\vartheta_{4}\left(P_{n}-2r\right)}\\ &=-\exp\left(i\Gamma y\right)\frac{\exp\left\{-\pi i\left(2P_{n}-4r+iy\right)\right\}}{\exp\left\{-\pi i\left(2P_{n}+iy\right)\right\}}\\ &=-\exp\left\{i\left(\Gamma y+4\pi r\right)\right\}=1.\end{split}$
-5.3. Evolution of curvature
-The curve constructed in Theorem 5.1 reveals an interesting recurrence equation for $\kappa$.
+Using (3.1)-(3.3), we see that (5.129) hold. Putting
+[…]
+in (A.10), we obtain (5.130). ∎
 Proposition 5.14.
@@
 The right-hand side of this equation depends neither on $n$ nor on $t$. Therefore, the quantity on the left-hand side can be considered a conserved quantity of the signed curvature angle.
-A proof of Proposition 5.12 is given in Appendix E. The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.7.
+Proof.
+Using (5.129) and (5.130), we see that (5.136) holds. ∎
+The evolution of the signed curvature angle satisfies the semi-discrete mKdV equation in Proposition 5.9. To obtain the semi-discrete mKdV equation, we use the following Lemma.
+Lemma 5.15.
+The $\tau$ functions (5.1) satisfy the following equation.
+(5.137) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\beta_{8}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right),$
+where
+(5.138) $\beta_{8}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv;y)-d_{3}(iv;y)\right).$
+Proof.
+After differentiating the equation (A.3) with
+[…]
+by $h$ and putting $h=0$, we obtain
+(5.144) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(\mu_{n-1}-\frac{1}{2}iv+it\right)\right\}.\end{split}$
+After rearranging (5.144), we obtain (5.137). ∎
 Theorem 5.16.
 The signed curvature angle (5.3) satisfies
-(5.28) $\dot{\kappa}_{n}=2\beta^{2}\eta_{3}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right),$
-where
-(5.29) $\eta_{3}=i\frac{\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}{\vartheta_{1}\left(2iv\right)\vartheta_{3}\left(0\right)}\left(d_{1}(iv)-d_{3}(iv)\right).$
-Combined with Proposition 5.7, (5.6), and the fact that $e^{i\lambda}\delta_{2}\in\mathbb{R}$, we see
-(5.30) $2\beta^{2}\eta_{3}=\left(2\beta\beta_{1}\cos\lambda+2\beta\beta_{2}\sin\lambda\right).$
-A proof of Theorem 5.13 is given in Appendix F.
+(5.145) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
+Proof.
+Differentiating the signed curvature angle yields the following.
+(5.146) $\dot{\kappa}_{n}=2\frac{d}{dt}\arctan\left(\frac{\widetilde{G}_{n}}{\beta_{1}F_{n}}\right)=2\beta_{1}\frac{D_{t}\widetilde{G}_{n}\cdot F_{n}}{\widetilde{G}_{n}^{2}+\beta_{1}^{2}F_{n}^{2}}.$
+Using (5.129) and (5.137), we see that the signed curvature angle satisfies
+(5.147) $\dot{\kappa}_{n}=2\beta_{1}^{2}\beta_{8}\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
+Since
+(5.148) $\beta_{1}\beta_{8}=\eta_{1}\cos\lambda+\eta_{2}\sin\lambda,$
+holds, this is nothing but the semi-discrete mKdV equation (5.145). ∎
 5.4. Potential function of curvature
@@
 Theorem 5.17.
-For fixed $v,y$, define the function $\Theta$ as follows.
+For fixed $v$ and $y$, define the function $\Theta$ as follows.
 (5.149) $\Theta_{n}\left(t\right)=\Theta_{n}\left(t;v,y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}.$
@@
 (5.150) $\kappa_{n}=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}.$
-A proof of Theorem 5.14 is given in Appendix G. The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation.
+Proof.
+From (A.17), (A.18) and (A.19), we see that
+(5.151) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2},\end{split}$
+holds. ∎
+The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation, establishing an explicit relation between our construction and the previous work [21] reviewed in §3.2.
+Since the semi-discrete sine-Gordon equation can be solved by almost the same calculation, we will solve only the semi-discrete potential mKdV equation here.
+We construct a solution to the equation using the theory of $\tau$ functions [17].
+For this purpose, we first define the following real-valued functions for fixed $v$ and $y$.
+(5.152) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
+Then the potential function can be written as follows
+(5.153) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
+where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$.
+$\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy.
+Then we have the following lemma.
+Lemma 5.18.
+The functions $A$ and $B$ defined by (5.152) satisfy the following equations.
+(5.154) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
+(5.155) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
+where
+(5.156) $J_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad J_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
+(5.157) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+J_{1}(v,y)+J_{2}(v,y),$
+(5.158) $\beta_{9}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-J_{1}(v,y)+J_{2}(v,y)}{L_{1}(v,y)}.$
+Proof.
+In this proof, all elliptic theta functions have $w=2i/y$ and we denote $\vartheta_{j}(v;2i/y)$ by $\vartheta_{j}(v)$. After differentiating the equation in (A.6) with
+[…]
+by $h$ and putting $h=0$, we obtain
+(5.164) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
+Also, after differentiating the equation in (A.7) with
+[…]
+by $h$ and putting $h=0$, we obtain
+(5.170) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
+Also, after differentiating the equation in (A.7) with
+[…]
+by $h$ and putting $h=0$, we obtain
+(5.176) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
+Also, after differentiating the equation in (A.7) with
+[…]
+by $h$ and putting $h=0$, we obtain
+(5.182) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
+Adding (5.164) and (5.170) yields (5.154), and subtracting (5.176) from (5.182) yields (5.155). ∎
+Then we have the following theorem.
 Theorem 5.19.
 Let $y>0$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$. The potential function $\Theta$ defined by (5.149) simultaneously satisfies the semi-discrete potential mKdV and sine-Gordon equations
+(5.183) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\left(1+\cos\lambda\right)\rho_{+}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
+(5.184) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\left(1-\cos\lambda\right)\rho_{-}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
+Proof.
+We first show that the potential function satisfies
 (5.185) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\frac{L_{1}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
-(5.34) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\frac{L_{2}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right),$
-where $Q_{1},Q_{2},L_{1},L_{2}\in\mathbb{R}$ are defined by
-(5.35) $Q_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad Q_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
-(5.36) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)+Q_{2}(v,y),$
-(5.37) $L_{2}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)-Q_{2}(v,y).$
-A proof of Theorem 5.15 is given in Appendix H.
-The following shows two equations are related by a parameter shift.
-Proposition 5.16.
-By the transformation of parameters $v\rightarrow v+y$ and $r\rightarrow\frac{1}{2}-r$, the signed curvature angle $\kappa$, the torsion angles $\lambda$, the potential functions $\Theta$, the coefficients of the semi-discrete mKdV equation $L_{1}/y$ and of the semi-discrete sine-Gordon equation $L_{2}/y$, the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ are transformed as
-(5.38) $\kappa_{n}\left(t\right)$ $\rightarrow(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2}\right),\quad\lambda\rightarrow\pi+\lambda,$
-(5.39) $\Theta_{n}\left(t\right)$ $\rightarrow(-1)^{n}\Theta_{n}\left(t-\frac{y}{2}\right),$
-(5.40) $\frac{L_{1}}{y}$ $\rightarrow\frac{L_{2}}{y},\quad\frac{L_{2}}{y}\rightarrow\frac{L_{1}}{y},$
-(5.41) $\widetilde{N}_{n}(t)$ $\rightarrow(-1)^{n+1}\widetilde{N}_{n}\left(t-\frac{y}{2}\right),\quad\widetilde{B}_{n}(t)\rightarrow(-1)^{n+1}\widetilde{B}_{n}\left(t-\frac{y}{2}\right).$
-A proof of Proposition 5.16 is given in Appendix I.
-The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{n}\Theta_{n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.33) with one of the semi-discrete sine-Gordon equation (5.34).
-Proposition 5.16 shows that this equation-level transformation is also achieved at the solution level by transforming the parameters.
+using Lemma 5.18. Then we show that
+(5.186) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
+Suppose that a solution to the equation (5.185) is expressed as (5.153). Then both sides of the equation can be rewritten as follows.
+(5.187) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
+and
+(5.188) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right).\end{split}$
+Then it can be seen that a sufficient condition for (5.153) to be a solution of (5.185) is that
+(5.189) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\beta_{9}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
+is satisfied. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
+(5.190) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
+The real part of this equation yields (5.154), and the imaginary part yields (5.155).
+Thus, Lemma 5.18 shows that the potential function satisfies the equation (5.185).
+We then calculate $L_{1}(v,y)/y$.
+From (A.15), (A.16) and (A.20), we see
+(5.191) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)-\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
+On the other hand, we see
+(5.192) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{1}}.\end{split}$
+We can see that (5.186) holds by rewriting (5.191) and (5.192) using the following.
+(5.193) $\begin{split}&\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{4}\left(iv\right)\vartheta_{2}\left(0\right)+\vartheta_{2}\left(iv\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\}\\ &=2\vartheta_{3}\left(r\right)^{2}\vartheta_{1}\left(r\right)^{2}.\end{split}$
 Remark 5.20.
 Rewriting (5.145) in terms of the potential function $\Theta$, we obtain
-(5.42) $\frac{d}{dt}\Theta_{n}=4\beta^{2}\eta_{3}\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
+(5.194) $\frac{d}{dt}\Theta_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
 which is also called the semi-discrete potential mKdV equation [16, 15, 7].
@@
 It does not usually happen that these equations are simultaneously solved.
-For example, it can be checked that the 2 and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [6, 7] are not solutions of the semi-discrete potential mKdV equation.
-The following describes an explicit relation between our construction and the previous work [19] reviewed in §3.2.
-Proposition 5.18.
-When $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), we have
-(5.43) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
-Similarly, when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), we have
-(5.44) $\frac{L_{2}(v,y)}{y}\frac{1}{1-\cos\lambda}=-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}=\rho_{-}.$
-A proof of Proposition 5.18 is given in Appendix J.
+For example, it can be checked that the 2- and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [7, 8] are not solutions of the semi-discrete potential mKdV equation.
 Note the difference in approaches of [21] and this work; [21] starts with a geometric property observed in the motion of Kaleidocycles, and derive evolution equations, whereas we aim to construct a particular motion of Kaleidocycles and investigate the equations satisfied by them.
 They turn out to arrive at the same place; the latter gives an explicit solution to the equations derived in the former.
+6. Existence of Kaleidocycles
+The constant torsion-angle and unit segment-length curves given in Theorem 5.1 are not necessarily closed.
+In this section, we derive a necessary and sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle.
+Furthermore, we prove that there exist parameters that satisfy the closure condition when $k\geq 6$.
+The overall approach is as follows: For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle has to have period $k$.
+From this, we see that $m:=kv/y$ is an integer (§6.1).
+Then, rewriting the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ assuming $m$ is an integer, we obtain two equations for v,r,yv,r,y6.2).
+By rearranging one of the equations, we verify that $r=r(y)$ is uniquely determined when $y$ is in a certain range (§6.3).
+Substituting this $r(y)$ into the other equation, we find that if $m\geq 3$, then there exists $y$ satisfying it (§6.4).
+By considering two transformations of $m$ and $n$ that preserve the shape of a curve, we see that it is sufficient to consider $0<m\leq k/2$.
+Finally, we show that under this range of $m$ and $k\geq 6$, there always exists a set of parameters with which the curve Theorem 5.1 is closed, and hence, gives a $1$-dimensional periodic motion of a Kaleidocycle (§6.5).
+6.1. Period of the curvature angle
+We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$.
+In this case, the absolute value of the curvature angles also has a period $k$.
+Hence, we obtain the necessary condition for the curve to be closed that
+(6.1) $\begin{split}\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}=0,\end{split}$
+holds for all $n$ and $t$. Note that there is no Kaleidocycle with period $1$, hence $k\geq 2$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
+(6.2) $\begin{split}&\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}\\ &=\frac{\vartheta_{1}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}}{\vartheta_{3}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{4}\left(0;\frac{i}{y}\right)^{2}}\left\{\mathrm{cn}^{2}\left[2\tilde{K}\left(\tilde{\mu}_{n}(t)+\frac{v}{y}k\right)\right]-\mathrm{cn}^{2}\left[2\tilde{K}\tilde{\mu}_{n}(t)\right]\right\}=0,\end{split}$
+where we note $\tilde{\mu}_{n}(t)=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}$.
+Since $\tilde{\mu}_{n}(t)$ can be regarded as a linear function of $t$ when $v,y$ and $n$ are fixed, $\tilde{\mu}_{n}(t)$ attains all reals as $t$ varies.
+Thus, we only need to check that (6.2) holds for $n=0$.
+Since $\mathrm{cn}^{2}$ has a real period $2\tilde{K}$, $m=kv/y$ must be an integer.
+Hereafter, let $v$ satisfy
+(6.3) $v=\frac{m}{k}y.$
+6.2. Closure conditions
+Here we obtain the necessary and sufficient conditions satisfied by $v,r,y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
+(6.4) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
+$n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
+We rewrite the left-hand side further with (A.12) under the condition that $m$ is an integer. Then we obtain
+(6.5) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
+$F_{n+k}=F_{n}\exp\left(k\frac{\Delta_{3}}{\Delta_{3}-\Delta_{1}}z\right)\exp\left[-\pi i\left\{2m\left(\mu_{n}-\frac{1}{2}iv\right)+m^{2}w\right\}\right],$
+where $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Using (6.5), (6.4) can be rewritten as
+(6.6) $\Lambda(v,r,y):=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
+(6.7) $\Xi(v,r,y):=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
+Note that from the definitions of $R_{j}$ and $\Delta_{j}$, we see that
+(6.8) $\frac{\partial}{\partial r}\log R_{j}=\Delta_{j}\quad(j=1,2,3,4),$
+holds. From this, we have
+(6.9) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi.$
+Since $\Delta_{3}-\Delta_{1}\neq 0$ holds, (6.7) can be written as
+(6.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=0.$
+On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v,r,y$ satisfy $kv=my$, (6.6) and (6.10).
+Thus, they give necessary and sufficient condition for the curve given by Theorem 5.1 to be a closed curve for every $t\in\mathbb{R}$.
+6.3. An explicit solution to the condition (6.10)
+Here we solve (6.10) with respect to $r$.
+From the period of elliptic theta function, we see the period and symmetry of $\Lambda(v,r,y)$ with respect to $r$ as follows: $\Lambda(v,r+1,y)=\Lambda(v,r,y)$, $\Lambda(v,-r,y)=\Lambda(v,r,y)^{*}$.
+Therefore, when searching for $(v,r,y)$ such that $\Lambda(v,r,y)=0$, we do not lose generality by assuming $0\leq r\leq 1/2$.
+We exclude the case $r=0,1/2$ because $\gamma$ forms a plane curve.
+Hereafter, we assume $0<r<1/2$.
+Since
+[…]
+holds, we have
+(6.12) $\begin{split}-&\frac{u_{1}}{k}\frac{\partial}{\partial r}\log(\Lambda+1)=-u_{1}\left(\Delta_{3}+\Delta_{1}\right)+4\pi i\frac{v}{y}u_{1}\\ &=\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\left(\vartheta_{4}\left(2r\right)\vartheta_{2}^{{}^{\prime}}\left(iv\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}^{{}^{\prime}}\left(iv\right)\right)\\ &+2\pi i\frac{v}{y}\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\left(\vartheta_{4}\left(2r\right)\vartheta_{2}\left(iv\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(iv\right)\right)=0.\end{split}$
+Arranging this we obtain
+(6.13) $\frac{\vartheta_{4}\left(0\right)\vartheta_{2}\left(2r\right)}{\vartheta_{2}\left(0\right)\vartheta_{4}\left(2r\right)}=\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{v}{y}\vartheta_{2}\left(iv\right)+\vartheta_{2}^{{}^{\prime}}\left(iv\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{v}{y}\vartheta_{4}\left(iv\right)+\vartheta_{4}^{{}^{\prime}}\left(iv\right)\right)}.$
+The right-hand side of this equation does not contain $r$. The left-hand side is nothing but Jacobi $\mathrm{cn}$-function.
+(6.14) $\mathrm{cn}\left[4Kr,\xi\right]=\frac{\vartheta_{4}\left(0\right)\vartheta_{2}\left(2r\right)}{\vartheta_{2}\left(0\right)\vartheta_{4}\left(2r\right)},$
+where $\xi=\vartheta_{2}\left(0\right)^{2}/\vartheta_{3}\left(0\right)^{2}$ is an elliptic modulus and $K=\pi\vartheta_{3}\left(0\right)^{2}/2$ is a complete elliptic integral of the first kind.
+We define the following function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
+(6.15) $\begin{split}\varphi\left(\frac{m}{k},y\right):=\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}&=\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}.\end{split}$
+When $1<\varphi\left(m/k,y\right)<1$ is satisfied, $r=r(y)$ is uniquely determined because the inverse function of $\mathrm{cn}$ exists.
+6.3.1. Existence of a inverse function
+Here we show that for a certain range of $y$, $1<\varphi\left(m/k,y\right)\leq 0$ holds when $0<m\leq k/2$.
+Then we find $r=r(y)$ explicitly using elliptic integrals.
+First, from the zeros and evenness of elliptic theta function we see the following: For any $y>0$, $\varphi(1/2,y)=0$ and $\varphi(m/k,y)\neq 0\quad(0<m/k<1/2)$.
+Therefore, when $m/k=1/2$, $r=1/4$.
+Next, consider the limit of $\varphi(m/k,y)$.
+From the definition of elliptic theta function, when $u$ is fixed and $y\rightarrow+0$, we see that
+(6.16) $\vartheta_{2}\left(u;\frac{i}{y}\right)$ $=2q^{\frac{1}{4}}\cos\pi u+\mathcal{O}(q^{\frac{9}{4}}),$
+$\vartheta_{4}\left(u;\frac{i}{y}\right)$ $=1+\mathcal{O}(q),$
+$\vartheta_{2}^{{}^{\prime}}\left(u;\frac{i}{y}\right)$ $=-2\pi q^{\frac{1}{4}}\sin\pi u+\mathcal{O}(q^{\frac{9}{4}}),$
+$\vartheta_{4}^{{}^{\prime}}\left(u;\frac{i}{y}\right)$ $=4\pi q\sin 2\pi u+\mathcal{O}(q^{4}),$
+where we denote $q=e^{-\frac{\pi}{y}}$. From this we have
+(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<1/2).$
+On the other hand, the following hold when $y\rightarrow\infty$
+(6.18) $p^{\left(\frac{m}{k}\right)^{2}}\vartheta_{2}\left(i\frac{m}{k}y;iy\right)$ $=p^{\left(\frac{1}{2}-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(\frac{1}{2}+\frac{m}{k}\right)^{2}}),$
+$p^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}\left(i\frac{m}{k}y;iy\right)$ $=p^{\left(\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1-\frac{m}{k}\right)^{2}}),$
+$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=\pi p^{\left(\frac{1}{2}-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(\frac{1}{2}+\frac{m}{k}\right)^{2}}),$
+$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<1/2),$
+where we denote $p=e^{-\pi y}$. From this we have
+(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<1/2).$
+Hence, when $m$ is fixed within $0<m<k/2$, there exists $y$ such that $\varphi\left(m/k,y\right)=-1$.
+We denote the smallest one among them by $y_{0}(m/k)$.
+Summarizing the above, when $m$ is fixed within $0<m<k/2$ and $y$ moves in the range $0<y<y_{0}(m/k)$, $r=r(y)$ satisfying (6.10) is uniquely determined within the range of $0<r<1/2$.
+By taking the inverse function of $\mathrm{cn}$ in (6.13), we obtain
+(6.20) $\begin{split}&r\left(y;\frac{m}{k}\right)=\frac{1}{4K}\int^{1}_{\varphi(m/k,y)}\frac{dx}{\sqrt{(1-x^{2})(1-\xi^{2}+\xi^{2}x^{2})}}.\end{split}$
+This also holds when $m/k=1/2$, hence we see that $1/4\leq r<1/2$. If we assume 1/2<r<0-1/2<r<0, the same calculation yields
+(6.21) $\begin{split}r\left(y;\frac{m}{k}\right)=-\frac{1}{4K}\int^{1}_{\varphi(m/k,y)}\frac{dx}{\sqrt{(1-x^{2})(1-\xi^{2}+\xi^{2}x^{2})}}.\end{split}$
+From this, we see $1/2<r\leq-1/4$ holds when $0<m\leq k/2$. The sign of $r$ is chosen so that $\sin\lambda=\langle\widetilde{B}_{n},\widetilde{N}_{n-1}\rangle>0$ holds.
+6.4. Existence of a solution to the condition $\Lambda=0$
+By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$. Note that $\Lambda+1$ can also be written as
+(6.22) $\begin{split}\Lambda+1&=\exp\left\{\pi im\left(4r+1\right)\right\}\left[\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}\right]^{k}\\ &=\exp{\left(\pi im\right)}\left[\frac{\vartheta_{1}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}{\vartheta_{1}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}\right]^{k},\end{split}$
+by transforming the parameter $y\rightarrow 1/y$. Since $\Lambda+1$ is a complex number of magnitude $1$, one may denote
+(6.23) $\Lambda+1=\exp{\left(i\Psi\right)},$
+using a real-valued function $\Psi=\Psi(v,r,y)$. When
+(6.24) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|,$
+is greater than $2\pi$, then there exists a solution to $\Lambda(v,r(y),y)=0$ in the range $0<y<y_{0}(m/k)$.
+Here, we compute the above quantities and show that such a $y$ exists if $m\geq 3$.
+From a representation by an infinite product of elliptic theta functions,
+(6.25) $\vartheta_{1}\left(u;\frac{i}{y}\right)$ $=2q^{\frac{1}{4}}\sin(\pi u)\prod_{n=1}^{\infty}\left(1-q^{2n}\right)\left(1-q^{2n}e^{2\pi iu}\right)\left(1-q^{2n}e^{-2\pi iu}\right),$
+$\vartheta_{3}\left(u;\frac{i}{y}\right)$ $=\prod_{n=1}^{\infty}\left(1-q^{2n}\right)\left(1+q^{2n-1}e^{2\pi iu}\right)\left(1+q^{2n-1}e^{-2\pi iu}\right),$
+we obtain
+(6.26) $\begin{split}&\frac{\vartheta_{1}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{1}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\frac{\sin\left(\pi\frac{r}{iy}+\pi\frac{m}{2k}\right)}{\sin\left(\pi\frac{r}{iy}-\pi\frac{m}{2k}\right)}\exp{\left(-2i\sum_{n=1}^{\infty}\frac{1}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{2n(1-r)}-q^{2n(1+r)}\right)\right)},\\ &\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\exp{\left(-2i\sum_{n=1}^{\infty}\frac{(-1)^{n}}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{n(1-2r)}-q^{n(1+2r)}\right)\right)}.\end{split}$
+Furthermore, from
+[…]
+we obtain
+(6.28) $\lim_{y\to+0}\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}=1.$
+Similarly, we obtain
+(6.29) $\lim_{y\to+0}\frac{\vartheta_{1}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{1}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}=\exp{\left(\pi i\frac{m}{k}\right)}.$
+From the above, we have
+(6.30) $\lim_{y\to+0}\Psi(v,r(y),y)=2\pi m.$
+From the pseudoperiod of elliptic theta function, we have
+(6.31) $\Psi(v,r(y_{0}),y_{0})=3\pi m.$
+Summarizing the above discussion, we have
+(6.32) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|=\pi m.$
+Therefore, when $m\geq 3$, there is always a solution in $\Lambda(v,r(y),y)=0$.
+6.5. Transformations of a parameter and a variable
+We show that the set of our explicit solutions admits two types of self-bijections that transform
+  • $m\rightarrow m+k$, $n\rightarrow k-n$,
+  • $m\rightarrow-m$, $n\rightarrow n$.
+A finite number of combinations of these transformations can make any $m\in\mathbb{Z}\setminus(k\mathbb{Z})$ fall in the interval $0<m\leq k/2$.
+Combining with the previous subsection, this proves the existence of Kaleidocycles for $k\geq 6$.
+In the following, we will construct the self-bijections and prove that the Kaleidocycles obtained by the self-bijections are essentially identical to the Kaleidocycles before the transformation.
+A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarised in Table 1.
+6.5.1. Transformation $m\rightarrow m+k$, $n\rightarrow k-n$
+Consider transformation $v\rightarrow v+y$.
+Then, $m$ changes to $k+m$.
+The period of elliptic theta function $\varphi(m/k+1,y)=-\varphi(m/k,y)$ and (6.21) implies the change $r\rightarrow-r-1/2$.
+From the period of elliptic theta function, we see that $\Lambda(v+y,-r-1/2,y)=\Lambda(v,r,y)^{*}$ holds.
+Thus, this transformation preserves the closure conditions.
+From the original binormal vectors $\widetilde{B}_{n}=\widetilde{B}_{n}(t;v,r,y)$, the new binormal vectors are computed as
+(6.33) $\begin{split}(-1)^{k-n}\widetilde{B}_{k-n}\left(t-y/2;v,r,y\right).\end{split}$
+That is, the new Kaleidocycle is obtained by shifting the time parameter, reversing the orientation of the central curve, and flipping every other hinge, which is equivalently encoded by the torsion angle update $\lambda\mapsto\pi-\lambda$.
+6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
+Consider the transformation $v\rightarrow-v$ and $r\rightarrow-r$.
+Then, $m$ changes to m-m.
+From the period of elliptic theta function, we see that $\Lambda(-v,-r,y)=\Lambda(v,r,y)$ holds.
+Thus, this transformation preserves the closure conditions.
+From the original binormal vectors $\widetilde{B}_{n}=\widetilde{B}_{n}(t;v,r,y)$, the new binormal vectors are computed as
+(6.34) $\begin{split}-\widetilde{B}_{n}\left(-t;v,r,y\right).\end{split}$
+That is, the new Kaleidocycle is obtained by reversing the time parameter and flipping all hinges.
+Summarising the results of this section, we have the following theorem.
+Theorem 6.1.
+For a fixed $k\geq 6$, let $0<m\leq k/2$.
+Then there exist $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $y>0$ satisfying $kv=my$, (6.6) and (6.10).
+Then, the curve defined by Theorem 5.1 is closed for every $t\in\mathbb{R}$ and forms a $k$-Kaleidocycle with the torsion angle given by (5.2).
+Remark 6.2.
+When a curve is deformed according to (3.24), a combination of the two transformations described above exchanges the deformation by the semi-discrete potential mKdV equation and the deformation by the semi-discrete sine-Gordon equation.
+Consider the transformation $v\rightarrow-v-y$, $r\rightarrow r+1/2$ and $n\rightarrow k-n$.
+Then the new $\Gamma$ and $C$ are computed as
+(6.35) $\begin{split}&\Gamma_{\pm}\left(r+1/2,y\right)=\Gamma_{\mp}\left(r,y\right),\\ &C_{\pm}\left(-v-y,r+1/2,y\right)=C_{\mp}\left(v,r,y\right).\end{split}$
+The new potential function is computed as
+(6.36) $(-1)^{k-n+1}\Theta_{k-n}\left(-t+y/2;v,y\right).$
+The coefficients of the semi-discrete potential mKdV equation and of the semi-discrete sine-Gordon equation are transformed as
+(6.37) $\left(1+\cos\lambda\right)\rho_{+}\rightarrow-\left(1-\cos\lambda\right)\rho_{-},\quad\left(1-\cos\lambda\right)\rho_{-}\rightarrow-\left(1+\cos\lambda\right)\rho_{+}.$
+The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{k-n+1}\Theta_{k-n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.183) with one of the semi-discrete sine-Gordon equation (5.184).
+We see that this equation-level transformation is also achieved at the solution level by transforming parameters and a variable.
+This is also consistent with the result of the transformation of $\Gamma$ and $C$.
 Table 1 summarises the relationships among the three deformation equations.
 Table 1. Relation between the three deformation equations
-Values of
-$C,\Gamma$
-Deformation equation of the curve Equation for potential function
+Deformation equation of the curve Equation for the potential function
 $C,\Gamma\in\mathbb{R}$
-$\dot{\gamma}_{n}=\footnotesize\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right)\normalsize$
-$+2\beta\left(\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
-$\dot{\Theta}_{n}=\left(4\beta\beta_{1}\cos\lambda+4\beta\beta_{2}\sin\lambda\right)$
-×$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
-$\downarrow$ Specialise the deformation equation by choosing rigid transformation parameters $C,\Gamma$.
+$\dot{\gamma}_{n}=\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right)$
+$2\beta_{1}\left(\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
+$\dot{\Theta}_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)$
+$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
+$\downarrow$ Specialise the deformation equation by choosing
+rigid transformation parameters $C$ and $\Gamma$ as in Theorem 5.11.
 $C_{+},\Gamma_{+}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{+}$
 $\frac{\dot{\Theta}_{n+1}+\dot{\Theta}_{n}}{2}=\left(1+\cos\lambda\right)\rho_{+}$
-×$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
-$\updownarrow$ By replacing the parameters of the solution with $v\rightarrow v+y,r\rightarrow 1/2-r$,
-these two deformation equations transform to each other.
+$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
+$\updownarrow$ As in Remark 6.2, these equations exchange with
+the transformation of parameters $v,r$ and a variable $n$ of the solution.
 $C_{-},\Gamma_{-}$ $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle=\rho^{2}_{-}$
 $\frac{\dot{\Theta}_{n+1}-\dot{\Theta}_{n}}{2}=\left(1-\cos\lambda\right)\rho_{-}$
-×$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
+$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
+Example 6.3.
+The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [9]) is obtained by $m=3$ and $(v,r,y)\approx(0.6400,-0.2500,1.279)$.
+Remark 6.4.
+By using (6.6) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. That is, the parity of $m$ corresponds to the orientability of the Kaleidocycle.
 7. Examples
 In this section, we present examples of Kaleidocycles constructed by Theorem 6.1 and the semi-discrete K-surfaces obtained as their motion trajectory as explained in Remark 5.12.
-Numerical experiments using our computer codes [17] suggest that the $8$-, $9$- and $15$-Kaleidocycles in Figs.
-2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, $9$-, and $15$-Kaleidocycles, respectively [18].
-In other words, they are the Möbius Kaleidocycles [20].
+Numerical experiments using our computer codes [19] suggest that the $8$-, 99- and $15$-Kaleidocycles in Figs.
+2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, 99-, and $15$-Kaleidocycles, respectively [20].
+In other words, they are the Möbius Kaleidocycles [22] (see Remark 2.1).
 For Fig.
@@
 We numerically find that the latter attains the local minima of the torsion angle.
-We present the following two conjectures based on our extensive numerical experiments:
+We present the following conjecture based on our extensive numerical experiments:
 Conjecture 7.1.
 Any Kaleidocycle obtained by Theorem 6.1 has a minimal or maximal torsion angle. In other words, the projection from the space defined by (1.1) onto the $\lambda$-coordinate attains a critical value.
-Conjecture 6.2.
-The self-linking number of the closed curve $\gamma$ constructed in Theorem 4.2 is $m/2$ when the curve is unknotted.
-As briefly discussed in §1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [18, 20, 30].
+As briefly discussed in §1 and Remark 2.1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [20, 22, 31].
 Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
 Figure 2.
-(Left) An example of 8-Kaleidocycle with $(v,r,y,m)\approx(0.400,0.717,1.067,3)$ and $\cos\lambda\approx 0.4700$.
+(Left) An example of 8-Kaleidocycle with $(v,r,y)\approx(0.4000,-0.2834,1.067)$ having $m=3$ and $\cos\lambda\approx 0.4700$.
 (Middle) 8-Kaleidocycle as a discrete curve.
@@
 Figure 3.
-(Left) An example of 9-Kaleidocycle with $(v,r,y,m)\approx(0.342,0.289,1.027,3)$ and $\cos\lambda\approx 0.5852$.
+(Left) An example of 9-Kaleidocycle with $(v,r,y)\approx(0.3422,-0.2890,1.027)$ having $m=3$ and $\cos\lambda\approx 0.5852$.
 (Middle) 9-Kaleidocycle as a discrete curve.
@@
 Figure 4.
-(Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.189,0.300,0.947,3)$ and $\cos\lambda\approx 0.8533$.
+(Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.1894,-0.2996,0.9470)$ having $m=3$ and $\cos\lambda\approx 0.8533$.
 (Middle) 15-Kaleidocycle as a discrete curve.
@@
 Figure 5.
-(Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.401,0.689,1.204,5)$ and $\cos\lambda\approx 0.6497$.
+(Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.4012,-0.3113,1.204)$ having $m=5$ and $\cos\lambda\approx 0.6497$.
 (Middle) 15-Kaleidocycle as a discrete curve.
 (Right) the semi-discrete K-surface obtained as its trajectory.
-Appendix
-In the appendix, we prove the theorems and propositions in the paper by using the properties of elliptic theta functions listed in Appendix A.
 Appendix A Functional identities of elliptic theta functions
-We collect some properties of elliptic theta functions which we rely heavily on our calculations.
+In the appendix, we collect some properties of elliptic theta functions which we rely heavily on our calculations.
 A.1. Four term identities
@@
 (A.19) $=\vartheta_{2}\left(X;\frac{iy}{2}\right)\vartheta_{2}\left(Y;\frac{iy}{2}\right)-\vartheta_{1}\left(X;\frac{iy}{2}\right)\vartheta_{1}\left(Y;\frac{iy}{2}\right).$
-(A.20) $\begin{split}&\vartheta_{3}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{4}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{4}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ &\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)=\vartheta_{1}\left(X;iy\right)\vartheta_{4}\left(0;iy\right),\\ -&\vartheta_{1}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}+\vartheta_{2}\left(\frac{X}{2};\frac{iy}{2}\right)^{2}=2\vartheta_{2}\left(X;iy\right)\vartheta_{3}\left(0;iy\right).\\ \end{split}$
-Appendix B Proofs of Theorem 4.1 and Theorem 5.1
-Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (3.1)-(3.6).
-Theorem 4.1 follows from Theorem 5.1 by setting $t=0$.
-We assume $z=0$ without loss of generality except for the proof of (3.5).
-Putting
-(B.5) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+r+it,\\ \vspace{2mm}Y=ivn-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
-in (A.2) yields (3.1). Putting
-(B.10) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=ivn+r+it,\\ \vspace{2mm}U=iv,\\ V=\frac{1}{2}iv-r,\\ \end{array}\right.$
-in (A.2) yields (3.2). Putting
-(B.15) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=iv,\\ V=ivn-r+it,\\ \end{array}\right.$
-in (A.3) yields (3.3). Putting
-(B.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=-\frac{1}{2}iv+r,\\ \end{array}\right.$
-in (A.4) yields
-(B.21) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
-Also, putting
-(B.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=\frac{1}{2}iv+r,\\ \vspace{2mm}U=ivn-r+it,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
-in (A.4) yields
-(B.27) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(ivn-r+it\right)\vartheta_{4}\left(ivn-r+it\right)\\ &-\vartheta_{3}\left(ivn-r+it\right)\vartheta_{1}\left(ivn-r+it\right).\end{split}$
-Subtracting (B.27) from (B.21) and substituting
-(B.28) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
-in the equation, we obtain
-(B.29) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn-r+it\right),\end{split}$
-from which we see (3.4) holds. Putting
-(B.34) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn+z+r+it,\\ \vspace{2mm}Y=ivn+z-r+it,\\ \vspace{2mm}U=\frac{1}{2}iv+r+h,\\ V=\frac{1}{2}iv-r+h,\\ \end{array}\right.$
-in (A.2) yields
-(B.35) $\begin{split}&\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
-Differentiating (B.35) with $h$ and then putting $h=0$, we obtain
-(B.36) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(ivn+z+r+it\right)\vartheta_{2}\left(ivn+z-r+it\right)\\ &+\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{4}\left(ivn+z-r+it\right).\end{split}$
-Using the equation (B.35) with $h=0$ and substituting (B.36), we obtain
-(B.37) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{D_{z}\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\cdot\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right.\\ &\left.+\left(\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\right.\\ &\left.\times\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+z+it\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+z+it\right)\right\}\\ &=\left(\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}+\frac{\vartheta^{{}^{\prime}}_{3}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}-\frac{\vartheta^{{}^{\prime}}_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\right)\\ &\times\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(ivn+z-r+it\right),\end{split}$
-from which we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows.
-Appendix C Proof of Proposition 5.5
-For the proof of Proposition 5.5, we use the following lemma:
-Lemma C.1 ([11, §3]).
-The following equation holds:
-(C.1) $f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}=f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}=\beta F_{n}F_{n}.$
-C.1. Proof of Proposition 5.5
-Here we prove that the $\tau$ functions given in Theorem 5.1 satisfy (5.8)-(5.14). We assume $z=0$ without loss of generality except for the proof of (5.14). Putting
-(C.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=\frac{1}{2}iv-r,\\ \vspace{2mm}U=ivn+r+it,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
-in (A.8) yields (5.8). Also, (5.9) follows from (C.1) and (5.8). Putting
-(C.11) $\left\{\begin{array}[]{l}\vspace{2mm}X=ivn-r+it,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r+h,\\ \vspace{2mm}U=\frac{1}{2}iv-r+h,\\ V=iv\left(n-1\right)-r+it,\\ \end{array}\right.$
-in (A.3) yields
-(C.12) $\begin{split}&\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)-2r+it+h\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it-h\right)\\ &=\vartheta_{4}\left(ivn-r+it\right)\vartheta_{3}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{1}\left(\frac{1}{2}iv-r+h\right)\vartheta_{2}\left(iv\left(n-1\right)-r+it\right)\\ &-\vartheta_{2}\left(ivn-r+it\right)\vartheta_{1}\left(-\frac{1}{2}iv-r+h\right)\\ &\times\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\vartheta_{4}\left(iv\left(n-1\right)-r+it\right),\end{split}$
-After differentiating (C.12) by $h$ and putting $h=0$, we obtain (5.13) by rearrangement. Also, rearranging the equation with $h=0$ in (C.12) yields (5.10). Putting
-(C.17) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv-r,\\ \vspace{2mm}U=iv\left(n-\frac{1}{2}\right)+it,\\ V=-iv\left(n-\frac{1}{2}\right)-it,\\ \end{array}\right.$
-in (A.5) yields (5.11). Also, (5.12) follows from (5.8) and (5.11). Putting
-(C.22) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}Y=iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ \vspace{2mm}U=h,\\ V=h,\\ \end{array}\right.$
-in (A.2) yields
-(C.23) $\begin{split}&\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}+h\right)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}-h\right)\\ &=-\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{1}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}\\ &+\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}}\right)^{2}\frac{\vartheta_{3}\left(h\right)^{2}}{\vartheta_{3}\left(0\right)^{2}}.\end{split}$
-After differentiating (C.23) twice with $h$ and putting $h=0$, we obtain (5.14).
-Appendix D Proofs of Theorem 5.6 and Theorem 5.9
-For the proofs of Theorem 5.6 and Theorem 5.9, we need the following lemmas.
-Lemma D.1 ([11, §3]).
-Using (3.2) and (3.3), we see that the following equation holds.
-(D.1) $g_{n}f_{n-1}-f_{n}g_{n-1}=G_{n}F_{n}.$
-Lemma D.2.
-The following relation holds.
-(D.2) $\omega=\delta_{2}-\delta_{2}^{*}.$
-Proof.
-Putting
-(D.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=-\frac{1}{2}iv+r,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r,\\ \vspace{2mm}U=\frac{1}{2}iv+r,\\ V=\frac{1}{2}iv+r,\\ \end{array}\right.$
-in (A.9) yields (D.2). ∎
-Lemma D.3.
-The following identity holds.
-(D.8) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=2\delta_{2}\delta_{3}^{*}+2\delta_{2}^{*}\delta_{3}+\epsilon_{2}=$ 0 .
-Proof.
-Putting
-(D.13) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{1}{2}iv+r+h,\\ \vspace{2mm}Y=-\frac{1}{2}iv+r-h,\\ \vspace{2mm}U=\frac{1}{2}iv+r-h,\\ V=-\frac{1}{2}iv+r+h,\\ \end{array}\right.$
-in (A.9) yields
-(D.14) $\begin{split}&\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(\frac{1}{2}iv+r-h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r+h\right)\\ &-\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(-\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(\frac{1}{2}iv+r-h\right)\vartheta_{3}\left(-\frac{1}{2}iv+r+h\right)\\ &=\vartheta_{3}\left(2r\right)\vartheta_{3}\left(0\right)\vartheta_{1}\left(iv\right)\vartheta_{1}\left(2h\right).\end{split}$
-After differentiating (D.14) by $h$ and setting $h=0$, we obtain (D.8). ∎
-For simplicity, we define $p$ and $q$ as follows
-(D.15) $p_{n}=g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*},$
-(D.16) $q_{n}=g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}.$
-Then we rewrite $D_{t}H_{n}\cdot F_{n}$ by using (5.6), (D.15) and (D.16) as follows
-(D.17) $D_{t}H_{n}\cdot F_{n}=\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}.$
-Also, we see the following hold.
-(D.18) $\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}=T_{n}-\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}q_{n}^{*}+q_{n}\\ \vspace{2mm}\frac{1}{i}\left(q_{n}^{*}-q_{n}\right)\\ -2\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\end{array}\right),$
-(D.19) $\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\frac{1}{i}\left(p_{n}-p_{n}^{*}\right)\\ \vspace{2mm}p_{n}+p_{n}^{*}\\ -\frac{2}{i}\omega|G_{n}|^{2}\end{array}\right).$
-We have the following lemma.
-Lemma D.4.
-The following relations hold.
-(D.20) $p_{n}p_{n}^{*}=F_{n}^{2}|G_{n}|^{2}+\omega^{2}|G_{n}|^{4},$
-(D.21) $\begin{split}p_{n}q_{n}^{*}-p_{n}^{*}q_{n}=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
-(D.22) $\begin{split}&q_{n}q_{n}^{*}=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+\left\{1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
-Proof.
-By definition of $p$, we see that
-(D.23) $\begin{split}p_{n}p_{n}^{*}&=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\\ &=\left(f_{n}^{*}g_{n-1}^{*}-g_{n}^{*}f_{n-1}^{*}\right)\left(f_{n}g_{n-1}-g_{n}f_{n-1}\right)\\ &+\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\end{split}$
-holds. Substituting (D.1) and (5.9) into (D.23) yields (D.20). By definition of $p$ and $q$, we see that
-(D.24) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\left(g_{n}^{*}f_{n-1}-f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}^{*}-f_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\\ &=2\left(g_{n}^{*}f_{n-1}f_{n}^{*}g_{n-1}-g_{n}f_{n-1}^{*}f_{n}g_{n-1}^{*}\right)\\ \end{split}$
-holds. On the other hand, from (C.1) and (5.9), the following holds.
-(D.25) $\begin{split}&\beta\omega F_{n}^{2}|G_{n}|^{2}\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(g_{n}^{*}g_{n-1}-g_{n}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+g_{n}g_{n-1}^{*}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\\ &=\left(f_{n}^{*}f_{n-1}+g_{n}^{*}g_{n-1}\right)\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right).\\ \end{split}$
-Using (D.25), we obtain
-(D.26) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=f_{n}f_{n-1}^{*}f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}f_{n}^{*}f_{n-1}+g_{n}g_{n-1}^{*}g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}g_{n}^{*}g_{n-1}\\ &=\left(f_{n}f_{n-1}^{*}\right)^{2}-\left(f_{n}^{*}f_{n-1}\right)^{2}+\left(g_{n}g_{n-1}^{*}\right)^{2}-\left(g_{n}^{*}g_{n-1}\right)^{2}\\ &=\left(f_{n}f_{n-1}^{*}-f_{n}^{*}f_{n-1}\right)\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\\ &+\left(g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right).\end{split}$
-Substituting (5.9), (5.11) and (5.12) into (D.26) gives
-(D.27) $\begin{split}&p_{n}q_{n}^{*}-p_{n}^{*}q_{n}\\ &=\omega|G_{n}|^{2}\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}-g_{n}g_{n-1}^{*}-g_{n}^{*}g_{n-1}\right)\\ &=2\left(2\delta_{1}-\beta\right)\omega F_{n}^{2}|G_{n}|^{2}+2\omega\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{4},\end{split}$
-thus (D.21). Also from (D.1), (5.11) and (5.12), we see that
-(D.28) $\begin{split}&q_{n}q_{n}^{*}-F_{n}^{2}|G_{n}|^{2}\\ &=\left(g_{n}^{*}f_{n-1}+f_{n}g_{n-1}^{*}\right)\left(g_{n}f_{n-1}^{*}+f_{n}^{*}g_{n-1}\right)\\ &-\left(g_{n}f_{n-1}-f_{n}g_{n-1}\right)\left(g_{n}^{*}f_{n-1}^{*}-f_{n}^{*}g_{n-1}^{*}\right)\\ &=\left(f_{n}f_{n-1}^{*}+f_{n}^{*}f_{n-1}\right)\left(g_{n}g_{n-1}^{*}+g_{n}^{*}g_{n-1}\right)\\ &=\left(2\delta_{1}F_{n}^{2}+\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\left(2\left(\beta-\delta_{1}\right)F_{n}^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}\right)\\ &=4\delta_{1}\left(\beta-\delta_{1}\right)F_{n}^{4}+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)F_{n}^{2}|G_{n}|^{2}-\left(\delta_{2}+\delta_{2}^{*}\right)^{2}|G_{n}|^{4}\end{split}$
-holds. From (D.28), we get (D.22). ∎
-Using Proposition 5.5, Lemma D.2, Lemma D.3 and Lemma D.4, Theorem 5.6 and Theorem 5.9 can be proved as follows.
-D.1. Proof of Theorem 5.6
-By differentiating $\gamma$ with respect to $t$, we get
-(D.29) $\begin{split}&\dot{\gamma}_{n}(t)=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle D_{t}H_{n}\cdot F_{n}+D_{t}H_{n}^{*}\cdot F_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(D_{t}H_{n}\cdot F_{n}-D_{t}H_{n}^{*}\cdot F_{n}\right)\\ -D_{t}D_{z}F_{n}\cdot F_{n}\end{array}\right)\\ &=\frac{1}{F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -\left(\epsilon_{1}+C\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
-On the other hand, from (D.18) and (D.19), we see
-(D.30) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}+2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)|G_{n}|^{2}+2i\beta_{2}\omega|G_{n}|^{2}\end{array}\right)\end{split}$
-holds. From (D.8), (D.30) can be rewritten as follows
-(D.31) $\begin{split}&\beta_{1}\left(T_{n}-\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}\right)+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\\ &=\frac{1}{2\beta F_{n}^{2}}\left(\begin{array}[]{c}\vspace{2mm}\displaystyle\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\\ \vspace{2mm}\displaystyle\frac{1}{i}\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ -2\beta_{1}\left(\beta-2\delta_{1}\right)F_{n}^{2}-\epsilon_{2}|G_{n}|^{2}\end{array}\right).\end{split}$
-Thus (5.15) is obtained.
-D.2. Proof of Theorem 5.9
-We compute the condition satisfied by $C$ and $\Gamma$ when
-(D.32) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,$
-(D.33) $\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},$
-where $\rho\in\mathbb{R}$ is a constant. From (D.19) and (D.29) we see that
-(D.34) $\begin{split}2iF_{n}^{3}\widetilde{G}_{n}\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=&\left(p_{n}-p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)\\ &+\left(p_{n}+p_{n}^{*}\right)\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)\\ &+2\omega|G_{n}|^{2}\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}\\ =&2F_{n}^{2}|G_{n}|^{2}\left\{2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)\right\}\\ &+2\omega|G_{n}|^{4}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\end{split}$
-holds. For (D.32) to be satisfied, it is sufficient that the following equations hold.
-(D.35) $2\beta_{1}\left(2\delta_{1}-\beta\right)\omega+2i\beta_{2}+\omega\left(\epsilon_{1}+C\right)=0,\$
-(D.36) $02\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}=$ 0 .
-By comparing (D.36) and (D.8) we see that (D.35) is a necessary and sufficient condition for the realization of (D.32).
-By substituting (5.6) into (D.35), it can be rewritten in terms of $C$ and $\Gamma$ as follows.
-(D.37) $i\Gamma\left(\omega\left(2\delta_{1}-\beta\right)-\frac{\delta_{2}+\delta_{2}^{*}}{\omega}\right)+\epsilon_{1}+C+\left(\delta_{3}+\delta_{3}^{*}\right)\left(2\delta_{1}-\beta\right)-\frac{\delta_{3}-\delta_{3}^{*}}{\omega}=0.$
-Also, since
-(D.38) $\begin{split}&F_{n}^{4}\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\rangle\\ =&\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}+\beta_{1}q_{n}-i\beta_{2}p_{n}\right)^{2}\\ &-\left(\beta_{1}q_{n}^{*}+i\beta_{2}p_{n}^{*}-\beta_{1}q_{n}+i\beta_{2}p_{n}\right)^{2}\\ &+\left\{\left(\epsilon_{1}+C\right)F_{n}^{2}+\epsilon_{2}|G_{n}|^{2}\right\}^{2}\\ =&F_{n}^{4}\left\{16\beta_{1}^{2}\delta_{1}\left(\beta-\delta_{1}\right)+\left(\epsilon_{1}+C\right)^{2}\right\}\\ &+F_{n}^{2}|G_{n}|^{2}\left\{4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)\right.\\ &\left.-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)\right\}\\ &+|G_{n}|^{4}\left(-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}\right),\end{split}$
-holds, we see that for (D.33) holds, it is sufficient that
-(D.39) $4\beta_{1}^{2}\left(1+2\left(\beta-2\delta_{1}\right)\left(\delta_{2}+\delta_{2}^{*}\right)\right)-8i\beta_{1}\beta_{2}\omega\left(2\delta_{1}-\beta\right)+4\beta_{2}^{2}+2\epsilon_{2}\left(\epsilon_{1}+C\right)=0,$
-(D.40) $-4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}-8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)+4\beta_{2}^{2}\omega^{2}+\epsilon_{2}^{2}=0,$
-hold. Since
-(D.41) $\begin{split}&4\beta_{1}^{2}\left(\delta_{2}+\delta_{2}^{*}\right)^{2}+8i\beta_{1}\beta_{2}\omega\left(\delta_{2}+\delta_{2}^{*}\right)-4\beta_{2}^{2}\omega^{2}-\epsilon_{2}^{2}\\ =&\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}+\epsilon_{2}\right\}\left\{2\beta_{1}\left(\delta_{2}+\delta_{2}^{*}\right)+2i\omega\beta_{2}-\epsilon_{2}\right\},\end{split}$
-holds, (D.40) is identically satisfied.
-Therefore, (D.39) is a necessary and sufficient condition for the realization of (D.33).
-By using (D.8), (5.6) and (D.37), (D.39) can be written in terms of $\Gamma$ as follows.
-(D.42) $i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)\Gamma^{2}+2\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\Gamma+i\left(\delta_{2}{\delta_{3}^{*}}^{2}-\delta_{2}^{*}\delta_{3}^{2}\right)=0.$
-(D.42) has the following solutions.
-(D.43) $\Gamma=\frac{-\delta_{2}\delta_{2}^{*}\left(\delta_{3}+\delta_{3}^{*}\right)\pm\sqrt{\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}}}{i\delta_{2}\delta_{2}^{*}\left(\delta_{2}-\delta_{2}^{*}\right)}.$
-Since
-(D.44) $\begin{split}\delta_{2}\delta_{2}^{*}\left(\delta_{2}\delta_{3}^{*}+\delta_{2}^{*}\delta_{3}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)^{2}=\left(\delta_{2}\delta_{2}^{*}\right)^{2}\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2},\end{split}$
-holds, (D.43) can be rewritten as
-(D.45) $\Gamma=\frac{-\left(\delta_{3}+\delta_{3}^{*}\right)\pm\left(\delta_{3}e^{i\lambda}+\delta_{3}^{*}e^{-i\lambda}\right)}{i\omega}.$
-Now, we prove equations (5.20) and (5.21). From (4.2) and (5.6), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
-(D.46) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
-Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), the following relation holds
-(D.47) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{3}(r)^{4}-\vartheta_{1}(r)^{4}\right)^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}.$
-From (D.38), (D.46) and (D.47), we see that
-(D.48) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(r)^{4}\vartheta_{1}(r)^{4}}=\rho_{+}^{2}$
-holds. (D.48) is equivalent to (5.20). From (4.2) and (5.6), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
-(D.49) $16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}=4\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}.$
-Also, from (4.2), (5.6), (D.35) and (D.36), we see that when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), the following relation holds
-(D.50) $\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{3}(2r)^{2}\vartheta_{3}(0)^{6}}\frac{\left(\vartheta_{4}(r)^{4}-\vartheta_{2}(r)^{4}\right)^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}.$
-From (D.38), (D.49) and (D.50), we see that
-(D.51) $\rho^{2}=16\beta_{1}^{2}\left(\beta-\delta_{1}\right)\delta_{1}+\left(\epsilon_{1}+C\right)^{2}=\left(i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\right)^{2}\frac{u_{2}^{2}}{\vartheta_{4}(r)^{4}\vartheta_{2}(r)^{4}}=\rho_{-}^{2}$
-holds. (D.51) is equivalent to (5.21).
-Appendix E Proof of Proposition 5.12
-Here we prove that the signed curvature angle $\kappa$ satisfies (5.27). First, we see
-Lemma E.1.
-The $\tau$ functions (5.1) satisfy the following bilinear equations.
-(E.1) $\widetilde{G}_{n}^{2}+\beta^{2}F_{n}^{2}=F_{n+1}F_{n-1},$
-(E.2) $\beta^{2}\widetilde{G}_{n}^{2}-\frac{\vartheta_{1}\left(iv\right)^{4}}{\vartheta_{3}\left(iv\right)^{4}}\beta^{4}F_{n}^{2}=\widetilde{G}_{n+1}\widetilde{G}_{n-1}.$
-Proof.
-Using (3.1)-(3.3), we see that (E.1) hold. Putting
-(E.7) $\left\{\begin{array}[]{l}\vspace{2mm}X=0,\\ \vspace{2mm}Y=0,\\ \vspace{2mm}U=iv\left(n+\frac{1}{2}\right)+it+\frac{z}{\Delta_{3}-\Delta_{1}},\\ V=-iv\left(n-\frac{3}{2}\right)-it-\frac{z}{\Delta_{3}-\Delta_{1}},\\ \end{array}\right.$
-in (A.10), we obtain (E.2).
-E.1. Proof of Proposition 5.12
-Using (E.1) and (E.2), we see that (5.27) holds.
-Appendix F Proof of Theorem 5.13
-Here, we show that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28). For the proof of Theorem 5.13, we first see
-Lemma F.1.
-The $\tau$ functions (5.1) satisfy the following bilinear equation.
-(F.1) $D_{t}\widetilde{G}_{n}\cdot F_{n}=\eta_{3}\left(\widetilde{G}_{n+1}F_{n-1}-\widetilde{G}_{n-1}F_{n+1}\right).$
-Proof.
-We assume $z=0$ without loss of generality. After differentiating the equation (A.3) with
-(F.6) $\left\{\begin{array}[]{l}\vspace{2mm}X=iv\left(n+\frac{1}{2}\right)+it,\\ \vspace{2mm}Y=iv+h,\\ \vspace{2mm}U=-iv+h,\\ V=iv\left(n-\frac{3}{2}\right)+it\\ \end{array}\right.$
-by $h$ and putting $h=0$, we obtain
-(F.7) $\begin{split}&D_{t}\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\cdot\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it\right)\right\}\\ &=\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\\ &-\frac{\vartheta_{1}(iv)\vartheta_{3}(iv)}{\vartheta_{1}(2iv)\vartheta_{3}(0)}\frac{\vartheta^{{}^{\prime}}_{1}\left(iv\right)\vartheta_{3}\left(iv\right)-\vartheta_{1}\left(iv\right)\vartheta^{{}^{\prime}}_{3}\left(iv\right)}{i\vartheta_{1}\left(iv\right)\vartheta_{3}\left(iv\right)}\\ &\times\left\{-i\vartheta_{1}(iv)\vartheta_{4}\left(iv\left(n+\frac{1}{2}\right)+it\right)\right\}\left\{\vartheta_{3}(iv)\vartheta_{2}\left(iv\left(n-\frac{3}{2}\right)+it\right)\right\}.\end{split}$
-After rearranging (F.7), we obtain (F.1).
-F.1. Proof of Theorem 5.13
-Using (E.1) and (F.1), we see that the signed curvature angle satisfies the semi-discrete mKdV equation (5.28).
-Appendix G Proof of Theorem 5.14
-From (A.17), (A.18) and (A.19), we see that
-(G.1) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2}\end{split}$
-holds.
-Appendix H Proof of Theorem 5.15
-We first construct the solution to the semi-discrete potential mKdV equation (5.33) in two stages by proving Lemma H.1 and Lemma H.2.
-Then we construct the solution to the semi-discrete sine-Gordon equation (5.34) in two stages by proving Lemma H.3 and Lemma H.4.
-This is based on the construction of the solution to the mKdV equation using the theory of $\tau$ functions [15].
-For fixed $v,y$, define the following real-valued functions
-(H.1) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
-Then the potential function can be written as follows
-(H.2) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
-where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$.
-$\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy.
-Lemma H.1.
-Assuming that the solution of (5.33) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.33).
-(H.3) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
-(H.4) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
-where $\eta_{1}$ is a real constant.
-Proof.
-Since
-(H.5) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
-(H.6) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right)\end{split}$
-hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.33) is that
-(H.7) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\eta_{1}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
-is satisfied, where $\eta_{1}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
-(H.8) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\eta_{1}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\eta_{1}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
-The real part of this equation yields (H.3), and the imaginary part yields (H.4). ∎
-In Lemma H.2, we construct a solution of (H.3) and (H.4) in terms of elliptic theta functions.
-Lemma H.2.
-The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.3) and (H.4) with
-(H.9) $\eta_{1}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)+Q_{4}(v,y)}{L_{1}(v,y)}.$
-Proof.
-In this proof, all elliptic theta functions have $w=2i/y$ and we denote $\vartheta_{j}(v;2i/y)$ by $\vartheta_{j}(v)$. After differentiating the equation in (A.6) with
-(H.14) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
-by $h$ and putting $h=0$, we obtain
-(H.15) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
-Also, after differentiating the equation in (A.7) with
-(H.20) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}U=\frac{v}{y}+h,\\ V=h,\\ \end{array}\right.$
-by $h$ and putting $h=0$, we obtain
-(H.21) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
-Also, after differentiating the equation in (A.7) with
-(H.26) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
-by $h$ and putting $h=0$, we obtain
-(H.27) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
-Also, after differentiating the equation in (A.7) with
-(H.32) $\left\{\begin{array}[]{l}\vspace{2mm}X=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y},\\ \vspace{2mm}Y=\frac{v}{y}+h,\\ \vspace{2mm}U=\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y},\\ V=h,\\ \end{array}\right.$
-by $h$ and putting $h=0$, we obtain
-(H.33) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}Q_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
-Adding (H.15) and (H.21) yields (H.3), and subtracting (H.27) from (H.33) yields (H.4). ∎
-Next, we construct the solution to the semi-discrete sine-Gordon equation (5.34).
-Lemma H.3.
-Assuming that the solution of (5.34) is expressed as (H.2), we obtain the following bilinear equations as a sufficient condition for satisfying (5.34).
-(H.34) $D_{t}A_{n+1}\cdot A_{n}-D_{t}B_{n+1}\cdot B_{n}=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right),$
-(H.35) $D_{t}A_{n+1}\cdot B_{n}+D_{t}B_{n+1}\cdot A_{n}=\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right),$
-where $\eta_{2}$ is a real constant.
-Proof.
-Since
-(H.36) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}-\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}\right),\end{split}$
-(H.37) $\begin{split}&\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}}\right)\end{split}$
-hold, it can be seen that a sufficient condition for (H.2) to be a solution of (5.34) is that
-(H.38) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}_{n}=-\frac{L_{2}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}^{*}_{n}+\frac{L_{2}}{2y}\eta_{2}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n},\end{split}$
-is satisfied, where $\eta_{2}$ is a real constant. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
-(H.39) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}-B_{n+1}\cdot B_{n}\right)+iD_{t}\left(A_{n+1}\cdot B_{n}+B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{2}}{2y}\left(\eta_{2}-1\right)\left(A_{n+1}A_{n}-B_{n+1}B_{n}\right)\\ &+i\frac{L_{2}}{2y}\left(\eta_{2}+1\right)\left(A_{n+1}B_{n}+B_{n+1}A_{n}\right).\end{split}$
-The real part of (H.39) yields (H.34), and the imaginary part yields (H.35). ∎
-In Lemma H.4, we construct a solution of (H.34) and (H.35) in terms of elliptic theta functions.
-Lemma H.4.
-The real functions $A$ and $B$ defined by (H.1) are the solutions of (H.34) and (H.35) with
-(H.40) $\eta_{2}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)-Q_{4}(v,y)}{L_{2}(v,y)}.$
-Proof.
-In this proof, we set $w=2i/y$ similarly to Lemma H.2. Subtracting (H.21) from (H.15) yields (H.34), and adding (H.33) and (H.27) yields (H.35). ∎
-Appendix I Proof of Proposition 5.16
-We first show (5.40) and (5.39) by using the periodicity of elliptic theta functions (A.11). Consider replacing the parameter with $\tilde{v}=v+y$. Then we see
-(I.1) $\begin{split}&Q_{1}(\tilde{v},y)=-Q_{1}(v,y),\quad Q_{2}(\tilde{v},y)=-Q_{2}(v,y),\\ &d_{1}(\tilde{v},y)=d_{1}(v,y),\quad d_{4}(\tilde{v},y)=d_{4}(v,y),\\ &L_{1}(\tilde{v},y)=L_{2}(v,y).\end{split}$
-and
-(I.2) $\begin{split}&\Theta_{n}\left(t;\tilde{v},y\right)=4\arctan\frac{\vartheta_{1}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}{\vartheta_{4}\left(\frac{\tilde{v}}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)}\\ &=4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}+n-\frac{1}{2}\right)}\\ &=(-1)^{n}4\arctan\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}-\frac{1}{2}\right)}\\ &=(-1)^{n}\Theta_{n}\left(t-\frac{y}{2};v,y\right).\end{split}$
-From the relation between the signed curvature angle $\kappa$ and the potential function $\Theta$, we see
-(I.3) $\begin{split}&\kappa_{n}\left(t;\tilde{v}\right)=(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2};v\right).\end{split}$
-Next, we check the transformation of $\lambda$. From (A.13) and (A.14), we have
-(I.4) $\begin{split}&\lambda\left(\tilde{v},\frac{1}{2}-r\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}{\vartheta_{1}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+\frac{1}{2}-r+\frac{1}{2}iy\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+\frac{1}{2}-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(\frac{\vartheta_{2}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}{\vartheta_{2}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}\frac{\vartheta_{4}\left(\frac{1}{2}iv-r+\frac{1}{2}iy\right)}{\vartheta_{4}\left(-\frac{1}{2}iv-r-\frac{1}{2}iy\right)}\right)\\ &=\frac{1}{i}\log\left(-\frac{\vartheta_{3}\left(-\frac{1}{2}iv-r\right)}{\vartheta_{3}\left(\frac{1}{2}iv-r\right)}\frac{\vartheta_{1}\left(\frac{1}{2}iv-r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv-r\right)}\right)\\ &=\pi+\frac{1}{i}\log\left(\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)}\frac{\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\right)\\ &=\pi+\lambda\left(v,r\right).\\ \end{split}$
-Finally, we see that the transformation of the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ follows from the Frenet-Serret formula (2.6).
-Appendix J Proof of Proposition 5.18
-First, we calculate $L_{1}(v,y)/y$ and $L_{2}(v,y)/y$. From (5.36), (A.15), (A.16) and (A.20), we see
-(J.1) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
-Similarly, from (5.37), (A.15), (A.16) and (A.20), we see
-(J.2) $\begin{split}\frac{L_{2}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)-\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}-\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;i\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y};i\frac{2}{y}\right)}{\vartheta_{4}\left(0;i\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y};i\frac{2}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}-i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)-\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
-On the other hand, from (4.2) and (5.2), we see
-(J.3) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}},\end{split}$
-(J.4) $\begin{split}&1-\cos\lambda\\ &=-\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)-\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{2}}.\end{split}$
-From (J.1) and (J.3), we see that (5.43) holds. Similarly, from (J.2) and (J.4), we see that (5.44) holds.
-Appendix K Special property of elliptic theta functions when $w$ is a pure imaginary number
-Lemma K.1.
-The following holds for any $z\in\mathbb{C}$ when $w=iy$.
-(K.1) $\vartheta_{j}(z)^{*}=\vartheta_{j}(z^{*}),\quad j=1,2,3,4.$
-Proof.
-Consider the following theta function with characteristics.
-(K.2) $\vartheta_{a,b}(z):=\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z+b)+\pi i(n+a)^{2}\omega\right),\quad a,b\in\left\{0,\frac{1}{2}\right\},$
-then,
-(K.3) $\vartheta_{1}(z)=-\vartheta_{\frac{1}{2},\frac{1}{2}}(z),\quad\vartheta_{2}(z)=\vartheta_{\frac{1}{2},0}(z),\quad\vartheta_{3}(z)=\vartheta_{0,0}(z),\quad\vartheta_{4}(z)=\vartheta_{0,\frac{1}{2}}(z)$
-hold. Since $w^{*}=-w$, we have
-(K.4) $\begin{split}\vartheta_{a,b}(z)^{*}=&\sum_{n\in\mathbb{Z}}\exp\left(-2\pi i(n+a)(z^{*}+b)-\pi i(n+a)^{2}w^{*}\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(-n-a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\sum_{n\in\mathbb{Z}}\exp\left(2\pi i(n+a)(z^{*}+b)+\pi i(n+a)^{2}w\right)\\ =&\vartheta_{a,b}(z^{*}).\end{split}$
-This lemma guarantees that some $\tau$ functions, such as $F_{n}$, are real-valued functions.
+[…]

### v4

@@
 Abstract.
-We consider the configuration space of ordered points on the two-dimensional sphere that satisfy a specific system of quadratic equations.
-We construct periodic orbits in this configuration space using elliptic theta functions and show that they simultaneously satisfy semi-discrete analogues of mKdV and sine-Gordon equations.
-The configuration space we investigate corresponds to the state space of a linkage mechanism known as the Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle.
-A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra greater than five.
-Our approach is founded on the relationship between the deformation of spatial curves and integrable systems, offering an intriguing example where an integrable system is explicitly solved to generate an orbit in the space of real solutions to polynomial equations defined by geometric constraints.
+We study a configuration space consisting of ordered points on the two-dimensional sphere satisfying a system of quadratic constraints.
+We construct explicit periodic orbits in the configuration space using elliptic theta functions.
+The constructed orbits simultaneously satisfy semi-discrete analogues of the modified KdV and sine-Gordon equations.
+This configuration space arises as the state space of a linkage mechanism called a Kaleidocycle, and the constructed orbits describe the characteristic motion of the Kaleidocycle.
+A key consequence of our construction is the proof that Kaleidocycles exist for any number of tetrahedra $k\geq 6$.
+Our approach is founded on the relationship between the deformation of spatial curves and integrable systems.
+The motion of the mechanism is interpreted as a deformation of a closed discrete spatial curve with constant torsion angle.
+This provides an explicit example in which an integrable system is solved to generate periodic orbits in a real solution set of polynomial equations arising from geometric constraints.
 Key words and phrases:
@@
 Primary 53A04, 53A70; Secondary 53A17, 70B15, 37K25, 37K10, 35Q53
-The first-named author was partially supported by KAKENHI, Grant-in-Aid for Scientific Research (B) 25K00921 and (S) 25H00399.
-The second and third-named authors were partially supported by KAKENHI Grant-in-Aid for Scientific Research (C) 21K3229.
+The first-named author was partially supported by JSPS KAKENHI Grant Numbers JP25K00921 and JP25H00399.
+The second-named author was partially supported by JSPS KAKENHI Grant Number JP25K21661.
+The third-named author was partially supported by JSPS KAKENHI Grant Numbers JP25K00921 and JP25K17297.
 1. Introduction
-Given a natural number $k$ and a real number $1<\cos\lambda<1$, we consider the following space of configurations of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
+Given a natural number $k$ and a real number $0<\lambda<\pi$, we consider the following configuration space of $k+1$ ordered points on the 2-sphere $S^{2}$ satisfying certain quadratic constraints:
 (1.1) $\mathcal{C}^{\pm}_{k,\lambda}=\{(b_{0},\ldots,b_{k})\in(S^{2})^{k+1}\mid$ $\langle b_{n-1},b_{n}\rangle=\cos\lambda,\,(1\leq n\leq k),$
@@
 $b_{k}=\pm b_{0}\}.$
-Our main result states that for any $k\geq 6$, there exists $\lambda$ such that a circle can be embedded in $\mathcal{C}^{\pm}_{k,\lambda}$.
-Moreover, we explicitly construct a circle in terms of elliptic theta functions.
-This result implies the existence of a particular class of linkage mechanisms called Kaleidocycles (Fig.
+This configuration space naturally appears as the state space of a linkage mechanism called a Kaleidocycle.
+A Kaleidocycle consists of $k$ identical equifacial tetrahedra connected by hinge joints along pairs of opposite edges so that they form a ring (Fig.
 1 (Left)).
-These mechanisms consist of $k$ identical copies of equifacial tetrahedra connected by their opposing edges as hinge joints, forming a ring.
-Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial ones has remained an open problem.
-Since an element of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a Kaleidocycle with $k$-tetrahedra in a specific position (Section 2.1) and its motion is described by a path in the configuration space, our main result provides an explicit construction and its periodic motion for Kaleidocycles of any $k\geq 6$.
-The key observation is to identify a Kaleidocycle with a closed spatial polygonal curve formed by connecting the midpoints of the hinge edges.
-Then, the angles between the hinge edges are reminiscent of the torsion angle of a curve, and the turning angles are reminiscent of the signed curvature angle (§2.1).
-An element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a sequence of binormals.
-In this geometric formulation using spatial curves, the motion of a Kaleidocycle is modelled by a segment-length and torsion-angle preserving deformation of the corresponding closed polygonal curve of a constant torsion angle $\lambda$.
-In our construction, it plays the central role to compare the functional identities of elliptic theta functions with the compatibility conditions of the framing of a discrete spatial curve.
-It was known in [21] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (see §3.2).
-The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12]5).
-Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the discrete constant negative curvature surface (K-surface) studied in [2, 18].
-Both views combined, a flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ is explicitly integrated using elliptic theta functions.
-Our work connects seemingly distant fields of integrable systems and kinematics of linkage via an interesting object, the Kaleidocycle.
+The mechanism exhibits a characteristic turning motion.
+Although Kaleidocycles have been studied by several authors [5, 9, 30, 31], the rigorous proof of their existence beyond trivial cases has remained an open problem.
+Our main result (Theorem 6.1) states that for any $k\geq 6$ there exists $\lambda$ for which the configuration space $\mathcal{C}^{\pm}_{k,\lambda}$ contains an embedded circle.
+We explicitly construct this circle using elliptic theta functions.
+Since a point of $\mathcal{C}^{\pm}_{k,\lambda}$ corresponds to a particular configuration of the mechanism, a closed path in this space describes a periodic motion of the Kaleidocycle.
+Therefore our result provides a constructive proof of the existence of Kaleidocycles consisting of $k$ tetrahedra for all $k\geq 6$.
+The key observation is that a Kaleidocycle can be identified with a closed discrete spatial curve obtained by connecting the midpoints of the hinge edges.
+In this interpretation the hinge edges correspond to the binormal directions of the curve.
+The angles between consecutive hinge edges correspond to a torsion angle, while the turning angles of the polygon correspond to signed curvature angles (Section 2.1).
+Hence an element $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$ can be regarded as a sequence of binormal vectors of a closed polygonal curve with constant torsion angle $\lambda$.
+In this geometric formulation the motion of a Kaleidocycle is modelled as a deformation of a closed polygonal curve preserving both the segment length and the torsion angle.
+In our construction, the compatibility conditions of the discrete Frenet frame describing such a deformation are compared with the functional identities of elliptic theta functions.
+It was known in [20] that the semi-discrete modified KdV (mKdV) and the sine-Gordon equations are defined on the configuration space (Section 3.2).
+The main result of this paper can also be viewed as providing a solution to the system through an explicit construction based on the $\tau$ functions of the two-component KP hierarchy [12] (Section 5).
+Moreover, the sweeping surface of the motion corresponds to a semi-discrete analogue of the fully-discrete constant negative curvature surface (K-surface) studied in [2, 18].
+Combining both views, we explicitly integrate the flow generated by an integrable system on a real-algebraic set $\mathcal{C}^{\pm}_{k,\lambda}$ using elliptic theta functions.
+Our work connects seemingly distant fields of integrable systems and the kinematics of linkages via an interesting object, the Kaleidocycle.
 The structure of the paper is as follows.
-Section 2 introduces the definition of Kaleidocycles.
-They are identified with a certain class of discrete spatial curves.
-As essential invariants to describe the state of a Kaleidocycle, the signed curvature angle and torsion angle are defined together with the Frenet frame.
-Section 3 reviews the connection between the theory of curves and integrable systems.
-The construction of discrete spatial curves and their Frenet frames in terms of the $\tau$ functions is presented.
-A previous result that describes the motion of a Kaleidocycle using semi-discrete integrable equations is recalled.
-Section 4 constructs a set of $\tau$ functions with three parameters $v,r,y$ using elliptic theta functions.
-The resulting curve is shown to have a constant torsion angle.
-In Section 5, a family of time-parametrised curves is constructed by introducing an additional parameter $t$.
-We then study the time evolution equations satisfied by the vertex positions and curvature angle of the curve and show connections to previous research.
-Section 6 investigates a necessary and sufficient condition for the resulting curve to be closed for every $t$.
-We prove there exists a set of parameters that satisfies the closure condition for a given $k\geq 6$, which proves the existence of Kaleidocycles.
-Section 7 provides examples of Kaleidocycles and corresponding semi-discrete surfaces with figures.
-We present a conjecture (Conjecture 7.1) suggested by numerical experiments: the Kaleidocycles constructed in this paper seem to be so-called Möbius Kaleidocycles.
+Section 2 introduces Kaleidocycles and their description as discrete spatial curves.
+The signed curvature angle and torsion angle are defined together with the associated Frenet frame.
+Section 3 reviews the relationship between discrete curves and integrable systems through the $\tau$ functions, and recalls previous results describing Kaleidocycle motion by semi-discrete equations.
+Section 4 constructs a set of $\tau$ functions that lead to discrete curves with constant torsion angle using elliptic theta functions.
+Section 5 introduces a time parameter and derives the evolution equations satisfied by the curve.
+Section 6 proves our main result (Theorem 6.1) that for every $k\geq 6$ there exist parameters for the $\tau$ functions for which the curve closes, establishing the existence of Kaleidocycles.
+Finally, Section 7 presents numerical examples and a conjecture that the constructed solutions correspond to the so-called Möbius Kaleidocycles.
+We briefly explain the strategy of the proof of the main result, Theorem 6.1, which establishes the existence of Kaleidocycles for every $k\geq 6$.
+The argument consists of three main steps.
+First, in 2.5, we identify Kaleidocycles with closed discrete curves of constant torsion angle.
+Next, in Theorem 5.1, we construct a family of motions of curves having the desired geometric properties.
+Finally, in Section 6, we determine a choice of parameters for which these curves become closed.
 2. Algebraic and geometric models of Kaleidocycle
@@
 It can be folded from a sheet of paper and is an example of rigid origami; it rotates like a bubble ring while keeping all faces rigid.
-The origin of Kaleidocycle is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R.
+The origin of Kaleidocycles is obscure, but one of the first appearances in the literature is in Bricard’s paper [3], where he considers an essentially equivalent closed kinematic chain, now known as the Bricard 6R.
 For nearly a hundred years, the geometry and kinematics of this object have attracted the attention of researchers (see [30] and the references therein).
@@
 Remark 2.1.
-The solution space of (2.1) admits an action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$, and the expected dimension of the quotient is $3\cdot 2k-5k-3-\dim(O(3))=k-6$, where $5k$ is the number of edges and $3$ comes from the equation $x_{0}+x_{1}=0$.
-This simple counting estimate is often called Maxwell’s degree-of-freedom counting law [9].
-However, this estimate is sometimes violated in interesting examples.
-The Bricard 6R is a case of $k=6$.
-Its configuration space has dimension one, which is greater than the expected dimension of 0, giving an example of an overconstrained system  [9].
-When $k\geq 7$, it is numerically observed that the configuration space degenerates to a circle when the angle between adjacent hinge edges is set to a critical value [22].
-[^1] Some of the findings of [22] are also published in [31].
-For $k>7$, this class of Kaleidocycles, named the Möbius Kaleidocycles, provides a rare example of an underconstrained linkage mechanism whose configuration space has a smaller dimension than the expected one.
+The solution space of (2.1) admits a natural action of the orthogonal group $O(3)$ of $\mathbb{R}^{3}$. A naive dimension count then predicts that the quotient space should have dimension
+$3\cdot 2k-5k-3-\dim(O(3))=k-6,$
+where $5k$ is the number of edges and the term $3$ comes from the constraint $x_{0}+x_{1}=0$. This heuristic is commonly referred to as Maxwell’s degree-of-freedom counting law [9].
+It is well known, however, that this estimate can fail in nontrivial ways.
+A classical example is the Bricard 6R linkage ($k=6$), which is equivalent to the classical $6$-Kaleidocycle.
+Its configuration space is one-dimensional, exceeding the expected dimension 0, and hence it provides a prototypical example of an overconstrained mechanism [9].
+For $k\geq 7$, numerical evidence indicates that, when the angle between adjacent hinge edges attains a critical value, the configuration space degenerates to a circle.
+The existence of this single degree of freedom was first identified by the first-named author together with collaborators, leading to a subsequent patent application [19].
+Shortly thereafter, related results were disseminated in the literature [31].
+For $k>7$, this family of Kaleidocycles–referred to as Möbius Kaleidocycles [19]–thus provides a rare example of an underconstrained linkage, whose configuration space has strictly smaller dimension than that predicted by Maxwell’s count.
 We use the following elementary geometric property of equifacial tetrahedron:
@@
 Let $\gamma_{n}=(x_{2n}+x_{2n+1})/2\ (0\leq n<k)$ be the mid-point of the hinge edge.
-There exists $l>0$ such that $l|b_{n+1}\times b_{n}|=|\gamma_{n+1}-\gamma_{n}|$.
 By 2.2, $\gamma_{n+1}-\gamma_{n}$ is orthogonal to both $b_{n}$ and $b_{n+1}$.
-Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$.
+Therefore, $b_{n+1}\times b_{n}$ is parallel to $\gamma_{n+1}-\gamma_{n}$, and we have $l(b_{n+1}\times b_{n})=\gamma_{n+1}-\gamma_{n}$ for some $l\in\mathbb{R}$.
+Then, $l\sum_{n=0}^{k-1}(b_{n+1}\times b_{n})=\sum_{n=0}^{k-1}(\gamma_{n+1}-\gamma_{n})=0$.
 This means $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$.
-Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=l\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron.
+Conversely, given $(b_{0},\ldots,b_{k})\in\mathcal{C}^{\pm}_{k,\lambda}$, define $\gamma_{n}=\frac{l}{\sin\lambda}\sum_{m=0}^{n-1}(b_{m+1}\times b_{m})$, where $l$ is the distance of the opposing hinge edges of the tetrahedron.
 We set $x_{2n}=\gamma_{n}-\frac{l_{01}}{2}b_{n}$ and $x_{2n+1}=\gamma_{n}+\frac{l_{01}}{2}b_{n}$.
 Then, $(x_{0},\ldots,x_{2k+1})$ satisfies Equation 2.1.
-It is clear that the above constructions are mutually inverse. ∎
+It is clear that the above constructions are mutually inverse and $O(3)$ equivariant. ∎
 While the shape of the tetrahedron is determined by the triangle edge lengths $(l_{01},l_{12},l_{02})$, the configuration space depends only on the hinge angle $\lambda$.
@@
 We adopt this convention to avoid discontinuous changes in the framing during the motion of the Kaleidocycle; its motion passes through the state where the curvature vanishes.
-As in the case of a smooth curve, a discrete spatial curve of a unit segment length is determined by its torsion and curvature angles.
-Therefore, the motion of a discrete curve of a constant torsion angle is determined by the set of functions $\kappa_{n}(t)$, which can be freely chosen.
-In this language, a Kaleidocycle is equivalent to a discrete closed ($\gamma_{k}=\gamma_{0})$ spatial curve with constant torsion angle $\lambda$ determined by the binormal $\widetilde{B}_{n}=b_{n}$.
-Our approach is to utilise this identification.
-The fact that a Kaleidocycle is closed enforces interesting topological constraints.
-For example, the self-linking number must be a half-integer, and hence, stay constant during the motion.
-By the Călugăreanu-Fuller-White theorem [1], the writhing number of the curve must remain constant as well.
-The self-linking number also tells that the configuration space of a Kaleidocycle is not generally connected.
+As an immediate corollary of 2.3, we obtain the following identification.
+Proposition 2.5.
+A state of a Kaleidocycle is identified with a closed discrete spatial curve $(\gamma_{k}=\gamma_{0})$ of constant torsion angle $\lambda$ whose binormal vectors $\widetilde{B}_{n}$ satisfy (1.1) with $b_{n}=\widetilde{B}_{n}$.
+As in the smooth case, a discrete spatial curve of unit segment length is determined by its curvature and torsion angles.
+Accordingly, the motion of a discrete curve with constant torsion angle is governed by the family of curvature functions $\kappa_{n}(t)$.
+The fact that a Kaleidocycle is closed imposes interesting topological constraints.
+The signed curvature angles $\kappa_{n}(t)$ are therefore subject to non-trivial restrictions.
+In particular, the self-linking number is a half-integer, and hence remains constant during the motion.
+By the Călugăreanu–Fuller–White theorem [1], the writhe of the curve must also remain constant.
+Moreover, the self-linking number shows that the configuration space of a Kaleidocycle is not connected in general.
 3. Curves and integrable systems
@@
 In particular, if the torsion is a constant, the integrable deformation is governed by the mKdV equation and its hierarchy [10, 25, 29].
-The connections are also investigated in a discrete setting, where curves are discretised as piecewise linear (polygonal) curves.
+The connections are also investigated in a discrete setting, where curves are discretized as piecewise linear (polygonal) curves.
 For example, continuous deformations of discrete plane and spatial curves have been considered in [6], where the deformations are governed by the semi-discrete analogue of the mKdV and the NLS equations.
@@
 In [20], a particular motion of a Kaleidocycle satisfying the following conditions is considered:
-(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\tilde{\rho}^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
+(3.24) $\left\langle\dot{\gamma}_{n},\widetilde{B}_{n}\right\rangle=0,\quad\left\langle\dot{\gamma}_{n},\dot{\gamma}_{n}\right\rangle=\rho^{2},\quad|{\gamma}_{n+1}-{\gamma}_{n}|=1\quad(0\leq\forall n\leq k),$
 where $\rho\neq 0$ is a constant and $\dot{\gamma}_{n}=\dfrac{d\gamma_{n}}{dt}$.
@@
 4. An explicit formula of a discrete curve with a constant torsion angle in terms of elliptic theta functions
-Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in Theorem 4.2.
+Based on the representation of curves discussed in §3.1, we define certain $\tau$ functions satisfying (3.1)-(3.6) using elliptic theta functions to construct a family of discrete curves with a constant torsion angle in 4.2.
 4.1. A family of curves with a constant torsion angle
@@
 we see that $\Delta_{3}-\Delta_{1}\neq 0$.
-Theorem 4.2.
+Proposition 4.2.
 Put $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$ and define
@@
 (4.13) $\begin{split}\kappa_{n}&=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta_{1}F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right).\end{split}$
-This theorem follows as a corollary of Theorem 5.1 and will be proven in §5.1.
+This proposition follows as a corollary of Theorem 5.1 and will be proven in §5.1.
+Note that $\widetilde{G}$ and $G$ have the same simple zeros.
+Therefore, (4.12) remains valid even at points where $\widetilde{G}$ vanishes.
+Moreover, it follows from the zero distribution of the theta functions that both $\beta_{1}$ and $F_{n}$ are non-vanishing.
+Combining this with (3.1), it follows that (3.10) and (4.13) hold for any $v$,$r$,$y$, and $n$.
 Remark 4.3.
-Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discritised curvature (see, for example, [14]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
+Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discretized curvature (see, for example, [15]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
 […]
@@
 By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$).
-We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature (§5.2, §5.3, §5.4).
+We then derive the evolution equation for the vertex positions of the curve in Theorem 5.85.2), which reduces to the previously known equations presented in §3.2 under the specific specialization of the global motion of the curve, as detailed in Theorem 5.11.
+Furthermore, we derive the evolution equations for the (potential of the) curvature in Theorem 5.16 and Theorem 5.195.3, §5.4).
 Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [20] are discussed.
-This is also summarised in Table 1 in §6.
-5.1. Explicit formula of the curve with deformation parameter
+This is also summarized in Table 1 in §6.
+5.1. Explicit formula for the curve with deformation parameter
 We introduce the deformation parameter $t$ so that only the signed curvature angles change while the torsion angle is kept constant.
@@
 In particular, this depends neither on $n$ nor on $t$, and $\gamma$ has a constant torsion angle $\lambda_{n}(t)=\lambda_{0}(0)$. Moreover, the signed curvature angle is given by
-(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right).\end{split}$
+(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
+and has a period in $t$:
+(5.4) $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right).$
 Proof.
@@
 (5.41) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right.\\ &\left.+\Delta_{3}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\left(\Delta_{3}-\Delta_{1}\right)\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
-From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Theorem 4.2 follows from Theorem 5.1 by setting $t=0$. ∎
+From this we see (3.5) holds.
+By the definition of $G$ and $\widetilde{G}$, (3.6) follows.
+Periodicity of the signed curvature angle (5.4) follows from (A.12).
+4.2 follows from Theorem 5.1 by setting $t=0$.
+∎
 Remark 5.2.
@@
 (5.96) $\begin{split}P_{n}Q_{n}^{*}-P_{n}^{*}Q_{n}=2\left(2\beta_{3}-\beta_{1}\right)\beta_{2}F_{n}^{2}|G_{n}|^{2}+2\beta_{2}\left(\beta_{4}+\beta_{4}^{*}\right)|G_{n}|^{4},\end{split}$
-(5.96) $\begin{split}&Q_{n}Q_{n}^{*}=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
+(5.97) $\begin{split}Q_{n}Q_{n}^{*}=&4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
 Proof.
@@
     (5.108) $\dot{\kappa}_{n}=2\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\left(\tan\frac{\kappa_{n+1}}{2}-\tan\frac{\kappa_{n-1}}{2}\right).$
-Furthermore, if the left-hand-side of the equation (5.136) does not depend on $n$, the following holds:
+Furthermore, assume that
+(5.109) $\tan\frac{\kappa_{n+1}}{2}\tan\frac{\kappa_{n-1}}{2}-\tan^{2}\frac{\kappa_{n}}{2}+\tan\frac{\kappa_{n+1}}{2}\tan^{2}\frac{\kappa_{n}}{2}\tan\frac{\kappa_{n-1}}{2}$
+is independent of $n$ (see 5.14). Then,
   1. (3)
@@
 Proof.
-The proofs of (i) and (ii) follow from the same calculation as in Section 5 of [18]. We prove (iii). From the deformation equation (5.103) and the discrete Frenet-Serret formula (2.5), we see that
+The proofs of (1) and (2) follow from the same calculation as in Section 5 of [18]. We prove (3). From the deformation equation (5.104) and the discrete Frenet-Serret formula (2.5), we see that
 (5.110) $\begin{split}&-\dot{\lambda}\sin\lambda=\left\langle\dot{\widetilde{B}}_{n},\widetilde{B}_{n+1}\right\rangle+\left\langle\widetilde{B}_{n},\dot{\widetilde{B}}_{n+1}\right\rangle\\ &=-\sin\lambda\left(2\beta_{1}\eta_{1}\sin\lambda-2\beta_{1}\eta_{2}\cos\lambda\right)\times\left(\frac{1}{\sin\kappa_{n+1}}\tan\frac{\kappa_{n+2}}{2}\right.\\ &\left.+\frac{1}{\tan\kappa_{n+1}}\tan\frac{\kappa_{n}}{2}-\frac{1}{\tan\kappa_{n}}\tan\frac{\kappa_{n+1}}{2}-\frac{1}{\sin\kappa_{n}}\tan\frac{\kappa_{n-1}}{2}\right),\end{split}$
-holds, which vanishes because of (5.136). Therefore $\dot{\lambda}=0$.
+holds. Put $a_{n}=\tan\frac{\kappa_{n}}{2}$. Using
+$\sin\kappa_{n}=\frac{2a_{n}}{1+a_{n}^{2}},\qquad\tan\kappa_{n}=\frac{2a_{n}}{1-a_{n}^{2}},$
+the bracket in (5.110) becomes
+$\frac{1}{2}\Bigl(a_{n+2}a_{n}-a_{n+1}^{2}+a_{n+2}a_{n+1}^{2}a_{n}-a_{n+1}a_{n-1}+a_{n}^{2}-a_{n+1}a_{n}^{2}a_{n-1}\Bigr).$
+Hence it is equal to one half of the difference between (5.109) at indices $n+1$ and $n$. Therefore it vanishes if the quantity in (5.109) is independent of $n$, implying $\dot{\lambda}=0$.
 Remark 5.10.
-Proposition 5.9 (iii) does not hold in general without the assumption on (5.136).
-Indeed, there is a motion of a Kaleidocycle satisfying (5.103) such that the left-hand side of (5.136) depends both on $n$ and $t$.
+Proposition 5.9 (3) does not hold in general without the assumption on (5.109). Indeed, there is a motion of a Kaleidocycle satisfying (5.104) such that (5.109) depends both on $n$ and $t$.
 By adjusting the global motion of $\gamma$ by $C$ and $\Gamma$, the deformation equation of $\gamma$ can be made to satisfy the equation considered in the previous studies presented in §3.2.
@@
 Set $C$ and $\Gamma$ to be
-(5.109) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0,\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.\end{split}$
+(5.111) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)\\ &-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}},\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)\\ &-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}.\end{split}$
 Then, the curve constructed in Theorem 5.1 deforms according to the equation (3.24). The magnitude of $\dot{\gamma}_{n}$ is a constant and is given by
@@
 6. Existence of Kaleidocycles
-The constant torsion-angle and unit segment-length curves given in Theorem 5.1 are not necessarily closed.
-In this section, we derive a necessary and sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle.
-Furthermore, we prove that there exist parameters that satisfy the closure condition when $k\geq 6$.
-The overall approach is as follows: For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle has to have period $k$.
-From this, we see that $m:=kv/y$ is an integer (§6.1).
-Then, rewriting the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ assuming $m$ is an integer, we obtain two equations for v,r,yv,r,y6.2).
-By rearranging one of the equations, we verify that $r=r(y)$ is uniquely determined when $y$ is in a certain range (§6.3).
-Substituting this $r(y)$ into the other equation, we find that if $m\geq 3$, then there exists $y$ satisfying it (§6.4).
-By considering two transformations of $m$ and $n$ that preserve the shape of a curve, we see that it is sufficient to consider $0<m\leq k/2$.
-Finally, we show that under this range of $m$ and $k\geq 6$, there always exists a set of parameters with which the curve Theorem 5.1 is closed, and hence, gives a $1$-dimensional periodic motion of a Kaleidocycle (§6.5).
+The curves given in Theorem 5.1 are not necessarily closed.
+In this section, we derive necessary and sufficient conditions on the parameters $v$, $r$, and $y$ for the curve $\gamma$ to be closed, and hence to define a Kaleidocycle.
+We also prove that such parameters exist when $k\geq 6$.
+The overall strategy is as follows.
+First, we derive a necessary condition for closedness.
+For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle must be periodic with period $k$.
+This implies that $m:=kv/y\in\mathbb{Z}\setminus(k\mathbb{Z})$6.1).
+Assuming $m\in\mathbb{Z}\setminus(k\mathbb{Z})$, we then rewrite the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ as two equations in the parameters $v$, $r$, and yy6.2).
+One of these equations determines $r=r(y)$ uniquely when $y$ lies in a certain range (§6.3).
+Substituting this expression into the remaining equation, we show that a solution $y$ exists whenever $3\leq m\leq k/2$6.4).
+Finally, by considering two transformations of $m$ and $n$ that preserve the shape of the curve, we may assume without loss of generality that $3\leq m\leq k/2$.
+Combining these results, we conclude that for $k\geq 6$, there always exists a choice of parameters for which the curve in Theorem 5.1 is closed, thereby establishing the existence of a Kaleidocycle (§6.5).
 6.1. Period of the curvature angle
-We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$.
+We say a curve $\gamma$ is closed with a period $k\geq 2$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$.
 In this case, the absolute value of the curvature angles also has a period $k$.
@@
 (6.1) $\begin{split}\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}=0,\end{split}$
-holds for all $n$ and $t$. Note that there is no Kaleidocycle with period $1$, hence $k\geq 2$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
+holds for all $n$ and $t$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
 (6.2) $\begin{split}&\tan^{2}\frac{\kappa_{n+k}(t)}{2}-\tan^{2}\frac{\kappa_{n}(t)}{2}\\ &=\frac{\vartheta_{1}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}}{\vartheta_{3}\left(\frac{v}{y};\frac{i}{y}\right)^{2}\vartheta_{4}\left(0;\frac{i}{y}\right)^{2}}\left\{\mathrm{cn}^{2}\left[2\tilde{K}\left(\tilde{\mu}_{n}(t)+\frac{v}{y}k\right)\right]-\mathrm{cn}^{2}\left[2\tilde{K}\tilde{\mu}_{n}(t)\right]\right\}=0,\end{split}$
@@
 Thus, we only need to check that (6.2) holds for $n=0$.
-Since $\mathrm{cn}^{2}$ has a real period $2\tilde{K}$, $m=kv/y$ must be an integer.
+Since $\mathrm{cn}^{2}$ has the real period $2\tilde{K}$, it follows that $m=kv/y$ must belong to $\mathbb{Z}\setminus(k\mathbb{Z})$.
 Hereafter, let $v$ satisfy
@@
 6.2. Closure conditions
-Here we obtain the necessary and sufficient conditions satisfied by $v,r,y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
+Here we obtain the necessary and sufficient conditions satisfied by $v$,$r$, and $y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
 (6.4) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
 $n+k-2\displaystyle\frac{\partial\log F_{n+k}}{\partial z}=n-2\displaystyle\frac{\partial\log F_{n}}{\partial z}.$
-We rewrite the left-hand side further with (A.12) under the condition that $m$ is an integer. Then we obtain
+We further rewrite the left-hand side using $m$ and (A.12). Then we obtain
 (6.5) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}}\left(R_{1}R_{3}\right)^{-k}\exp\left(4mr\pi i\right)(-1)^{m},$
@@
 (6.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=0.$
-On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v,r,y$ satisfy $kv=my$, (6.6) and (6.10).
+On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v$, $r$, and $y$ satisfy $kv=my$, (6.6) and (6.10).
 Thus, they give necessary and sufficient condition for the curve given by Theorem 5.1 to be a closed curve for every $t\in\mathbb{R}$.
-6.3. An explicit solution to the condition (6.10)
+6.3. An explicit solution to the condition $\frac{\partial}{\partial r}\log(\Lambda+1)=0$
 Here we solve (6.10) with respect to $r$.
@@
 where $\xi=\vartheta_{2}\left(0\right)^{2}/\vartheta_{3}\left(0\right)^{2}$ is an elliptic modulus and $K=\pi\vartheta_{3}\left(0\right)^{2}/2$ is a complete elliptic integral of the first kind.
-We define the following function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
+We define the following continuous real-valued function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
 (6.15) $\begin{split}\varphi\left(\frac{m}{k},y\right):=\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}&=\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}.\end{split}$
-When $1<\varphi\left(m/k,y\right)<1$ is satisfied, $r=r(y)$ is uniquely determined because the inverse function of $\mathrm{cn}$ exists.
-6.3.1. Existence of a inverse function
+When $1<\varphi\left(m/k,y\right)<1$ is satisfied, (6.13) determines a unique $r=r(y)$ in $(0,1/2)$ because on the interval $0<r<1/2$, the function $r\mapsto\mathrm{cn}(4Kr,\xi)$ is strictly decreasing from $1$ to 1-1.
+6.3.1. Existence of an inverse function
 Here we show that for a certain range of $y$, $1<\varphi\left(m/k,y\right)\leq 0$ holds when $0<m\leq k/2$.
@@
 where we denote $q=e^{-\frac{\pi}{y}}$. From this we have
-(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<1/2).$
+(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<\frac{1}{2}).$
 On the other hand, the following hold when $y\rightarrow\infty$
@@
 $ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=\pi p^{\left(\frac{1}{2}-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(\frac{1}{2}+\frac{m}{k}\right)^{2}}),$
-$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<1/2),$
+$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<\frac{1}{2}),$
 where we denote $p=e^{-\pi y}$. From this we have
-(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<1/2).$
+(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<\frac{1}{2}).$
 Hence, when $m$ is fixed within $0<m<k/2$, there exists $y$ such that $\varphi\left(m/k,y\right)=-1$.
@@
 6.4. Existence of a solution to the condition $\Lambda=0$
-By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$. Note that $\Lambda+1$ can also be written as
+By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that when $3\leq m\leq k/2$, $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$.
+Note that $\Lambda+1$ can also be written as
 (6.22) $\begin{split}\Lambda+1&=\exp\left\{\pi im\left(4r+1\right)\right\}\left[\frac{\vartheta_{1}\left(\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(\frac{1}{2}iv+r;iy\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r;iy\right)\vartheta_{3}\left(-\frac{1}{2}iv+r;iy\right)}\right]^{k}\\ &=\exp{\left(\pi im\right)}\left[\frac{\vartheta_{1}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}{\vartheta_{1}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)\vartheta_{3}\left(-\frac{1}{2}\frac{m}{k}+\frac{r}{iy};\frac{i}{y}\right)}\right]^{k},\end{split}$
@@
 (6.23) $\Lambda+1=\exp{\left(i\Psi\right)},$
-using a real-valued function $\Psi=\Psi(v,r,y)$. When
+using a continuous real-valued function $\Psi=\Psi(v,r,y)$. When
 (6.24) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|,$
 is greater than $2\pi$, then there exists a solution to $\Lambda(v,r(y),y)=0$ in the range $0<y<y_{0}(m/k)$.
-Here, we compute the above quantities and show that such a $y$ exists if $m\geq 3$.
+Here, we compute the above quantities and show that such a $y$ exists if $3\leq m\leq k/2$.
 From a representation by an infinite product of elliptic theta functions,
@@
 (6.26) $\begin{split}&\frac{\vartheta_{1}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{1}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\frac{\sin\left(\pi\frac{r}{iy}+\pi\frac{m}{2k}\right)}{\sin\left(\pi\frac{r}{iy}-\pi\frac{m}{2k}\right)}\exp{\left(-2i\sum_{n=1}^{\infty}\frac{1}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{2n(1-r)}-q^{2n(1+r)}\right)\right)},\\ &\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}\\ &=\exp{\left(-2i\sum_{n=1}^{\infty}\frac{(-1)^{n}}{n}\frac{\sin(\pi n\frac{m}{k})}{1-q^{2n}}\left(q^{n(1-2r)}-q^{n(1+2r)}\right)\right)}.\end{split}$
-Furthermore, from
+From the explicit expression of $r$ in (6.20), it follows that $1/4\leq r<1/2$ when $0<m\leq k/2$, which yields the following limit.
 […]
-we obtain
+This calculation remains valid even when $r$ is determined by (6.21), that is, in the case where $1/2<r\leq-1/4$. Therefore, we obtain
 (6.28) $\lim_{y\to+0}\frac{\vartheta_{3}\left(\frac{r}{iy}+\frac{m}{2k};\frac{i}{y}\right)}{\vartheta_{3}\left(\frac{r}{iy}-\frac{m}{2k};\frac{i}{y}\right)}=1.$
@@
 (6.32) $\left|\Psi(v,r(y_{0}),y_{0})-\lim_{y\to+0}\Psi(v,r(y),y)\right|=\pi m.$
-Therefore, when $m\geq 3$, there is always a solution in $\Lambda(v,r(y),y)=0$.
+Therefore, when $3\leq m\leq k/2$, there is always a solution in $\Lambda(v,r(y),y)=0$.
 6.5. Transformations of a parameter and a variable
@@
 In the following, we will construct the self-bijections and prove that the Kaleidocycles obtained by the self-bijections are essentially identical to the Kaleidocycles before the transformation.
-A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarised in Table 1.
+A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarized in Table 1.
 6.5.1. Transformation $m\rightarrow m+k$, $n\rightarrow k-n$
@@
 That is, the new Kaleidocycle is obtained by shifting the time parameter, reversing the orientation of the central curve, and flipping every other hinge, which is equivalently encoded by the torsion angle update $\lambda\mapsto\pi-\lambda$.
+The argument remains valid when $r$ is calculated by (6.20).
 6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
@@
 That is, the new Kaleidocycle is obtained by reversing the time parameter and flipping all hinges.
-Summarising the results of this section, we have the following theorem.
+Summarizing the results of this section, we have the following theorem.
 Theorem 6.1.
-For a fixed $k\geq 6$, let $0<m\leq k/2$.
-Then there exist $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $y>0$ satisfying $kv=my$, (6.6) and (6.10).
-Then, the curve defined by Theorem 5.1 is closed for every $t\in\mathbb{R}$ and forms a $k$-Kaleidocycle with the torsion angle given by (5.2).
+For $k\geq 6$ and an integer satisfying $3\leq m\leq\frac{k}{2}$, there exist $y>0$, $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ such that $kv=my$ and (6.6) and (6.10) hold.
+Hence the curve given in Theorem 5.1 is closed for every $t\in\mathbb{R}$ and defines a $k$-Kaleidocycle whose torsion angle is given by (5.2).
 Remark 6.2.
@@
 This is also consistent with the result of the transformation of $\Gamma$ and $C$.
-Table 1 summarises the relationships among the three deformation equations.
+Table 1 summarizes the relationships among the three deformation equations.
 Table 1. Relation between the three deformation equations
@@
 $\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
-$\downarrow$ Specialise the deformation equation by choosing
+$\downarrow$ Specialize the deformation equation by choosing
 rigid transformation parameters $C$ and $\Gamma$ as in Theorem 5.11.
@@
 Appendix A Functional identities of elliptic theta functions
-In the appendix, we collect some properties of elliptic theta functions which we rely heavily on our calculations.
+In this appendix, we collect several properties of elliptic theta functions on which our calculations rely heavily.
 A.1. Four term identities
@@
 […]
+Conflicts of Interest
+The authors declare no conflicts of interest.
+Data Availability Statement
+The computer program used in the experiments is available in the associated GitHub repository https://github.com/shizuo-kaji/Kaleidocycle/.